### v1

# Demystifying Behavior-Based Malware Detection at Endpoints

Abstract.
Malware detection is one of the most ubiquitous applications of machine learning (ML) in security. Prior ML-based detectors can consume program features extracted statically from binary code or dynamically from execution traces. Dynamic detectors most commonly rely on executing programs in synthetic environments (sandboxes) that allow controlled malware analysis. However, sandbox traces are unavailable to the last line of defense offered by security vendors: real-time malware detection at endpoints. A detector at endpoints consumes the traces of programs running on hosts (i.e., real-world environments), as sandbox analysis, typically done in the cloud, might introduce intolerable delays. Despite their success in the sandbox setting, research hints at potential challenges for ML methods at endpoints, e.g., highly variable malware behaviors in the wild. Nonetheless, the impact of these challenges on existing approaches and how their excellent sandbox performance translates to the endpoint scenario remain unquantified.
We present the first measurement study of the performance of ML-based malware detectors at real-world endpoints. Leveraging a dataset of sandbox traces we collected and a dataset of in-the-wild program traces from Avllazagaj et al. (avllazagaj2021malware, ); we evaluate two scenarios where the endpoint detector was trained on (i) sandbox traces (convenient and accessible); and (ii) endpoint traces (less accessible due to needing to collect telemetry data). This allows us to identify a wide gap between prior methods’ sandbox-based detection performance—over 90%—and endpoint performance—about 20% (scenario (i)) to 50% (scenario (ii)). We pinpoint and characterize the challenges contributing to this gap, including label noise, behavior variability, sandbox evasion, or environment-specific features. Building on this, we suggest strategies to close the gap, yielding a relative improvement of 5–30% over the baselines. Our evidence suggests that applying detectors trained on sandbox data to endpoint detection is challenging. The most promising direction is training detectors directly on endpoint data, which marks a departure from current practice. To promote research, we implement a leaderboard for realistic detector evaluations—a format that has fueled progress in many ML tasks.
malware detection, dynamic analysis, machine learning
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Security and privacyccs: Security and privacy Intrusion/anomaly detection and malware mitigationccs: Security and privacy Malware and its mitigation
1. Introduction
Detection of malware threats is crucial for governments, enterprises, and end users as there are significant (and growing) financial and safety harms of malware infections (reddick_2022, ), which has created a \$7 Billion industry in 2022 with many players (imarcreport, ). Malware detection appears to be remarkably effective: industry-standard evaluations of commercial anti-malware products (avcomparisons, ), and prior research on malware detection using machine learning (ML) methods (raff2018malware, ; karbab2019maldy, ), routinely report that over 99% of malware samples in a standard corpus can be detected, with very few false-positives.
Training data Test data Method TPR @ 1%FPR
Sandbox traces Sandbox traces Standard classifier $\sim$95(karbab2019maldy, ; jindal2019neurlux, ; trizna2023nebula, )
Sandbox traces Endpoint traces Standard classifier $\sim$17(ours)
Endpoint traces Endpoint traces Standard classifier $\sim$49(ours)
Sandbox traces Endpoint traces Training set resampling + invariant learning $\sim$22(ours)
Endpoint traces Endpoint traces Soft labels + invariant learning $\sim$52(ours)
Table 1. Performance of ML-based behavioral malware detection under different settings.
Malware detection is often performed at endpoints where an endpoint security solution (quarantine, ) monitors host devices to detect threats in real time. In practice, these solutions rely on a chain of techniques (vmray, ), including static analysis and dynamic analysis. Conventional static methods, such as blocklists, reputation systems, and signatures, operate without executing the program. As static methods can be bypassed through obfuscation or polymorphism (moser2007limits, ; song2007infeasibility, ), dynamic analysis has become standard, offered by most vendors (kasperskyrealtime, ).
Dynamic analysis relies on observing the execution behaviors of a sample to detect whether it displays malicious behaviors. The typical paradigm for obtaining such behavior-based detectors is to detonate (execute) a large set of known samples (malware and benign) in a controlled environment (a sandbox), collect their execution traces, and learn to separate malicious and benign behaviors. Several ML models have demonstrated exceptional performance in this task (mariconti2016mamadroid, ; karbab2019maldy, ; jindal2019neurlux, ; trizna2023nebula, ). Unfortunately, sandbox traces are unavailable to models in endpoint detection as detonation cannot be done in real time, despite advances such as on-premise sandboxes (onpremsandbox, ). Consequently, these models must rely on traces observed on real-world hosts to detect malware.
Research has outlined several challenges associated with this task. First, program behaviors are environment-sensitive, and, especially for malware, two traces of the same sample from two different hosts often have little overlap (avllazagaj2021malware, ). As a result, a model trained on a trace captured on one host might fail to generalize to other hosts. Second, techniques frequently employed to evade sandbox analysis (lindorfer2011detecting, ; balzarotti2010efficient, ; galloro2022systematical, ) can mislead models trained on sandbox traces into learning features that do not capture actual malware behaviors. Although these challenges were acknowledged in prior work (barecloud, ; kirat2015malgene, ; jindal2019neurlux, ; avllazagaj2021malware, ), we lack understanding of how they manifest, their implications for endpoint detectors, and potential solutions.
Our work conducts the first quantitative study into the efficacy of ML-based methods for endpoint detection to demystify these challenges. We specifically focus on two scenarios: (i) an endpoint detector trained using only sandbox traces and (ii) an endpoint detector trained using real-world endpoint traces. Existing techniques can conveniently gather large-scale datasets of sandbox traces, making (i) a practical and accessible option. On the other hand, collecting endpoint traces is attainable mainly by vendors who receive telemetry data from the wild as it is still an unsolved problem to simulate such data in controlled environments (such as sandboxes) at scale (schreuders2017security, ; miramirkhani2017spotless, ; mills2020investigating, ). These requirements make (ii) a less accessible but likely a more successful option. Consequently, both scenarios are relevant in practice, which motivates us to study their challenges.
We employ three deep-learning-based approaches (karbab2019maldy, ; jindal2019neurlux, ; trizna2023nebula, ) that can achieve $\sim$95true-positive rate (TPR) at 1% false-positive rate (FPR) when evaluated on sandbox traces. We use a dataset from Avllazagaj et al. (avllazagaj2021malware, ) of around 1M endpoint traces from over 25K samples, recorded on real-world hosts by a commercial anti-malware vendor. The malicious samples in this dataset were undetected by the vendor’s defenses at that time, corresponding to real-world infections. Consequently, this data reflects the realistic threats that behavior-based detectors must combat. Additionally, to study scenario (i), we collect a dataset of traces (contemporaneous to our endpoint dataset) from two sandboxes.
Our initial finding is that ML methods exhibit lower performance in both scenarios (i) ($\sim$17TPR) and (ii) ($\sim$49TPR), in contrast with their excellent performance on sandbox traces (see Table 1). We study the low performance in (i) from the lens of distribution shift, a problem that plagues ML in many applications (koh2021wilds, ). Others in the security community have articulated the challenge of concept drift, where the data distribution shifts over time (jordaney2017transcend, ). We explore whether the problem can be addressed using existing tools for concept drift. However, we found that existing concept drift methods are not enough to address the differences in data distribution between sandbox vs endpoint traces: we discovered that endpoint traces of benign samples are detected as “drifting” (whereas in prior work concept drift has mostly affected malware (barbero2022transcending, )), and rejecting drifting samples does not improve the performance enough.
Investigating the root causes of the low performance, we discovered that endpoint detectors are applied to a different distribution of samples than considered in prior work: they are typically applied only to the hardest-to-classify samples. In past research, researchers have trained and evaluated detectors on a corpus of samples from repositories (karbab2019maldy, ; jindal2019neurlux, ; dambra2023decoding, ) such as MalwareBazaar (bazaar, ). However, endpoint malware detection systems employ a pipeline, where samples are first categorized using basic methods (e.g., signatures, static analysis), and ML-based classifiers are only applied to samples that cannot be resolved using basic methods (vmray, ). Consequently, in the wild, endpoint classifiers are only applied to samples that tend to be harder to classify than those found in a standard corpus. Prior research has not considered this factor, and we find that it causes a significant drop in performance: it reduces the performance of classifiers trained on sandbox traces from 95% (for samples from a standard corpus) to $\sim$60TPR (when the distribution of samples is adjusted to take into account earlier stages in the pipeline). This shows that prior evaluations of behavior-based detectors have greatly over-estimated their effectiveness in the wild.
Second, we study the impact of variable program behaviors across different environments. We measure large differences between a sample’s sandbox trace and its endpoint traces. Sandbox traces lack diversity: collecting multiple traces by running a group of related samples, e.g., from the same malware family, in a sandbox yields very similar traces. This introduces spurious features that do not generalize to other environments. Conversely, endpoint traces are highly diverse, making a model’s predictions on different traces of the same sample inconsistent, which hurts detection performance. Moreover, we provide the first evidence that sandbox-evading malware (40–80% of all malware (galloro2022systematical, )) skews a model in scenario (i) towards classifying very short traces (often an indication of evasion (kuchler2021does, )) as malware. This correlation is spurious as it is absent in endpoint traces. Thus, this correlation causes prior evaluations using sandbox traces to over-estimate the accuracy of endpoint classifiers.
Our characterization of these challenges allows us to pinpoint avenues for improvement in endpoint detection. To improve the performance on the distribution of difficult-to-classify samples, we explore soft-labeling (effective against label noise (lukasik2020does, )) and using more accurate distributions for training. Against variable behaviors, we employ a technique popular in other areas of ML: invariant learning (zhao2019learning, ). In particular, we train our models to make consistent predictions on different traces of the same sample to prevent them from learning unreliable, environment-specific features. When combined, these strategies lead to moderate but promising gains, from 17% to 22% TPR in scenario (i) (30% relative improvement) and from 49% to 52% TPR in scenario (2) (5% relative improvement).
We believe our results should serve as a call to action for the research community. Previous research has suggested that it is possible to achieve 95% TPR, which might leave an impression that progress is saturated, there is not much room for further improvement, and the problem of ML-based malware detection is more or less solved. Our results show that the reality is different: the problem is not solved, the actual performance on malware in the wild is far worse, and there are major unsolved challenges and significant room for further improvement. To stimulate further research on this challenging problem, we have designed a leaderboard (malwaredetectioninthewild.github.io (leaderboard, )), a popular format in ML (croce2021robustbench, ), for realistic evaluations of behavior-based malware detectors. Participants receive access to our sandbox dataset and all our metadata and can submit their detectors, which we will rank according to their endpoint performance. This will hopefully allow the community to identify promising directions.
Contributions. (I) We measure the performance discrepancy of ML-based approaches between sandbox-only and endpoint settings (Section 5). (II) We identify the challenges in endpoint detection to demystify this discrepancy and propose improvements (Section 6 and Section 7). (III) We organize a leaderboard for realistic evaluations of detectors (Section 8).
2. Background and Related Work
Dynamic Malware Analysis. Most work in dynamic analysis focuses on analyzing the behaviors of a sample detonated in controlled environments, such as sandboxes (willems2007toward, ; bayer2009view, ). As dynamic analysis has become a staple, malware has started including checks that suppress malicious activities if the environment is fingerprinted as a sandbox, known as evasive malware (lindorfer2011detecting, ). Although many strategies have been developed to prevent fingerprinting (liu2022enhancing, ), this is still an ongoing arms race (miramirkhani2017spotless, ). Although researchers have explored methods to analyze samples in bare metal environments (balzarotti2010efficient, ; barecloud, ), they fall short in preventing evasion (miramirkhani2017spotless, ). To our knowledge, we have taken the first step toward measuring the implications of evasion for ML-based detectors at endpoints. A recent large-scale measurement study over variable program behaviors in the wild has found that a given malware sample can behave significantly differently across time and in different real-world hosts (avllazagaj2021malware, ). Our work connects to this line of work as we are interested in quantifying the impact of these challenges, such as evasion or variability, on a malware detector deployed for endpoint detection at hosts in the wild.
ML for Malware Detection. ML-based methods are widely used in research and practice with static (raff2018malware, ) and dynamic features (huang2016mtnet, ). In detection with dynamic features, methods that essentially treat a program’s execution trace as a natural language document (a sequence of tokens) and adapt off-the-shelf methods from the ML community have been successful (miller2016reviewer, ; karbab2019maldy, ; jindal2019neurlux, ; kuchler2021does, ; trizna2023nebula, ). Most of these methods are trained and evaluated on the traces from a pre-configured sandbox on samples collected from public repositories; some popular ones are listed here (awesomemalware, ). We aim to understand the implications of these practices and the efficacy of popular ML approaches for detecting malware in real-world hosts with dynamic features.
Distribution Shifts in ML. Distribution shift occurs when a model’s training and testing distributions have significant differences, which hurts the performance (koh2021wilds, ). For example, as new malware variants emerge and old ones disappear over time, the performance of a detector that has not been kept up-to-date will deteriorate, known as concept drift (pendlebury2019tesseract, ). A popular paradigm to deal with concept drift is deploying drift detectors to reject samples that would have been misclassified and training on them later to update the model (jordaney2017transcend, ; barbero2022transcending, ; chen2023continuous, ). We deploy a drift detector to characterize the differences between the distribution shifts in our problem and concept drift. Although there is no silver bullet, the ML community has also proposed many ideas to tackle distribution shifts in different contexts, such as learning domain invariant features (ganin2015unsupervised, ; 2013gongconnecting, ), distributionally robust optimization (sagawa2020investigation, ), or continuous learning (chen2023continuous, ). We adapt some of these ideas, such as invariant learning, to our problem to measure their benefits for malware detection at endpoints.
Limitations of ML for Security. Research suggests that popular ML methods have many pitfalls when applied to security tasks (arp2020and, ). For example, in malware detection, most work has been found to overestimate the success of the methods due to impossible time splits of training and testing sets (pendlebury2019tesseract, ). A line of work also focuses on how ML methods successful in lab-only evaluations break down when they are deployed in the real world due to distribution shifts, for example, in the context of network anomaly detection (jacobs2022ai, ), website fingerprinting (cherubin2022online, ) or malware detection (das2019sok, ; aghakhani2020malware, ). In these contexts, models are known to learn spurious features, such as specific IP addresses (jacobs2022ai, ), that are artifacts of the experimental setup in the lab and do not apply to realistic settings. Our work also investigates the limitations of ML methods in malware detection with dynamic features when they are trained and evaluated in controlled settings (e.g., using sandboxes) and deployed in the wild.
3. Endpoint Malware Detection
Terminology. A sample is an executable program, identified by its unique SHA-256 hash. A trace is a sequence of behavioral actions (such as file accesses or process creations) performed by a sample when it is executed in a computing environment. Our work focuses on Windows environments and executable samples. We refer to an in-the-wild endpoint environment as a host. In our endpoint dataset, the anti-malware system in each host recorded the traces of samples executed by the host that were not determined to be benign or malicious in an earlier stage of the detection pipeline. Each sample can have multiple traces recorded at multiple hosts at different timestamps. A sandbox (SB) is a synthetic environment that provides tools for analyzing samples and recording their traces through controlled execution. Our sandbox dataset contains traces collected from two sandboxes from two vendors, namely Tencent HABO (habo, ) and an in-house version of Cuckoo Sandbox (cuckoo, ). In the rest of the paper, we will refer to these sandboxes as SB1 and SB2, respectively.
Notation. We denote the set of samples in our dataset as $\mathbf{P}$. $P_{i}$ is the $i$-th program and $y_{i}$ is its ground truth label, where $y\!=\!0$ and $y\!=\!1$ indicate a benign and malware sample, respectively. If $P_{i}$ is a malware sample, it is also tagged with a family label $s_{i}$ that is useful for grouping samples with similar characteristics. A trace of $P_{i}$ is $x_{i}^{j}$, where $j$ denotes the execution environment, specifically, $j\!=\!0$ refers to the endpoint hosts and $j\!\in\!\{1,2\}$ refers to the sandboxes. As there are multiple endpoint traces per sample, we refer to the $k$-th one as $x_{i,k}^{0}$, and the endpoint traces of a sample are enumerated by their timestamps. The timestamp of its earliest observed trace—$x_{i,0}^{0}$—marks the first time $P_{i}$ was first seen in the wild, and $\mathbf{x}^{0}_{i,(t<h)}$ is the set of all endpoint traces recorded within $h$ hours of this.
An ML-based detection model takes a trace $x_{i}$ of $P_{i}$ as its input and aims to infer $y_{i}$. We split a model $f$ into two parts: an encoder enc and a classifier $g$. The encoder produces a vector embedding $z_{i}=\text{enc}(x_{i})$ of the input trace and $f(x_{i})=g(z_{i})$ is the model’s predicted probability (score) that $P_{i}$ is malware. The predicted label $\hat{y_{i}}$ is $\hat{y_{i}}\!=\!1$ if $f(x_{i})\!\geq\!\tau$, or $\hat{y_{i}}\!=\!0$, otherwise. Here, $\tau$ is a threshold tuned based on the requirements, where a higher $\tau$ results in fewer false positives in exchange for fewer true positives. In Section 3.3, we discuss the tuning of $\tau$ in more detail.
Figure 1. An overview of endpoint malware detection.
3.1. Malware Detection Workflow
Figure 1 provides an overview of the workflow for malware detection at endpoints. We focus on behavior-based detectors deployed in step 3. There are two main phases from the vendor’s perspective: training and testing. In the training phase, the vendor collects a set of samples. These samples are often collected from large-scale repositories, either public, such as MalwareBazaar (bazaar, ), or private (ugarte2019close, ). Most prior work in this domain has followed this practice (jindal2019neurlux, ; miller2016reviewer, ; karbab2019maldy, ; dambra2023decoding, ; kuchler2021does, ). The vendor can rely on a public platform, such as VirusTotal (virustotal, ) to label these samples, or for more recent or advanced malware, can allocate malware analysts for manual labeling (yong2021inside, ). Most commonly, the collected samples are executed in sandboxes that can conveniently produce large-scale datasets of traces for training. In contrast, collecting a large-scale dataset of endpoint traces from hosts requires a telemetry infrastructure, a permissive user agreement, and a large enough user base—conditions that commercial anti-malware vendors can meet but researchers typically cannot. The training phase concludes with training an ML model on the collected traces for detecting malware from runtime behaviors. During the test phase, the model is deployed locally at the hosts. Although we view them as sequential, these phases can be interwoven as vendors continuously update their models on new data (chen2023continuous, ).
In practice, vendors deploy a chain of techniques (avcomparisons, ; vmray, ) to detect if a sample $P$ encountered by a host (e.g., from an e-mail attachment) is malware—step 2 in Figure 1. First the sample is analyzed with static techniques, which operate without executing the sample. This is sufficient for classifying most existing samples (dambra2023decoding, ) either as malware ($P$ is prevented from executing) or benign ($P$ is allowed to execute). If $P$ remains undetected after step 2, it will be executed if the host demands it, e.g., by clicking on the attachment. Critically, our endpoint dataset contains only the traces recorded in the wild after this step, i.e., the traces of undetected samples. The challenges in performing behavior-based detection on such samples are unknown, as prior work has not considered it a criterion for constructing evaluation sets. If $P$ is malware, executing it leads to a real-world compromise—an infection. An endpoint detector is the last line of defense that operates at the host in real time to identify whether $P$ is malware by consuming the trace resulting from this execution—step 3. If $P$ is detected as malware during this step, anti-malware products often apply remedies such as quarantining, deleting files, or alerting (quarantine, ).
Endpoint detectors must classify the sample based on an endpoint trace observed on the end host. It is not feasible to first execute the sample in a sandbox and then classify it based on the sandbox trace, as running in a sandbox (whether in the cloud (cloudsandbox, ) or on-premise (onpremsandbox, )) introduces unacceptable delays. Prior research has only evaluated ML detectors on sandbox traces, but in deployment, the detectors will be applied to endpoint traces, so their actual performance has not been evaluated in the research literature to date. We tackle this issue in this paper.
3.2. Model Training and Evaluation Scenarios
Table 2 summarizes the detection scenarios, each defined along four dimensions, in our focus. The Samples dimensions indicate the step of the workflow training or testing samples originate from: 1—a broad distribution of all programs hosts encounter—or 3—a distribution of programs that remained undetected after 2. The Env. (short for environment) dimensions encode where the traces are collected—from a sandbox or endpoint hosts.
Deployment Scenario Training Phase Testing Phase Samples Env. Samples Env. SB->SB Step 1 Sandbox Step 1 Sandbox SB->EP Step 1 Sandbox Step 3 Endpoint EP->EP Step 3 Endpoint Step 3 Endpoint
Table 2. Scenarios for behavior-based malware detection.
SB->SB is the scenario most prior work in our domain considers: both training and testing samples are collected from sample repositories and corpora, which represent the distribution of all samples seen in the wild, and these samples are executed in a sandbox (karbab2019maldy, ; jindal2019neurlux, ; miller2016reviewer, ; kuchler2021does, ). It is already known (and also corroborated by our results) that detectors excel (90%+ TPR) in this scenario (dambra2023decoding, ).
In SB->EP and EP->EP scenarios, detectors are deployed at 3 in Figure 1, i.e., on the endpoint traces of undetected samples. In SB->EP, the detector is trained only on sandbox-based data (same as SB->SB) but deployed for endpoint detection. In EP->EP, the model is trained on the real-world endpoint traces of undetected samples. This departs from the norm in published research, as most prior detectors have relied on sandbox traces for training. We study this scenario to assess the success of a model explicitly trained for endpoint detection. These scenarios represent a trade-off between practicality (SB->EP is easiest to implement) and expected detection performance (where EP->EP is favored).
3.3. Success Metrics for Endpoint Detection
A conventional detector aims to detect as many malware samples as possible, i.e., to maximize the true-positive rate (TPR), while minimizing the number of benign samples misclassified, i.e., the false-positive rate (FPR). The defender can control the TPR and FPR by tuning the detection threshold $\tau$; a higher $\tau$ implies a lower TPR and FPR. Following prior work (arp2014drebin, ), we evaluate our models regarding their TPR @1% FPR on the test sets (reported as TPR). Additionally, we report the area under the TPR-FPR curve, i.e., the receiver operating characteristic curve, to gauge a model’s overall ability to separate malware and benign classes (reported as AUC). Note that a uniform random predictor, regardless of the class proportions in the test set, would achieve 1% TPR @1% FPR and 50% AUC.
Measuring TPR and FPR in SB->SB is straightforward as there is a one-to-one mapping from samples to traces. However, in endpoint detection (SB->EP and EP->EP), there is a one-to-many mapping as multiple hosts may execute the same sample at different times, and these resulting traces often have high variability (avllazagaj2021malware, ). For example, a few malware samples have over 400 traces, each corresponding to a real-world infection. We present a histogram of the number of traces per sample in Figure 6(a). To obtain a one-to-one mapping, we consider “an average trace” by selecting one random endpoint trace of each sample and measuring the TPR (averaged over 100 runs). This gives us a fair way to compare performance across different scenarios. We refine this metric further based on practical considerations. 67% of traces of an average malware sample (52% for benign) are seen within 24 hours of its first execution at a host. As this ratio drops rapidly after the first day (see Figure 6(b)) and malware samples die quickly within days (yong2021inside, ), we only randomly select from the first-day traces—$\mathbf{x}^{0}_{i,(t<24)}$—unless stated otherwise.
4. Technical Setup
4.1. Datasets
Training Testing Dataset Mal. Ben. Mal. Ben. EP 0.5K 16.5K 0.4K 8.1K EP (Traces) 19.4K 531.6K 9.7K 412.5K SB1 46.4K 16.7K 31.0K 16.4K SB2 34.5K 7.8K 18.1K 7.3K SB1 $\cap$ SB2 9.9K 5.3K 8.9K 5.5K
Table 3. A summary of the samples in our datasets.
Endpoint (EP) Dataset. Provided by Avllazagaj et al. (avllazagaj2021malware, ), we use a dataset of program traces recorded on real Windows hosts of a commercial anti-malware vendor from over 100 countries between January and July 2018. The vendor did not know whether the samples were benign or malicious at the time of execution. The samples were executed by the users, who interacted with them naturally, and the vendor’s behavior-based component recorded the traces in a last-ditch effort to discover unknown threats. We have reprocessed this dataset specifically for our problem and relabeled it by querying VirusTotal (virustotal, ) for more accurate labels. Our processed dataset contains around 1M execution traces from 900 malware samples and 25K benign samples. Each trace includes high-level actions (file, registry, process, and mutex actions) of a sample that is executed until its termination. To our knowledge, this is the only dataset that allows us to evaluate endpoint detection at scale.
Sandbox (SB) Dataset. Studying SB->EP realistically requires collecting a sandbox dataset contemporaneous to our EP dataset. Because the sandbox dataset will serve as the training set and the EP dataset as the test set, we must collect samples seen in or before 2018 to respect causality (pendlebury2019tesseract, ). Moreover, we cannot detonate samples in a sandbox, as it is commonly done (karbab2019maldy, ; jindal2019neurlux, ; miller2016reviewer, ; kuchler2021does, ), as behaviors of old programs today would differ from their original behaviors. This is because malware stops functioning when, for example, its remote infrastructure dies (yong2021inside, ) or it starts behaving differently over time (avllazagaj2021malware, ). Our solution is collecting traces from VirusTotal (virustotal, ), where third-party sandbox vendors publicly share behavior reports on samples. To this end, we curate a list of SHA-256 hashes of Windows samples from popular public malware detection corpora, including EMBER (anderson2018ember, ), and SOREL (harang2020sorel20m, ), released in 2017 or 2018. We then query VirusTotal with these hashes to collect their sandbox traces when available. We discard traces that came more than six months after the sample was first seen to capture close-to-original behaviors. This process results in traces from two sandbox vendors for around 110K malware and 40K benign samples. These vendors are well-known in their countries of origin (SB1 is from China, and SB2 is from the USA). There is a class imbalance as vendors on VirusTotal are more inclined to share traces of malware samples.
Rank Test Set Trained on SB1 Trained on SB2 Trained on EP SB1->SB1 SB1->SB2 SB1->EP SB2->SB1 SB2->SB2 SB2->EP EP->EP NGR HYB ATT NGR HYB ATT NGR HYB ATT NGR HYB ATT NGR HYB ATT NGR HYB ATT NGR HYB ATT SB1 95.0 94.2 93.2 60.7 53.7 56.7 11.2 8.0 4.5 48.5 31.8 32.7 85.5 59.2 83.7 14.2 7.2 12.3 43.5 34.0 25.7 SB2 94.0 93.1 92.2 70.4 61.2 64.3 13.9 7.1 5.5 17.6 10.0 14.7 92.6 91.4 90.8 7.5 6.3 11.4 43.4 37.6 26.4 EP 93.2 93.1 87.8 63.5 48.0 52.9 16.7 10.2 8.3 34.5 16.6 15.7 79.3 79.5 88.7 17.3 10.9 13.1 49.5 43.5 42.8
Table 4. The performance (TPR%) of three ML approaches (NGR, HYB, ATT) in seven detection scenarios (based on Table 2). In each row, the models are ranked based on the test set, and the average TPR of top-20 models is reported (see Table 15 for AUC).
Trace Standardization. Our EP traces contain high-level information that the vendor can seamlessly collect and analyze at endpoint hosts, such as file, process, and mutex creations, as well as registry key creations and deletions. Sandboxes also record lower-level information (such as memory dumps (sandboxmemory, )) as they are less computationally constrained. Moreover, our traces from three sources (SB1, SB2, and EP) have different formats and conventions. This prompts us to implement routines to keep, in our sandbox traces, only the type of information available in EP traces and covert all traces to a standardized format (see Appendix A for details). As shown in Section 5, the information remaining in sandbox traces is still enough to obtain performance comparable to prior work ($\sim$95TPR and 99% AUC) in the SB->SB setting. This standardization is also critical for our leaderboard design (Section 8).
Temporal Splitting and Labeling. We split our datasets into two portions based on the timestamps of the samples (the first-seen dates) to ensure that the train and test samples are temporally disjoint, avoiding a common pitfall in prior work (pendlebury2019tesseract, ; arp2020and, ). Samples seen before April 1st, 2018, along with their traces, are in the Training portion, and the samples seen after that date are in the Testing portion, on which we never train. As labels of samples are known to change mildly over time (labeldynamics, ), we make the best effort to assign the training samples historical labels that were available before April 1st, 2018. We could find such historical labels for $\sim$24and 100% of the samples in the EP and SB training portions, respectively. For testing samples, we query VirusTotal to obtain the most recent detection reports and label samples detected by over five anti-malware engines as malware, following the advice in (labeldynamics, ). Further, we use AVClass2 (sebastian2020avclass2, ) to assign family labels to our malware samples based on their detection reports. We tag malware samples that did not receive a family label as Generic malware. Overall, this methodology ensures that our evaluation is realistic and reflects the conditions the anti-malware vendor had to work under when our endpoint data was collected.
4.2. Machine Learning Details
We experiment with three ML approaches: (i) a bag-of-n-grams-based model (NGR), (ii) a hybrid model that combines convolution and attention (HYB), and (iii) a self-attention-based sequence model (ATT). These three approaches cover a wide range of designs proposed by prior work in behavior-based malware detection with ML, respectively, MalDy (karbab2019maldy, ), Neurlux (jindal2019neurlux, ) and, most recently, Nebula (trizna2023nebula, ). Please find the details on these models in Appendix B. We address the class imbalance in our training sets by oversampling the underrepresented classes (malware in EP and benign in SB datasets) (chawla2002smote, ). Note that our goal is not to develop a better model architecture or feature processing routine but to evaluate the effectiveness of state-of-the-art architectures at endpoint malware detection. We make our ML improvements without changing the fundamentals of these existing approaches.
[…]
5. Endpoint Detection Evaluations
Table 4 presents the endpoint detection performance of ML models (based on the methodology in Section 4.2), using three different data sets (SB1, SB2, and EP) for training and testing, following the scenarios in Table 2. Here, for example, SB1->SB2 and SB1->EP refer to the scenarios where the models are trained on the SB1 training set and evaluated on the SB2 and EP testing sets, respectively.
5.1. Evaluating the SB->EP Scenario
First, we see significant gaps between SB->SB and SB->EP scenarios across the board. This gap is the largest when models are trained and ranked on the same sandbox, e.g., 95.0% vs. 11.2% for NGR trained and ranked on SB1. When models are ranked on a different sandbox, it often shrinks, e.g., 94.0% vs. 13.9% for NGR trained on SB1 and ranked on SB2. Nevertheless, the gap is still substantial even when we rank the models on their EP performance, e.g., 93.2% vs. 16.7% for NGR trained on SB1 and ranked on EP. Although there is a gap between the performances on different sandboxes, e.g., 94.0% vs. 70.4% for NGR trained on SB1 and ranked on SB2, it is significantly smaller than the SB->EP gap. Overall, this gap persists regardless of the training sandbox or the model type, hinting at the limitations of sandbox-trained models for endpoint detection. Notably, this gap is larger for models that perform better on the test traces from a sandbox that also produced the training set. In consequence, studies that rely on a single sandbox for evaluation in SB->SB (miller2016reviewer, ; karbab2019maldy, ; kuchler2021does, ; trizna2023nebula, ) are at a higher risk of producing worse models for SB->EP.
Trained on SB1 Trained on SB2 SB1-SB2 SB1-EP SB2-EP SB1-SB2 SB1-EP SB2-EP NGR 0.47 0.03 0.45 --0.24 0.39 --0.46 HYB 0.88 0.84 0.75 --0.17 0.59 0.33 ATT 0.56 --0.17 0.31 0.51 0.58 0.35
Table 5. Ranking correlations of sandbox-trained models according to their performances (TPR) on different test sets.
Next, we verify the possibility of finding a better model for SB->EP by ranking the models using traces from an unseen sandbox (not used for training). We rank our models (100+ in each setting) based on their SB1, SB2, and EP performances (TPR) and compute Spearman’s correlation coefficients between these rankings. Table 5 reveals that (except for one setting) (i) the rankings based on the training sandbox correlate poorly (sometimes negatively) with EP rankings, (ii) rankings based on an unseen sandbox correlate more strongly with EP rankings. For example, for ATT trained on SB1, the rankings based on SB1 and SB2 have $0.03$ and $0.45$ correlation with the rankings based on EP, respectively. We believe this is because a model that performs better on the traces from an unseen sandbox is less likely to have overfitted to features specific to the training sandbox. We discuss such features in more detail in Section 7. As a result, to get higher performance in SB->EP, we recommend performing model selection using traces from a different, unseen sandbox.
Finally, we experiment with combining traces from two sandboxes (SB1 and SB2) for training. However, this brings only minor improvements for SB->EP over training on traces from a single sandbox. When we use EP traces to rank the models, training on both sandboxes achieves at most 17.5% TPR on the EP test set for NGR (11.7% for HYB, and 13.9% for ATT). In Section 7, we show that the diversity between the traces of a sample from two sandboxes is much lower than between its traces from two endpoint hosts. This lack of diversity makes combining SB1 and SB2 traces ineffective for increasing the EP performance.
Applying Drift Detection. We study the low performance in SB->EP through the lens of distribution shift, which occurs if the process generating inputs, i.e., program traces, changes between the training and testing phases. Concept drift is a highly-studied type of distribution shift in malware detection that stems from malware evolution (pendlebury2019tesseract, ). In past work on malware concept drift, the distribution shift builds up over time in concept drift, e.g., over 50% performance drop in 2 years (barbero2022transcending, ). In contrast, we find that in SB->EP the shift manifests immediately after the model is deployed.
One way to combat malware concept drift is to use a drift detector (DD) to identify samples that might have drifted and abstain from classifying them (jordaney2017transcend, ; barbero2022transcending, ; yang2021cade, ; chen2023continuous, ). DDs aim to prevent mistakes by rejecting out-of-distribution (OOD) samples that diverge from the training distribution while accepting in-distribution (ID) samples on which predictions are reliable (jordaney2017transcend, ; yang2021cade, ). We apply drift detection to quantify the distribution shift present in SB->EP. We employ a modern technique ReAct (sun2021react, ) to assign an outlierness score to each sample, among which the highest-scoring ones are rejected. ReAct rectifies a model’s penultimate layer activations, which is shown to make the prediction probabilities of a model a more calibrated indicator of outlierness. We selected ReAact due to its flexible approach that is agnostic to the underlying model as long as it’s a neural network.
Figure 2. The rejection rates resulting from drift detection.
[…]
Finally, in Table 6, we present the TPR on the accepted (remaining) traces in EP and SB2 for increasing values of $K$. Rejecting the traces in EP or SB2 increases the performance, but this increase is less pronounced for EP, e.g., gaining $\sim$10TPR (over the baselines in Table 4) takes rejecting 67% and 22% of malware traces in EP and SB2, respectively. This also distinguishes the distribution shift in SB->EP from concept drift; rejection rates must be high for meaningful performance gains due to the magnitude of the shift.
($K$) ID RejSB1->EP SB1->SB2 TPR MalBenTPR MalBen5% 20.2% 31.1% 20.5% 79.2% 22.0% 21.5% 10% 23.5% 53.2% 39.0% 89.4% 38.2% 36.5% 15% 25.6% 67.7% 46.4% 88.3% 52.4% 53.6% 20% 42.6% 72.2% 49.0% 87.7% 60.8% 60.6% 25% 53.1% 78.5% 52.5% 82.9% 68.5% 70.8%
Table 6. The TPR of an NGR model in SB1->SB2 and SB1->EP scenarios, with increasing ID rejection rates, and the corresponding rejection rates of malware and benign traces.
5.2. Evaluating the EP->EP Scenario
For our models in EP->EP, we select up to 4 EP traces per sample to form our training set. As we will show in Section 7, the behavior variability among the real-world traces of a sample (avllazagaj2021malware, ) impacts the model’s predictions. Consequently, including more traces per sample for training provides better coverage of the behaviors in the wild. We find that using more than 4 traces per sample provides diminishing returns, e.g., for 1, 2, 4, and 16 traces per sample, the TPR of NGR is 43.0%, 44.8%, 49.5%, and 49.9%, respectively.
In Table 4, we see a much higher EP performance in EP->EP than in SB->EP, as expected. Although training and testing sets are from similar distributions in EP->EP (as opposed to SB->EP), the performance still lags behind the SB->SB performance, e.g., 95% vs. 50% TPR and 99% vs. 88% AUC (Table 15). In the following sections, we shed light on this discrepancy and what makes endpoint detection particularly more challenging than a sandbox-only scenario.
6. Endpoint Detection Deals With Difficult-To-Classify Samples
As described in Section 3.1, an endpoint detector operates on samples that cannot be classified statically—a subset of all samples that exist in the wild. The filtering effect this introduces on the distribution of samples relevant to endpoint detection has not been systematically measured before. To illustrate, we explore (based on Table 13) the top malware families in our EP test set (undetected samples) and the SB test set (samples from public corpora) Although both datasets include only the samples first seen in 2017-2018, they encode vastly different priors over samples. For example, generic malware (i.e., malware that could not be placed into any known family) covers 23.8% of the EP samples vs. 4.8% of the SB samples. In the EP set, there are families such as Khalesi and Emotet that were rampant in 2018 (emotet, ). Conversely, in the SB set, we see older families such as Sivis and Upatre (circa 2013-2014), whose variants still spread to this day. Moreover, in Table 13, we share top benign sample publishers (extracted from the code-signing certificates when available), indicating a similar distribution difference in benign samples. Overall, the typical distribution of training and testing samples used by prior work in SB->SB (karbab2019maldy, ; jindal2019neurlux, ; dambra2023decoding, ) is substantially different than the distribution of testing samples in SB->EP and EP->EP.
Borderline Samples and Label Noise. A common practice in related work in malware detection with ML is discarding borderline samples on which the anti-malware engines deployed by VirusTotal are not in consensus. Different works use different criteria; for example, they discard a sample if it is detected by less than 20 engines (mantovani2020prevalence, ) or five engines (kuchler2021does, ), or by more than 0 and less than 40 engines (lucasadversarial, ). Even for works that do not explicitly discard borderline samples, their data collection results in few of them. For example, recent related work has suggested that dynamic features have limited benefits over static features (dambra2023decoding, ). The benign samples used in this work were from a trusted Windows software repository, and 93% of their malware samples were detected by 20+ engines.
Based on prior definitions, we consider samples detected by more than 0 and less than 20 engines borderline. This results in 27.7% borderline samples in our EP test set (vs. 15.9% in the SB test set). The prior practice would have discarded over a quarter of the samples that a real-world endpoint detector encountered. Next, we will show that this artificially inflates the measured success of a model. As described in Section 4.1, we do not discard any samples and label a sample as malware if more than five engines detect it. This threshold poses a trade-off for the ground truth: as it increases, more malware samples may be labeled as benign, and as it decreases, more benign samples may be labeled as malware. We present a case study in Section 6.2 to demonstrate how this trade-off plays out.
6.1. The Impact of Sample Distributions
Here, we measure how the distribution of samples in the test set changes a model’s measured performance. To this end, we resample our test sets according to some criteria, e.g., following the malware family distribution in the EP test set, and evaluate our models on these new test sets. This simulates different distributions over a model’s test samples and allows us to do controlled experiments. We focus on NGR models due to their superior EP performance in Section 5.
EX# SB1->SB1 SB1->EP EP->EP
TPR AUC TPR AUC TPR AUC
EX0 No Resampling - Orig. Distributions
93.2 99.0 16.7 78.4 49.5 87.5
EX1 Orig. Distributions w/o Generic Malware
95.6 99.5 18.6 81.4 56.8 91.4
EX2 Only Generic Malware
38.4 85.8 9.8 67.8 23.9 74.2
EX3 Discard Borderline Samples
96.7 99.5 22.6 83.9 66.9 94.2
EX4 Malw. Resampled Following EP Test (w/ Generic)
63.2 94.2 16.7 78.4 49.5 87.5
EX5 Malw. Resampled Following EP Test (w/o Generic)
72.1 97.2 18.6 81.4 56.8 91.4
Table 7. The impact of sample distributions in the test set on model evaluations. Each experiment is denoted as EX0-5.
Table 7 presents the results of our resampling experiments. We resample the test sets according to five criteria—denoted as EX0--5 where EX0 is testing on the original test set of each respective scenario. In EX1, leaving out generic malware samples from original test sets leads to a significant boost in performance, whereas keeping only generic malware in EX2 causes a massive drop. This aligns with prior claims that generic malware is at the borderline and much harder to classify (ugarte2019close, ). The fact that Generic has 5$\sim$ 5$\times$ 5 × higher coverage in the EP test set than in the SB test set (23.8% vs. 4.8%) also demonstrates the filtering effect that funnels more difficult samples to the endpoint detector. Next, in EX3, we discard the borderline samples (identified in the previous section) from all test sets, simulating a common prior practice (kuchler2021does, ; mantovani2020prevalence, ; dambra2023decoding, ). This yields a significant boost in all scenarios: 5.9% TPR boost in SB1->EP and 17.4% in EP->EP. In EX4 and EX5, we resample the test sets to match the distribution of malware families in the EP test set, with and without generic malware, respectively. Note that, for SB1->EP and EP->EP, EX4 is equivalent to EX0 and EX5 is equivalent to EX1. We observe a great decrease in the original SB1->SB1 performance, e.g., from 93.2% TPR to 63.2% in EX4 and to 72.1% in EX5. These results suggest that the difference in distribution of malware families explains some of the performance drop when applying these classifiers to EP traces, perhaps because EP traces contain more harder-to-classify malware families, but there remains a large gap that is not explained by differences in the distribution of families. Despite neither including any generic malware, the SB1->SB1 performance in EX5 is significantly lower than EX1. This suggests that non-generic malware samples relevant to endpoint detection are generally more difficult to classify correctly, whether based on their sandbox or endpoint traces. Overall, these experiments reveal how the performance of prior behavior-based detectors may be over-estimated because they are evaluated on samples that are not representative of the distribution of samples faced in endpoint detection.
6.2. Case Studies
Following our measurements, we present two case studies illustrating the sample distribution challenges in endpoint detection.
Difficult Benign Samples. Table 13 shows that Microsoft is the top benign publisher in our EP test set, covering 7.8% of all benign samples. This is counterintuitive as Microsoft is a trusted publisher, and our EP dataset only records samples that could not be classified statically as malware or benign. We discovered that almost all (98%) Microsoft samples in our EP test set share the AM_Delta prefix in their filenames, corresponding to periodic patches to Windows Defender Anti-Malware. The NGR model in EP->EP outputs an average score of 37% on AM_Delta endpoint traces, almost 2.5$\times$ of the average score across all benign traces. Note that the score quantifies the model’s confidence that the input trace belongs to a malware sample. In EP->EP, on a test set containing only AM_Delta samples as the benign samples, the model achieves only 62.6% AUC, compared to 87.5% with all benign samples. Moreover, there are online reports (amdeltareport, ) about anti-malware software falsely flagging AM_Delta files. A popular configuration repository for Sysmon—a tool to monitor process activities—includes an allow-list rule specifically for AM_Delta (amdelta_syzmon, ). Our evidence suggests that AM_Delta samples are difficult to classify correctly due to making sensitive modifications on hosts. Conversely, our SB test set contains only two AM_Delta samples (less than 0.01% of all benign samples). This highlights that an endpoint detector often faces difficult—false positive prone—benign samples in addition to difficult malware families, which is a root cause behind lower performance in SB->EP and EP->EP.
Label Noise in Borderline Samples. Roblox is a popular game creation platform. Our EP train and test sets contain 19 and 29 samples from Roblox. Despite being from a trusted publisher, there are reports about anti-malware products flagging Roblox samples. Six (31%) of our 19 training samples were labeled as malware because they were detected by over five engines on VirusTotal, whereas none of the 29 test samples were labeled as malware. This causes a model to associate Roblox samples with malware-ness during training, potentially introducing false positives during testing. For example, our NGR model in EP->EP outputs an average score of 40% on the endpoint test traces from Roblox samples. This demonstrates how borderline samples such as Roblox can introduce label noise, particularly in endpoint detection, where they are more common.
Before After NGR HYB NGR HYB TPR AUC TPR AUC TPR AUC TPR AUC Soft Labeling (EP->EP) 49.5 87.5 43.5 86.8 50.6 87.8 49.2 88.2 Resampled Training Set — Uniform (SB1->EP) 16.7 78.4 10.2 74.1 15.9 77.2 8.6 74.0 Resampled Training Set — EP Training (SB1->EP) 16.7 78.4 10.2 74.1 20.0 78.6 11.4 74.9
Table 8. Our strategies to improve the performance against sample distribution challenges in endpoint detection.
6.3. Combating Sample Distribution Challenges
Building an intuition from our previous experiments, we propose two strategies to improve the performance in SB->EP and EP->EP.
Soft-Labeling Against Label Noise in EP->EP. Malware detectors are generally trained using hard binary labels—0 for benign and 1 for malware. However, hard labels are hazardous when noisy as they force the model to overfit to a wrong prediction (li2019learning, ). The high frequency of borderline, i.e., potentially noisy, samples exacerbates this problem in endpoint detection. Research suggests that soft labels can be effective against label noise (lukasik2020does, ) by preventing the model from getting too confident on noisy labels during training. Hoping to improve the EP->EP performance, we implement a function to assign soft labels to our borderline training samples.
[…]
More Accurate Training Distributions in SB->EP. Our models in SB->EP were trained on the distribution of samples captured in public corpora. Although this is standard for obtaining the best SB->SB performance (jindal2019neurlux, ; trizna2023nebula, ), it is not ideal for SB->EP where the test set follows a different distribution. ML models are heavily influenced by the priors encoded in their training sets due to the empirical risk minimization (ERM) paradigm, which minimizes the average loss over all training samples. For example, 10% of the malware in the SB training set is from the Virut family, and, therefore, that family will be prioritized over Emotet, which accounts for 0.3%.
We first attempt to alleviate this discrepancy in SB->EP by resampling the SB training set to include an equal number of samples from each family. Unfortunately, as presented in Table 8 (second segment), this uniform sampling strategy ends up hurting the performance in SB->EP. Considering over 700 malware families are in the SB training set (vs. around 100 in the EP training set), families relevant to endpoint detection remain a minority in the new family-balanced training set. This explains the low performance and raises doubts about whether such an uninformed strategy could work.
Finally, we experiment with an informed strategy where the SB->EP training set is resampled to follow the family distribution in the EP training set. Widely used threat intelligence platforms can guide practitioners about which families to include in training (ciscothreatintel, ). Table 8 (third segment) shows that this strategy brings moderate performance improvements: 3.3% and 1.2% higher TPR for NGR and HYB, respectively. We will continue using this strategy to train models in the rest of our experiments. We leave further improvements to future work, such as handling notably noisy family labels (sebastian2020avclass2, ; kotzias2019mind, ).
7. Environment-Sensitive Program Behaviors Hurt Endpoint Detection
We have shown the performance implications of the distribution of samples encountered by endpoint detectors. However, even when we account for distribution differences (EX4 in Table 7), the performance in SB->SB (63.2% TPR) is higher than the performance in SB->EP (16.7% TPR) and in EP->EP (49.5% TPR). In this section, we aim to illuminate the remaining performance gap between by studying the impact of variable program behaviors.
It is known that program behaviors are environment sensitive and, as a result, a sample can exhibit varying behaviors in different environments (balzarotti2010efficient, ; lindorfer2011detecting, ; avllazagaj2021malware, ). Moreover, malware samples are more environment-sensitive than benign samples because, for example, they execute only when specific vulnerabilities are present in the host (avllazagaj2021malware, ) or do not execute if they detect they are running in a sandbox (lindorfer2011detecting, ; miramirkhani2017spotless, ). In this section, we build on these insights to understand the impact of variable program behaviors on endpoint detection performance. We focus on the variability stemming from environment differences between sandboxes and endpoint hosts (for the SB->EP scenario) and between different endpoint hosts (for the EP->EP scenario). We perform our experiments using the improved NGR models from Section 6.3.
7.1. The Diversity Among Endpoint Traces
Table 9 shows how much traces vary when similar sets of samples (e.g., from the same family) are run in different environments. We aggregate all the traces from three popular malware families and two benign publishers from our SB1, SB2, and EP datasets. To cover different types of samples, we select Wannacry (ransomware), Emotet (banking trojan), and Khalesi (info-stealer) as the families, and Opera (web browser), Roblox (game platform) as the publishers. We then measure the average pairwise distances between these traces within an environment, e.g., between the traces from SB1 (SB1-SB1), and between two environments, e.g., between the traces from SB1 and EP (SB1-EP). We use normalized compression distance metric, which has been used in similar contexts (bailey2007automated, ; avllazagaj2021malware, ).
Within Environment Between Environments SB1-SB1 SB2-SB2 EP-EP SB1-SB2 SB1-EP SB2-EP Wcry 0.16 0.11 0.44 0.28 0.67 0.50 Emot 0.48 0.35 0.50 0.55 0.65 0.53 Khal 0.33 0.29 0.48 0.52 0.62 0.57 Oper 0.05 0.07 0.69 0.43 0.70 0.75 Rblx 0.18 0.04 0.63 0.38 0.67 0.69
Table 9. Average pairwise distances between traces of samples from three malware families and two benign publishers.
We observe: (i) traces from the same sandbox are almost always very similar; (ii) traces from two endpoint hosts are dissimilar (also shown in (avllazagaj2021malware, )); (iii) traces from different sandboxes are dissimilar but much less so than one trace from a sandbox and one trace from an endpoint host. The observation (iii) is potentially detrimental for SB->EP, as the features learned from SB traces might be irrelevant to EP traces, and (ii) is potentially detrimental for both scenarios. The dissimilarity between SB and EP traces implies that the training set of a model in SB->EP follows a different distribution than its test set, which explains the high drift detection in Section 5.1.
Next, we demonstrate how (iii)—the distance between SB traces and EP traces—affects a model in SB1->EP. Figure 3 shows that the model is more confident for traces that are more similar to SB traces. Here, each marker represents a group of 30 traces of a malware family from endpoint hosts, and the x-axis quantifies the average distance of these traces to all traces of this family from SB1. The lower this distance, the more similar a group of endpoint traces is to traces of this family from SB1. For example, the model outputs a 90% score for Emotet EP traces with an average distance of 0.60 to Emotet SB1 traces, which drops to 60% for an EP trace with a distance of 0.75. Consequently, the less similar the input EP trace is to the SB traces in the training set, the worse decisions the model makes and performs poorly in SB->EP.
Figure 3. Relationship between the average distance of an input EP trace to the SB traces and the model’s output score in SB->EP. Each marker represents the avg. of 30 input traces.
[…]
Figure 4. Prediction score std. deviations on the endpoint traces of each sample in the EP test set; in SB1->EP and EP->EP.
7.2. Case Studies
After our quantitative results on behavior variability and its impact on the model, we present two case studies demonstrating its nature.
Behavior Variability in Benign Samples. RobloxLauncher and OperaPatcher are two benign samples on the opposite ends of the variability spectrum. RobloxLauncher accepts user input through a graphical interface, such as mouse clicks on menu items. We clustered the EP traces of this sample to discover two main execution paths: (i) downloading a long list of assets (such as .mp3 or .jpeg files), presumably to update the game; (ii) creating temporary Internet files and starting a process, presumably for launching the game. Our EP->EP model makes more accurate predictions on the traces in (i) than (ii) (26% vs 38% average prediction score). In its SB2 trace, the sample downloads a similar list of assets, whereas, in its SB1 trace, it creates only a log file and stops, likely because it could not access the Internet. Neither sandbox triggers the execution of the game launch code path, so we expect that a model trained on sandbox traces would struggle to classify endpoint traces that launch the game. Applied to a model trained on SB1 traces, a drift detector (Section 5.1) consistently rejects the endpoint traces of this sample. We believe this illustrates a broader challenge: emulating user interactions in sandboxes is challenging (liu2022enhancing, ), causing major differences in the data distribution of sandbox vs endpoint traces from benign samples. On the other hand, OperaPatcher performs almost the same actions in all its SB1, SB2, and EP traces: creating localization and library files to update the browser. The fact that this sample executes without any user input eliminates behavior (and prediction) variability across traces.
Non-Malicious Malware Traces. In behavior-based detection, most commonly, the label of a sample (e.g., obtained from VirusTotal) is transferred to all traces from this sample, which are then used for training and testing. This practice implicitly assumes that all traces of a malware sample contain some discernible malicious activity. However, this can be problematic in cases where the sample has failed to execute, e.g., because its remote infrastructure is down (yong2021inside, ).
To assess this approach, we study the Wannacry ransomware family. We select Wannacry as it is thoroughly dissected (unlike most families), and its indicators-of-compromise (IOCs) are well-known, allowing us to gauge whether a particular trace is associated with compromise. Using multiple sources, we create a list of IOCs for Wannacry and look for them in our SB1, SB2, and EP traces. We separate the traces into two sets based on whether they contain any IOC and then measure the average prediction scores of our models on each set. Based on Table 14, we observe (i) most traces contain at least one IOC (e.g., 97.7% of the EP traces), (ii) the model’s average prediction score is lower on the traces with no IOC (e.g., 97% vs. 52% in EP->EP); however, it is still significantly higher than the averages on benign traces (13%). The ability to associate malware traces with no IOC with maliciousness is a strength of ML-based methods over IOC-based detectors. Any malware that reaches the execution stage is a potential danger, regardless of whether it successfully compromises the host.
7.3. Sandbox Features That Do Not Generalize
We have shown that sandbox and endpoint traces for a given sample are different. In this section, we examine why they are different and how this difference might hurt the model in the SB->EP scenario.
Sandbox-Specific Features. We found several sandbox-specific features that are highly prevalent (seen in many traces) and predictive (seen only in malware traces), but occur only in one sandbox and not in the other environments. ML models typically exploit such features to minimize the loss and may overlook other predictive features (pezeshki2021gradient, ), causing the model to perform poorly on traces collected from a different sandbox or from endpoints. Table 10 presents four of these features (called artifacts) that we have found in SB1 and SB2 traces. We have identified over 100 such artifacts, most of which frequently co-occur. For each artifact, we report its prevalence (Prv.) and its malware ratio (MalR.) in SB1, SB2, and EP training sets. Prv. is the percentage of samples in which the artifact is present, and MalR. is the percentage of these samples labeled as malware.
File Name SB1 Traces SB2 Traces EP Traces Prv. MalR. Prv. MalR. Prv. MalR. SogouExp. 9.0% 100.0% 0.2% 0.0% PersonalBankP. 3.3% 100.0% Spotify 0.1% 0.0% 3.0% 100% 0.2% 0.0% Python 3.1% 100% 0.1% 0.0%
Table 10. Some strong malware features found in specific sandboxes that do not generalize to other environments.
These artifacts appear to be a result of the specific configuration and particular apps pre-installed on the specific sandbox (see Appendix C). As real-world hosts and other sandboxes do not share the same configuration, this explains why an endpoint classifier trained on traces from one sandbox might fail to generalize to endpoint traces seen in the wild.
The Impact of Sandbox Evasion. Recent works have measured that 40-80% of malware uses at least one evasive technique (galloro2022systematical, ) to avoid analysis or terminate early if it is running in a sandbox (mcafeeevasion, ; miramirkhani2017spotless, ). Evasion makes sandbox traces dissimilar to endpoint traces. Although sandbox evasion is well understood, its implications for endpoint detection have not been measured.
We explore the impact of sandbox evasion. We use a standard heuristic and treat a sample as evasive if the number of actions it performs is too low (kuchler2021does, ; balzarotti2010efficient, ; miramirkhani2017spotless, ). We first find the malware families common between our EP and SB1 test sets (30 total). We then compute the average trace length of each common family using the traces of the samples belonging to it. We count only registry actions, as they can be recorded unambiguously, unlike actions such as process injections, which might have vendor-specific definitions. We then find the length differences between SB1 and EP traces of each family. For example, Wannacry and Gandcrypt have differences of $+11$ and $-6$, respectively. We then split the families into two sets: the 21 families whose SB traces are longer (e.g., Wannacry) are deemed less likely to be evasive, and the 9 families whose SB traces are shorter (e.g., Gandcrypt) are deemed more likely. The median length differences for the non-evasive and evasive families are $+4.4$ and $-4.3$, respectively, with a few outliers on both sides, e.g., $+81$ for Vobfus and $-85$ for Bypassuac. Finally, we measure the TPR of our model in SB1->SB1 and SB1->EP on these two malware sets individually while keeping the benign samples the same.
Families #Fams #EP #SB1 SB1->EP SB1->SB1 All 30 169 5.8K 23.1% 86.4% Evasive 9 49 2.9K 20.4% 89.8% Non-Eva. 21 120 2.9K 23.3% 83.0%
Table 11. The performance (TPR) on malware families more (Evasive) or less (Non-Eva) likely to be evading sandboxes.
The trace length of a sample monotonically increases as its execution continues. This means our measurements might be confounded by the discrepancies between the execution durations in SB1 and endpoint hosts. Before presenting our results, we confirm this is unlikely to be the case: 82% of our EP traces are from executions that lasted less than 90 seconds. Although we do not know the exact configuration of SB1, allowing one to two minutes of execution is standard for most sandboxes in practice (kuchler2021does, ).
Based on the results in Table 11, we observe: (i) SB1->EP performance is higher (by 3%) on non-evasive families than on evasive families; (ii) SB1->SB1 performance is significantly higher (by 7%) on evasive families than on non-evasive families. This suggests that the classifier exploits short traces (evasiveness) as a feature, which is beneficial for SB->SB, though it does not generalize to SB->EP. However, the non-trivial SB1->EP performance on evasive families might hint that there are still features useful for endpoint detection in the sandbox traces of evasive samples. Although we are limited to observational data, a controlled study on evasive malware to disentangle these features is a promising direction.
The Trace Length Bias. Building on the previous experiment, we hypothesize that a model in SB->EP might learn an inverse correlation between trace length and malware-ness, i.e., evasive malware creates short traces, and, therefore, short traces are more likely to be malware. In Figure 8, we present the model’s average predicted scores on traces with a certain length in SB1 and EP test sets. For both SB and EP test sets, we also measure the ratio of malware traces of a given length labeled among all traces of that length. For example, if there are 100 total traces of length $L$ and 70 of them are labeled as malware in the ground truth, the malware ratio would be 0.7 for the traces of length $L$.
These plots support our hypothesis: the model predicts a higher score for the shortest traces in both test sets, i.e., it has a length bias. This bias leads to accurate predictions in the SB set, where very short traces are much more likely to be malware (the malware ratio on the leftmost side of the upper plot is high). However, this bias leads to inaccurate predictions in the EP set, where very short are not more likely to be malware. On the EP traces shorter than 10 actions ($\sim$20of all EP traces), the model achieves 15% TPR (vs. 20% when all traces are kept). Ultimately, trace length is a spurious correlation learned from sandbox traces that fails to generalize to endpoint traces. Evidence suggests this is introduced by evasive malware that tends to produce short sandbox traces in the same execution time. Methods preventing the model from learning such correlations (kirichenko2023last, ; pezeshki2021gradient, ) offer a promising next step.
7.4. Combatting Environment-Sensitivity With Environment-Invariant Features
We have shown that the environment-sensitivity of program behaviors impacts the model. When the classifier exploit features that are specific to the training environment (e.g., trace length), its predictions fail to generalize. Table 10 suggests that the artifacts of SB1 are absent in SB2, and vice versa, as they are configured sufficiently differently. This highlights an opportunity: learning features that are invariant across two environments (e.g., both sandboxes) might yield better generalization to endpoint detection.
Prior work has studied invariant learning in self-supervised learning, e.g., viewpoint-invariant visual features (purushwalkam2020demystifying, ). We employ the Siamese loss (chen2021exploring, ), which forces the model to produce similar embeddings (measured by a distance metric) on pairs of semantically related inputs. In our case, each pair consists of two traces of the same sample from different sandboxes (in SB->EP) or from different EP hosts (in EP->EP). To gain robustness to variable behaviors, we aim to make the model invariant to the differences in these pairs (such as sandbox-specific artifacts). The critical question is whether invariance across SB traces translates to invariance across EP traces. We represent a pair of traces of a sample $P_{i}$ as ($x_{i,0},x_{i,1})$ and the set of all pairs as $\mathbf{D}$; and define the following loss function that is added to the standard ERM loss as a regularizer:
$\mathcal{L}_{\text{inv}}=\dfrac{1}{\lvert\mathbf{D}\rvert}\sum_{(x_{i,0},x_{i,1})\in\mathbf{D}}-\mathrm{cos}(\text{enc}(x_{i,0}),\text{enc}(x_{i,1}))$
$\mathcal{L}_{\text{final}}=\mathcal{L}_{\text{ERM}}+\alpha\mathcal{L}_{\text{inv}}$
[…]
Table 12. The impact of $\mathcal{L}_{\text{inv}}$ on endpoint detection TPR.
[…]
Figure 5. The model’s embeddings with increasing values of $\alpha$ for invariant learning. Sep quantifies the separability of the SB1 and SB2 embeddings. t-SNE (van2008visualizing, ) visualization.
[…]
8. Endpoint Malware Detection Leaderboard
Our investigation has relied on a dataset of endpoint traces from the wild provided to us by Avllazagaj et al., who collected and analyzed it in their prior work (avllazagaj2021malware, ). We have pinpointed the salient challenges in ML-based endpoint malware detection using this data. Performing evaluations only on lab-based data can obscure these challenges and produce biased solutions inapplicable to the real world (cherubin2022online, ; jacobs2022ai, ). To minimize this bias, we designed a leaderboard for realistic evaluations of behavior-based malware detectors (malwaredetectioninthewild.github.io (leaderboard, )). Prior leaderboards such as RobustBench (croce2021robustbench, ) or TrojAI (trojai, ) have motivated research into critical and difficult problems and fueled many innovations in ML.
Anyone may submit a pre-trained detector and a feature extractor that converts a trace in the standardized format into an input to the detector. We evaluate each submitted detector on our endpoint traces and list it on our leaderboard. Although we are unable to share Avllazagaj et al.’s dataset, we will release the sandbox dataset we collected and the metadata for over 200K samples used in our work to assist the participants. The details of these artifacts can be found in §E. This format offers a blueprint for security vendors to guide researchers without publicly sharing their sensitive data. We hope to spearhead it into more applications of ML for security.
9. Discussion
Threats to Validity and Limitations. Our study has some limitations due to the nature of our endpoint data. First, this data was collected at the hosts of a single security vendor. Although we cannot rule out selection bias, the fact that these hosts are located in over 100 countries in both enterprise and consumer settings suggests that our results have broad applicability. Second, our data was collected five years ago. This is a common limitation in malware detection studies as collecting large-scale up-to-date data has not been feasible for researchers (yang2021cade, ; barbero2022transcending, ; chen2023continuous, ). We focus on the gaps between sandbox-based and endpoint detection and not on the specifics of the threats that existed at the time, making our observations relevant to today’s malware detection landscape. Third, our data consists of only Windows hosts and lacks any network-related actions. As 95% of malware is aimed at Windows (windowsmalware, ) and ML-based detectors can perform well without network actions (jindal2019neurlux, ), we do not expect this to affect our findings.
Early Detection of Malware. In our study, a detector uses the whole execution trace of a sample that terminates in a host. Although vendors offer ways to undo the damage from malware after its execution, e.g., quarantining (quarantine, ), it is more desirable to catch malware in its tracks. This would allow for containing the harm of a malware sample by cutting its execution short, e.g., before a ransomware sample starts encrypting personal files. As the detector has strictly less information available in this task, our results can be viewed as an upper bound on the performance of early malware detection. In consequence, we do not expect a detector that is trained only on sandbox traces to perform well in this setting.
Fine-Tuning a Sandbox-Based Model on Endpoint Traces. In Appendix D, we present an experiment where we first train a NGR model on sandbox traces and then fine-tune this model on increasing amounts of data from our EP training set. This hybrid approach simulates a realistic scenario where low-cost sandbox data is supplemented with high-cost endpoint traces from the wild. We evaluate two strategies: fine-tuning all layers of the model and fine-tuning only the last layer. We find that adding a few EP traces improves performance, and when very few EP traces are available, fine-tuning outperforms training from scratch (3–5% higher TPR).
Attacks on ML. All ML-based detectors, especially deep learning ones, are subject to adversarial attacks due to their sensitivity to input perturbations (szegedy2014intriguing, ; grosse2017adversarial, ). We expect our models will also be vulnerable to such attacks, e.g., an adversary can inject dummy actions into the behaviors of their malware sample designed to fool a model. We consider a defense against these attacks to be out of scope for this work; instead, we aim to show that naturally occurring factors, such as the environment variability, also degrade the performance of ML-based detectors in the wild.
10. Conclusion
Malware detection with dynamic features serves as a last line of defense at endpoints, providing security when all other measures have failed. Malware samples that breach this last line cause real-world infections and harm. A long-standing ambition is to find the best way to detect such samples. Through a systematic exploration of different scenarios, we provide some clarity. Existing detectors trained on sandbox executions degrade when deployed at endpoints. Training on real-world behaviors performs much better, though it necessitates collecting data from real hosts. We pinpoint various challenges in endpoint detection and propose ML techniques targeting them, yielding moderate improvements. Ultimately, this task remains challenging, and through our leaderboard, we hope to guide the community to develop effective and applicable solutions.
11. Acknowledgements
This research was supported by the Intelligence Community Postdoctoral Fellowship, the U.S. Department of Defense, EPSRC Grant EP/X015971/1, National Science Foundation Grant CNS-2154873, and generous gifts from Google and Amazon. We would like to thank VirusTotal for granting us academic access to their platform, which allowed us to curate and label our datasets. Finally, we thank Omer Faruk Akgul and David Acs for their support in the earlier phases of our research and Erin Avllazagaj for providing us with access to their program traces in the wild dataset. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the supporting organizations.
Appendix A Trace Standardization
First, we only keep the action types that all formats in our datasets share, which are file creation, registry key creation and deletion, process creation and injection, and mutex creation. Second, we clean up the strings (e.g., file names) in each trace by removing white spaces, capitalization, punctuation, non-ASCII characters, and so on. Third, we replace certain file and directory names to minimize differences caused by operating system versions or logging conventions. Fourth, we use regular expressions to replace MAC addresses, Windows security and resource identifiers, hashes (SHA-256, SHA-1, MD5), and epoch timestamps with special tokens (e.g. <macaddress>), to prevent introducing artifacts to our ML models. Fifth, we tokenize the entries in each trace (e.g., split a full file path into its components) and save a trace as a sequence of tokens, following (jindal2019neurlux, ). This standardized format allows us to implement a leaderboard for endpoint malware detection (detailed in Section 8) by providing all participants with an expected input format that their detectors should accept. When we train a model, we only keep the top-10K tokens in terms of frequency in the training traces and replace the remaining tokens with special tokens, such as <rare_file_name>, again following (jindal2019neurlux, ; trizna2023nebula, ). This makes the task more suitable for learning by eliminating uninformative tokens and reducing the feature dimensionality.
Appendix B Further Machine Learning Details
NGR (based on Maldy (karbab2019maldy, )). We convert each trace into a list of 2-grams. For example, the file path a/b/c.jpg is turned into three 2-grams: <a/b>, <b/c> and <c.jpg>. As this results in an intractable number of unique 2-grams (mostly very rare), we apply the hashing trick (weinberger2009feature, ) that assigns a numerical value up to $2^{14}$ to each 2-gram. The final feature vector—$x$—for a trace is a $2^{14}$-dimensional vector and each dimension is set to the number of occurrences of the corresponding 2-gram in the trace. We use a ResNet-based architecture (gorishniy2021revisiting, ) to train on these features.
HYB (based on Neurlux (jindal2019neurlux, )). We treat a trace as a natural language document and train an attention-convolution-hybrid sequence classification model. This approach eliminates the need for feature engineering (unlike the n-gram approach) and can extract useful features from long sequences thanks to the attention mechanism.
ATT (based on Nebula (trizna2023nebula, )). We treat a trace the same way as HYB (a sequence) and train a self-attention transformer-based architecture that is claimed to be robust to heterogeneous information (  e.g., different report format).
Overall, these approaches represent an increasing level of complexity, NGR being the simplest and most traditional and ATT being the most advanced. Although more advanced models seem to perform better in sandbox-based scenarios, we are interested in whether this trend changes in the endpoint scenario.
Appendix C Case Studies on Sandbox-Specific Artifacts
Here, we dive deeper into the sandbox-specific artifacts we identified in Table 10.
SogouExplorer is a Chinese web browser that exists only in malware traces (100% MalR.) in SB1; whereas it does not exist in any SB2 traces and very few EP traces. The samples that interact with it mostly belong to families such as Sivis and Memery, all tagged as file infectors that attach their code to other programs. Considering that SB1 was developed by a Chinese vendor, we believe that they pre-installed this browser on their sandboxes to generate an analysis environment representative of Chinese hosts. This, however, creates features specific to SB1 as samples interact with the programs in the environment. Although this artifact exists in a few endpoint traces from hosts in China, its prevalence is almost zero.
PersonalBankPortal, according to our research, is a program distributed by a Chinese bank to its customers. The samples that inject into this program belong to families such as Tinba and Ramnit, all considered as banking trojans that specifically ex-filtrate banking data. We believe the vendor pre-installs this program to lure malware samples into exhibiting their behaviors. Although this practice is useful for analyzing a sample (yong2021inside, ) (and for SB->SB), it causes artifacts that are rarely observed in the wild.
Among the artifacts found in SB2, Spotify is a popular music streaming service, and Python is the interpreter for Python programming language. Both programs are targeted and injected by file infectors, similar to SogouExplorer in SB1. These programs are much less prevalent in endpoint traces than in SB2 traces. We believe the SB2 vendor, which is based in the US, pre-installs them to create an environment representative of the hosts in the US.
Appendix D Fine-Tuning a Sandbox-Based Model on Endpoint Traces
For the experiments in Figure 10, we select two endpoint traces per sample and implement two fine-tuning strategies: (i) freezing the encoder layers $enc$ of our model and tuning only the classification layer $g$, and (ii) tuning all layers without freezing. We train models on increasing portions of the samples in our EP training set (randomly selected) and average the results over 10 models. We make the following observations. In low-data regimes (below 30% of the EP data), fine-tuning only $g$ outperforms the other options in terms of TPR by $\sim$3–5%. However, with more data, it starts to perform significantly worse, due to being less flexible in learning from the EP traces. Finally, fine-tuning all layers is generally the worst option, and training from scratch is the best when more EP data is available. These experiments show that starting from a well-performing model in SB->EP is beneficial in regimes with limited EP data, highlighting a promising direction for future work.
Appendix E More Details on the Leaderboard
The following artifacts will be released to the participants:
Sandbox Dataset. We will release the training portion of our sandbox datasets in Table 3, stored in the standardized format discussed in Section 4.1. We avoid releasing the testing portion to prevent participants from obtaining an impractical advantage by training their detectors on it (will be released at a later date). This will level the playing field for participants and ensure they all have access to the same sandbox data for development. Further, as this data contains traces from two sandboxes, participants can leverage our observation in Section 5 and tune their models on a second sandbox, not seen during training, for potentially higher endpoint performance. Moreover, they can also apply invariance learning techniques, which have been shown promising in Section 7.4.
Sample Metadata. We will release the metadata relating to the training samples, including their SHA-256 hashes, ground truth labels, family tags (if malware), and first-seen timestamps. Moreover, our metadata also tags the source of a sample, e.g., EMBER (anderson2018ember, ), SOREL (harang2020sorel20m, ), or our endpoint dataset. This allows participants to easily obtain realistic priors over malware families for endpoint detection, and make use of our findings and improvements in Section 6.3. Moroever, they can also create realistic testing distributions over samples using these malware family priors (and avoid the problems we discussed in Section 6.
Appendix F Additional Tables and Figures
Statistics on the Endpoint Traces
(a) The number of traces per sample.
(b) The average ratio of seen traces per sample as a function of time.
Figure 6. Statistics on the endpoint traces in our dataset.
Top Malware Families and Benign Publishers
EP Test Set SB Test Set Name EPSB1% Name SB1% EPGENERIC 23.8% 4.8% GENERIC 4.8% 23.8% Chindo 9.0% 0.0% Gepys 4.4% 0.0% Emotet 7.6% 0.5% Sivis 4.3% 0.3% Gandcrab 6.0% 3.6% Flystudio 4.2% 0.3% Loadmoney 6.0% 0.1% Upatre 3.8% 0.8% Khalesi 5.0% 0.3% Gandcrab 3.6% 6.0% Installcore 4.5% 0.0% Shipup 3.5% 0.0% UNSIGNED 29.6% 47.2% UNSIGNED 47.2% 28.4% Microsoft 7.8% 0.4% Google 4.7% 0.2% Tencent 2.6% 0.5% Mozilla 3.7% 0.1% Qihoo 2.2% 0.2% Digital R. 3.5% 0.0% Zoho 1.3% 0.1% Yandex 2.7% 1.0% Opera 1.2% 0.1% ScreenC. 2.2% 0.0% Yandex 1.0% 2.7% Zoom 2.0% 0.3%
Table 13. Top malware families (top) and benign (bottom) publishers in our EP (left) and SB (right) test sets, along with their shares in each dataset.
Soft Labeling Function Refer to caption
Figure 7. Our function for assigning probabilistic (soft) labels to each sample based on the number of VirusTotal detections.
Case Study on Wannacry and Its Indicators-Of-Compromise
Trace Type Ratio Avg. Pred. Score SB1 SB2 EP SB1->EP SB2->EP EP->EP Any IOC 89.9% 100% 97.7% 82.0% 75.6% 97.4% No IOC 10.1% 0.0% 2.3% 77.7% 67.7% 51.8%
Table 14. For Wannacry traces in our SB1, SB2, and EP datasets, we first measure the ratio of traces with at least one and no IOC. We then measure the average prediction score of our NGR models in SB1->EP, SB2->EP and EP->EP on these traces.
Rank Test Set Trained on SB1 Trained on SB2 Trained on EP SB1->SB1 SB1->SB2 SB1->EP SB2->SB1 SB2->SB2 SB2->EP EP->EP NGR HYB ATT NGR HYB ATT NGR HYB ATT NGR HYB ATT NGR HYB ATT NGR HYB ATT NGR HYB ATT SB1 99.1 99.0 98.9 93.8 91.3 92.6 76.1 70.7 72.8 89.5 83.4 89.1 97.7 94.1 96.9 72.0 68.2 70.8 85.6 85.2 84.3 SB2 99.0 98.6 98.8 94.9 93.2 93.5 76.6 71.3 74.4 81.7 71.5 86.4 98.7 98.5 98.2 71.6 66.7 72.7 85.4 85.9 84.8 EP 99.0 98.6 98.7 92.9 91.2 92.9 78.4 74.1 74.9 84.4 71.8 88.3 98.0 97.8 98.0 74.6 72.0 73.4 87.5 86.8 86.8
Table 15. The performance (AUC%) of three ML approaches (NGR, HYB, ATT) in seven detection scenarios (based on Table 2). In each row, the models are ranked based on the test set, and the average AUC of top-$20$ models is reported.
Trace Length Bias of the Sandbox-Based Model
Figure 8. Comparing the correlations between trace length and malware-ness prediction scores of the model. Malware ratio is the ground truth ratio of malware traces of a certain length among all traces of that length in a dataset.
Histogram of Prediction Score Standard Deviations on Endpoint Traces of the Same Sample for Models Trained With the Invariance Loss.
Figure 9. Prediction score standard deviations on the endpoint traces of each sample in the EP test set; for a model in SB1->EP (left), and EP->EP (right).
Fine-Tuning a Sandbox-Based Model on Endpoint Traces Refer to caption
Figure 10. The results of fine-tuning a sandbox-based model in SB->EP on endpoint traces. The dashed line indicates our best model SB->EP. Experiments on NGR.

### v2

@@
-Demystifying Behavior-Based Malware Detection at Endpoints
-Abstract.
-Malware detection is one of the most ubiquitous applications of machine learning (ML) in security.
-Prior ML-based detectors can consume program features extracted statically from binary code or dynamically from execution traces.
-Dynamic detectors most commonly rely on executing programs in synthetic environments (sandboxes) that allow controlled malware analysis.
-However, sandbox traces are unavailable to the last line of defense offered by security vendors: real-time malware detection at endpoints.
-A detector at endpoints consumes the traces of programs running on hosts (i.e., real-world environments), as sandbox analysis, typically done in the cloud, might introduce intolerable delays.
-Despite their success in the sandbox setting, research hints at potential challenges for ML methods at endpoints, e.g., highly variable malware behaviors in the wild.
-Nonetheless, the impact of these challenges on existing approaches and how their excellent sandbox performance translates to the endpoint scenario remain unquantified.
+ML-Based Behavioral Malware Detection Is Far From a Solved Problem
+Abstract
+Malware detection is a ubiquitous application of Machine Learning (ML) in security.
+In behavioral malware analysis, the detector relies on features extracted from program execution traces.
+The research literature has focused on detectors trained with features collected from sandbox environments and evaluated on samples also analyzed in a sandbox.
+However, in deployment, a malware detector at endpoint hosts often must rely on traces captured from endpoint hosts, not from a sandbox.
+Thus, there is a gap between the literature and real-world needs.
 We present the first measurement study of the performance of ML-based malware detectors at real-world endpoints.
-Leveraging a dataset of sandbox traces we collected and a dataset of in-the-wild program traces from Avllazagaj et al.
-(avllazagaj2021malware, ); we evaluate two scenarios where the endpoint detector was trained on (i) sandbox traces (convenient and accessible); and (ii) endpoint traces (less accessible due to needing to collect telemetry data).
-This allows us to identify a wide gap between prior methods’ sandbox-based detection performance—over 90%—and endpoint performance—about 20% (scenario (i)) to 50% (scenario (ii)).
-We pinpoint and characterize the challenges contributing to this gap, including label noise, behavior variability, sandbox evasion, or environment-specific features.
-Building on this, we suggest strategies to close the gap, yielding a relative improvement of 5–30% over the baselines.
+Leveraging a dataset of sandbox traces and a dataset of in-the-wild program traces, we evaluate two scenarios: (i) an endpoint detector trained on sandbox traces (convenient and easy to train), and (ii) an endpoint detector trained on endpoint traces (more challenging to train, since we need to collect telemetry data).
+We discover a wide gap between the performance as measured using prior evaluation methods in the literature—over 90%—vs.
+expected performance in endpoint detection—about 20% (scenario (i)) to 50% (scenario (ii)).
+We characterize the ML challenges that arise in this domain and contribute to this gap, including label noise, distribution shift, and spurious features.
+Moreover, we show several techniques that achieve 5–30% relative performance improvements over the baselines.
 Our evidence suggests that applying detectors trained on sandbox data to endpoint detection is challenging.
 The most promising direction is training detectors directly on endpoint data, which marks a departure from current practice.
-To promote research, we implement a leaderboard for realistic detector evaluations—a format that has fueled progress in many ML tasks.
-malware detection, dynamic analysis, machine learning
-copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Security and privacyccs: Security and privacy Intrusion/anomaly detection and malware mitigationccs: Security and privacy Malware and its mitigation
-1. Introduction
-Detection of malware threats is crucial for governments, enterprises, and end users as there are significant (and growing) financial and safety harms of malware infections (reddick_2022, ), which has created a \$7 Billion industry in 2022 with many players (imarcreport, ).
-Malware detection appears to be remarkably effective: industry-standard evaluations of commercial anti-malware products (avcomparisons, ), and prior research on malware detection using machine learning (ML) methods (raff2018malware, ; karbab2019maldy, ), routinely report that over 99% of malware samples in a standard corpus can be detected, with very few false-positives.
+To promote progress, we will facilitate researchers to perform realistic detector evaluations against our real-world dataset.
+I Introduction
+Detection of malware threats is crucial for governments, enterprises, and end users as there are significant (and growing) financial and safety harms of malware infections [1], which has created a \$7 Billion industry in 2022 with many players [2].
+Malware detection appears to be remarkably effective: industry-standard evaluations of commercial anti-malware products [3], and prior research on malware detection using machine learning (ML) methods [4, 5], routinely report that over 99% of malware samples in a standard corpus can be detected, with very few false positives.
+  This work has been accepted for publication in the IEEE Conference on Secure and Trustworthy Machine Learning (SaTML). The final version will be available on IEEE Xplore.
 Training data Test data Method TPR @ 1%FPR
-Sandbox traces Sandbox traces Standard classifier $\sim$95(karbab2019maldy, ; jindal2019neurlux, ; trizna2023nebula, )
+Sandbox traces Sandbox traces Standard classifier $\sim$95[5, 6, 7]
 Sandbox traces Endpoint traces Standard classifier $\sim$17(ours)
@@
 Endpoint traces Endpoint traces Soft labels + invariant learning $\sim$52(ours)
-Table 1. Performance of ML-based behavioral malware detection under different settings.
-Malware detection is often performed at endpoints where an endpoint security solution (quarantine, ) monitors host devices to detect threats in real time.
-In practice, these solutions rely on a chain of techniques (vmray, ), including static analysis and dynamic analysis.
-Conventional static methods, such as blocklists, reputation systems, and signatures, operate without executing the program.
-As static methods can be bypassed through obfuscation or polymorphism (moser2007limits, ; song2007infeasibility, ), dynamic analysis has become standard, offered by most vendors (kasperskyrealtime, ).
+TABLE I: Performance of ML-based behavioral malware detection under different settings.
+Malware detection is often performed at endpoints, where an endpoint security solution [8] monitors host devices to detect threats.
+In practice, these solutions employ a chain of techniques [9], including static analysis and dynamic analysis.
+Static methods, such as blocklists and signatures, operate without executing the program.
+As static methods can readily be bypassed via obfuscation or polymorphism [10, 11], dynamic analysis has become a standard offering [12].
 Dynamic analysis relies on observing the execution behaviors of a sample to detect whether it displays malicious behaviors.
-The typical paradigm for obtaining such behavior-based detectors is to detonate (execute) a large set of known samples (malware and benign) in a controlled environment (a sandbox), collect their execution traces, and learn to separate malicious and benign behaviors.
-Several ML models have demonstrated exceptional performance in this task (mariconti2016mamadroid, ; karbab2019maldy, ; jindal2019neurlux, ; trizna2023nebula, ).
-Unfortunately, sandbox traces are unavailable to models in endpoint detection as detonation cannot be done in real time, despite advances such as on-premise sandboxes (onpremsandbox, ).
-Consequently, these models must rely on traces observed on real-world hosts to detect malware.
+The typical paradigm for obtaining a behavioral detector is to detonate (execute) a large set of known samples (malware and benign) in a controlled environment (a sandbox), collect their execution traces, and learn to separate malicious and benign behaviors.
+Several ML models have been effective in this task [13, 5, 6, 7].
+Unfortunately, endpoint detection models cannot use sandbox traces to decide whether a sample under observation is malicious, as detonation cannot be done in real time, despite advances such as on-premise sandboxes [14].
+Consequently, these models must rely on traces observed at real-world hosts to detect malware.
 Research has outlined several challenges associated with this task.
-First, program behaviors are environment-sensitive, and, especially for malware, two traces of the same sample from two different hosts often have little overlap (avllazagaj2021malware, ).
-As a result, a model trained on a trace captured on one host might fail to generalize to other hosts.
-Second, techniques frequently employed to evade sandbox analysis (lindorfer2011detecting, ; balzarotti2010efficient, ; galloro2022systematical, ) can mislead models trained on sandbox traces into learning features that do not capture actual malware behaviors.
-Although these challenges were acknowledged in prior work (barecloud, ; kirat2015malgene, ; jindal2019neurlux, ; avllazagaj2021malware, ), we lack understanding of how they manifest, their implications for endpoint detectors, and potential solutions.
+First, program behaviors are environment-sensitive, and, especially for malware, two traces of the same sample from two different hosts often have little overlap [15].
+As a result, an ML model trained on a trace captured on one host might fail to generalize to others.
+Second, techniques frequently employed to evade sandbox analysis [16, 17] can mislead models trained on sandbox traces into learning spurious features that do not capture actual malware behaviors.
+Although these challenges have been previously articulated in the research literature [18, 19, 6, 15], we lack an understanding of how they manifest and can be addressed.
 Our work conducts the first quantitative study into the efficacy of ML-based methods for endpoint detection to demystify these challenges.
 We specifically focus on two scenarios: (i) an endpoint detector trained using only sandbox traces and (ii) an endpoint detector trained using real-world endpoint traces.
-Existing techniques can conveniently gather large-scale datasets of sandbox traces, making (i) a practical and accessible option.
-On the other hand, collecting endpoint traces is attainable mainly by vendors who receive telemetry data from the wild as it is still an unsolved problem to simulate such data in controlled environments (such as sandboxes) at scale (schreuders2017security, ; miramirkhani2017spotless, ; mills2020investigating, ).
-These requirements make (ii) a less accessible but likely a more successful option.
+Existing techniques can conveniently gather large datasets of sandbox traces, making (i) a practical and accessible option.
+On the other hand, collecting endpoint traces is attainable mainly by vendors who receive telemetry data from the wild, as simulating such data at scale in controlled environments (such as sandboxes) is still an open problem [20, 21, 22].
+This makes (ii) a less accessible but likely a more effective option.
 Consequently, both scenarios are relevant in practice, which motivates us to study their challenges.
-We employ three deep-learning-based approaches (karbab2019maldy, ; jindal2019neurlux, ; trizna2023nebula, ) that can achieve $\sim$95true-positive rate (TPR) at 1% false-positive rate (FPR) when evaluated on sandbox traces.
+We evaluate three ML approaches [5, 6, 7] that reach 95% true-positive rate (TPR) at 1% false-positive rate (FPR) when evaluated on sandbox traces.
 We use a dataset from Avllazagaj et al.
-(avllazagaj2021malware, ) of around 1M endpoint traces from over 25K samples, recorded on real-world hosts by a commercial anti-malware vendor.
-The malicious samples in this dataset were undetected by the vendor’s defenses at that time, corresponding to real-world infections.
-Consequently, this data reflects the realistic threats that behavior-based detectors must combat.
-Additionally, to study scenario (i), we collect a dataset of traces (contemporaneous to our endpoint dataset) from two sandboxes.
-Our initial finding is that ML methods exhibit lower performance in both scenarios (i) ($\sim$17TPR) and (ii) ($\sim$49TPR), in contrast with their excellent performance on sandbox traces (see Table 1).
-We study the low performance in (i) from the lens of distribution shift, a problem that plagues ML in many applications (koh2021wilds, ).
-Others in the security community have articulated the challenge of concept drift, where the data distribution shifts over time (jordaney2017transcend, ).
-We explore whether the problem can be addressed using existing tools for concept drift.
-However, we found that existing concept drift methods are not enough to address the differences in data distribution between sandbox vs endpoint traces: we discovered that endpoint traces of benign samples are detected as “drifting” (whereas in prior work concept drift has mostly affected malware (barbero2022transcending, )), and rejecting drifting samples does not improve the performance enough.
-Investigating the root causes of the low performance, we discovered that endpoint detectors are applied to a different distribution of samples than considered in prior work: they are typically applied only to the hardest-to-classify samples.
-In past research, researchers have trained and evaluated detectors on a corpus of samples from repositories (karbab2019maldy, ; jindal2019neurlux, ; dambra2023decoding, ) such as MalwareBazaar (bazaar, ).
-However, endpoint malware detection systems employ a pipeline, where samples are first categorized using basic methods (e.g., signatures, static analysis), and ML-based classifiers are only applied to samples that cannot be resolved using basic methods (vmray, ).
-Consequently, in the wild, endpoint classifiers are only applied to samples that tend to be harder to classify than those found in a standard corpus.
-Prior research has not considered this factor, and we find that it causes a significant drop in performance: it reduces the performance of classifiers trained on sandbox traces from 95% (for samples from a standard corpus) to $\sim$60TPR (when the distribution of samples is adjusted to take into account earlier stages in the pipeline).
-This shows that prior evaluations of behavior-based detectors have greatly over-estimated their effectiveness in the wild.
+[15] containing around 1M endpoint traces from 26K samples, recorded on real-world hosts by an anti-malware vendor.
+The malicious samples in this dataset were undetected by the vendor’s defenses at that time and caused real-world infections.
+Consequently, this data reflects the realistic threats that behavioral detectors must combat.
+Additionally, we collect a dataset of traces (contemporaneous to our endpoint data) from two sandboxes to realistically study scenario (i).
+Our initial finding is that ML methods perform poorly in both scenarios (i) ($\sim$17TPR) and (ii) ($\sim$49TPR), in contrast with their excellent performance on sandbox traces (see Table I).
+We study the low performance in (i) from the lens of distribution shift, a problem that plagues ML in many applications [23].
+Similarly, the security community has articulated the challenge of concept drift, where the data distribution shifts over time [24].
+We attempt to address the differences in data distribution between sandbox vs.
+endpoint traces using classification with rejection (an existing tool against concept drift [24]) and found that it does not improve the performance enough.
+Interestingly, endpoint traces of benign samples are also rejected as “drifting” in high proportions, whereas in prior work concept drift has mostly affected malware [25].
+Investigating why existing methods perform poorly, we first discovered that endpoint detectors are applied to a different distribution of samples than considered in prior work: they are typically applied only to the hardest-to-classify samples.
+Past research has trained and evaluated detectors on a corpus of samples from repositories [5, 6, 26] such as MalwareBazaar [27].
+However, endpoint malware detection systems employ a pipeline where basic methods (e.g., static signatures) attempt to categorize samples first, and ML-based classifiers are applied only to samples that remain unresolved [9].
+Consequently, in the wild, endpoint classifiers are only applied to samples that tend to be harder to classify than an average sample in a standard corpus.
+Prior research has overlooked this factor, and we find that it causes a significant drop in performance: it reduces the performance of classifiers trained on sandbox traces from 95% (for samples from a standard corpus) to $\sim$60TPR (when we adjust the distribution of samples to take into account earlier stages in the pipeline).
+This shows that prior evaluations have vastly overestimated the effectiveness of behavioral detectors in the wild.
 Second, we study the impact of variable program behaviors across different environments.
-We measure large differences between a sample’s sandbox trace and its endpoint traces.
-Sandbox traces lack diversity: collecting multiple traces by running a group of related samples, e.g., from the same malware family, in a sandbox yields very similar traces.
+We discovered large differences between a sample’s sandbox trace and its endpoint traces.
+Sandbox traces lack diversity: collecting multiple traces by running a group of related samples, e.g., from the same malware family, in a sandbox produces very similar traces.
 This introduces spurious features that do not generalize to other environments.
-Conversely, endpoint traces are highly diverse, making a model’s predictions on different traces of the same sample inconsistent, which hurts detection performance.
-Moreover, we provide the first evidence that sandbox-evading malware (40–80% of all malware (galloro2022systematical, )) skews a model in scenario (i) towards classifying very short traces (often an indication of evasion (kuchler2021does, )) as malware.
-This correlation is spurious as it is absent in endpoint traces.
-Thus, this correlation causes prior evaluations using sandbox traces to over-estimate the accuracy of endpoint classifiers.
-Our characterization of these challenges allows us to pinpoint avenues for improvement in endpoint detection.
-To improve the performance on the distribution of difficult-to-classify samples, we explore soft-labeling (effective against label noise (lukasik2020does, )) and using more accurate distributions for training.
-Against variable behaviors, we employ a technique popular in other areas of ML: invariant learning (zhao2019learning, ).
-In particular, we train our models to make consistent predictions on different traces of the same sample to prevent them from learning unreliable, environment-specific features.
-When combined, these strategies lead to moderate but promising gains, from 17% to 22% TPR in scenario (i) (30% relative improvement) and from 49% to 52% TPR in scenario (2) (5% relative improvement).
-We believe our results should serve as a call to action for the research community.
-Previous research has suggested that it is possible to achieve 95% TPR, which might leave an impression that progress is saturated, there is not much room for further improvement, and the problem of ML-based malware detection is more or less solved.
-Our results show that the reality is different: the problem is not solved, the actual performance on malware in the wild is far worse, and there are major unsolved challenges and significant room for further improvement.
-To stimulate further research on this challenging problem, we have designed a leaderboard (malwaredetectioninthewild.github.io (leaderboard, )), a popular format in ML (croce2021robustbench, ), for realistic evaluations of behavior-based malware detectors.
-Participants receive access to our sandbox dataset and all our metadata and can submit their detectors, which we will rank according to their endpoint performance.
-This will hopefully allow the community to identify promising directions.
+In contrast, endpoint traces are diverse, which hurts performance by making a model’s predictions on different traces of the same sample inconsistent.
+Moreover, we provide the first evidence that sandbox-evading malware (40–80% of all malware [17]) biases a model in scenario (i) to classify very short traces (often an indication of evasion [28]) as malware.
+This correlation is spurious as it is absent in endpoint traces; thus, it causes prior evaluations using sandbox traces to overestimate the accuracy of endpoint classifiers.
+Characterizing these challenges allows us to explore avenues for performance improvements.
+To improve the performance on the distribution of difficult-to-classify samples, we use soft labeling (effective against label noise [29]) and more accurate distributions for training.
+Against variable behaviors, we employ a technique popular in other areas of ML: invariant learning [30].
+In particular, we train our models to make consistent predictions on different traces of the same sample, forcing them to learn robust, environment-independent features.
+Together, these strategies lead to moderate gains, from 17% to 22% TPR in scenario (i) and from 49% to 52% in scenario (ii) (30% and 5% relative improvement, respectively).
+We believe our results should serve as a call to action for the community.
+Prior research has suggested that it is possible to achieve over 95% TPR, which might leave the impression that progress is saturated and there is not much room for improvement in behavioral malware detection using ML.
+We show that the reality is different: the problem is not solved, the actual performance on malware in the wild is far worse, and there are major unsolved challenges and significant room for further improvement.
+Moreover, ML methods that have shown success on standard benchmarks, such as group robustness [31], struggle due to the complexities of this domain.
+We plan to stimulate progress on this problem by releasing our sandbox dataset and metadata and offering a pipeline that allows researchers to evaluate their behavioral malware detectors against our real-world endpoint data.
+Our public website (https://malwaredetectioninthewild.github.io/) includes the details on this evaluation pipeline and data release.
 Contributions.
-(I) We measure the performance discrepancy of ML-based approaches between sandbox-only and endpoint settings (Section 5).
-(II) We identify the challenges in endpoint detection to demystify this discrepancy and propose improvements (Section 6 and Section 7).
-(III) We organize a leaderboard for realistic evaluations of detectors (Section 8).
-2. Background and Related Work
+(I) We measure the performance discrepancy of ML-based malware detectors between sandbox-only and endpoint settings.
+(II) We characterize the ML challenges, such as distribution shifts, behind this discrepancy.
+(III) We explore ML methods to improve the endpoint performance.
+II Background and Related Work
 Dynamic Malware Analysis.
-Most work in dynamic analysis focuses on analyzing the behaviors of a sample detonated in controlled environments, such as sandboxes (willems2007toward, ; bayer2009view, ).
-As dynamic analysis has become a staple, malware has started including checks that suppress malicious activities if the environment is fingerprinted as a sandbox, known as evasive malware (lindorfer2011detecting, ).
-Although many strategies have been developed to prevent fingerprinting (liu2022enhancing, ), this is still an ongoing arms race (miramirkhani2017spotless, ).
-Although researchers have explored methods to analyze samples in bare metal environments (balzarotti2010efficient, ; barecloud, ), they fall short in preventing evasion (miramirkhani2017spotless, ).
+Most work in dynamic analysis focuses on analyzing the behaviors of a sample detonated in controlled environments, such as sandboxes [32, 33].
+As dynamic analysis has become a staple, malware has started including checks that suppress malicious activities if the environment is fingerprinted as a sandbox, known as evasive malware [34].
+Although many strategies have been developed to prevent fingerprinting [35], this is still an ongoing arms race [21].
+Even the methods that analyze samples in bare metal environments [16, 18] struggle to prevent evasion [21].
 To our knowledge, we have taken the first step toward measuring the implications of evasion for ML-based detectors at endpoints.
-A recent large-scale measurement study over variable program behaviors in the wild has found that a given malware sample can behave significantly differently across time and in different real-world hosts (avllazagaj2021malware, ).
+A recent large-scale measurement study over variable program behaviors in the wild has found that a given malware sample can behave significantly differently across time and in different real-world hosts [15].
 Our work connects to this line of work as we are interested in quantifying the impact of these challenges, such as evasion or variability, on a malware detector deployed for endpoint detection at hosts in the wild.
 ML for Malware Detection.
-ML-based methods are widely used in research and practice with static (raff2018malware, ) and dynamic features (huang2016mtnet, ).
-In detection with dynamic features, methods that essentially treat a program’s execution trace as a natural language document (a sequence of tokens) and adapt off-the-shelf methods from the ML community have been successful (miller2016reviewer, ; karbab2019maldy, ; jindal2019neurlux, ; kuchler2021does, ; trizna2023nebula, ).
-Most of these methods are trained and evaluated on the traces from a pre-configured sandbox on samples collected from public repositories; some popular ones are listed here (awesomemalware, ).
-We aim to understand the implications of these practices and the efficacy of popular ML approaches for detecting malware in real-world hosts with dynamic features.
+ML-based methods are widely used in research and practice with static [4] and behavioral features [36].
+In behavioral detection, methods that treat a program’s execution trace as a text document (a sequence of tokens) and adapt existing ML architectures have shown promise [5, 6, 7].
+Most of these methods are trained and tested on the traces from a pre-configured sandbox for samples collected from public repositories (see [37] for the popular ones).
+We aim to understand the implications of these practices and the efficacy of popular ML approaches for detecting malware with behavioral features in real-world hosts.
 Distribution Shifts in ML.
-Distribution shift occurs when a model’s training and testing distributions have significant differences, which hurts the performance (koh2021wilds, ).
-For example, as new malware variants emerge and old ones disappear over time, the performance of a detector that has not been kept up-to-date will deteriorate, known as concept drift (pendlebury2019tesseract, ).
-A popular paradigm to deal with concept drift is deploying drift detectors to reject samples that would have been misclassified and training on them later to update the model (jordaney2017transcend, ; barbero2022transcending, ; chen2023continuous, ).
-We deploy a drift detector to characterize the differences between the distribution shifts in our problem and concept drift.
-Although there is no silver bullet, the ML community has also proposed many ideas to tackle distribution shifts in different contexts, such as learning domain invariant features (ganin2015unsupervised, ; 2013gongconnecting, ), distributionally robust optimization (sagawa2020investigation, ), or continuous learning (chen2023continuous, ).
-We adapt some of these ideas, such as invariant learning, to our problem to measure their benefits for malware detection at endpoints.
+Distribution shift occurs when a model’s training and testing distributions have significant differences, which hurts the performance [23].
+For example, as new malware variants emerge and old ones disappear over time, the performance of a detector that has not been kept up-to-date will deteriorate, known as concept drift [38].
+An effective way to deal with concept drift is deploying drift detectors to reject samples that would have been misclassified and training on them later to update the model [24, 25, 39].
+We deploy a drift detector to characterize how the distribution shifts in our problem differ from concept drift.
+The ML community has also proposed other ideas to tackle distribution shifts in different contexts, such as domain invariant features [40, 41], distributionally robust optimization [31], or continuous learning [39].
+We adapt some of these ideas, such as invariant learning, to assess their benefits in our problem.
 Limitations of ML for Security.
-Research suggests that popular ML methods have many pitfalls when applied to security tasks (arp2020and, ).
-For example, in malware detection, most work has been found to overestimate the success of the methods due to impossible time splits of training and testing sets (pendlebury2019tesseract, ).
-A line of work also focuses on how ML methods successful in lab-only evaluations break down when they are deployed in the real world due to distribution shifts, for example, in the context of network anomaly detection (jacobs2022ai, ), website fingerprinting (cherubin2022online, ) or malware detection (das2019sok, ; aghakhani2020malware, ).
-In these contexts, models are known to learn spurious features, such as specific IP addresses (jacobs2022ai, ), that are artifacts of the experimental setup in the lab and do not apply to realistic settings.
-Our work also investigates the limitations of ML methods in malware detection with dynamic features when they are trained and evaluated in controlled settings (e.g., using sandboxes) and deployed in the wild.
-3. Endpoint Malware Detection
+Research suggests that popular ML methods have many pitfalls when applied to security tasks [42].
+For example, in malware detection, the success of many methods has been overestimated due to impossible time splits of training and testing sets [38].
+Another line of work focuses on how ML methods successful in lab-only evaluations break down in the real world due to distribution shifts, e.g., in the context of network anomaly detection [43], website fingerprinting [44] or malware detection [45, 46].
+In these contexts, models are known to learn spurious features, such as specific IP addresses [43], that are artifacts of the experimental setup and do not apply to realistic settings.
+We investigate the limitations of ML methods specifically in behavioral malware detection when they are trained and evaluated in controlled settings (e.g., using sandboxes) but deployed in the wild.
+III Endpoint Malware Detection
 Terminology.
-A sample is an executable program, identified by its unique SHA-256 hash.
+A sample is an executable program identified by its unique SHA-256 hash.
 A trace is a sequence of behavioral actions (such as file accesses or process creations) performed by a sample when it is executed in a computing environment.
-Our work focuses on Windows environments and executable samples.
 We refer to an in-the-wild endpoint environment as a host.
-In our endpoint dataset, the anti-malware system in each host recorded the traces of samples executed by the host that were not determined to be benign or malicious in an earlier stage of the detection pipeline.
-Each sample can have multiple traces recorded at multiple hosts at different timestamps.
-A sandbox (SB) is a synthetic environment that provides tools for analyzing samples and recording their traces through controlled execution.
-Our sandbox dataset contains traces collected from two sandboxes from two vendors, namely Tencent HABO (habo, ) and an in-house version of Cuckoo Sandbox (cuckoo, ).
-In the rest of the paper, we will refer to these sandboxes as SB1 and SB2, respectively.
+Our work focuses on Windows hosts and executable samples.
+In our endpoint dataset, the anti-malware system in each host recorded the traces of samples executed by the host that were not determined to be benign or malicious at an earlier stage of the detection pipeline.
+Each sample can have multiple traces recorded at multiple hosts at different times.
+A sandbox (SB) is a synthetic environment that provides tools for analyzing samples and recording their traces with controlled executions.
+Our sandbox dataset contains traces collected from two commercial sandboxes: Tencent HABO [47] and VirusTotal [48], which we will refer to as SB1 and SB2, respectively.
 Notation.
-We denote the set of samples in our dataset as $\mathbf{P}$.
-$P_{i}$ is the $i$-th program and $y_{i}$ is its ground truth label, where $y\!=\!0$ and $y\!=\!1$ indicate a benign and malware sample, respectively.
+We denote an individual sample in our dataset as $P_{i}$ and its ground truth label as $y_{i}$, where $y_{i}\!=\!0$ and $y_{i}\!=\!1$ indicate a benign and malware sample, respectively.
 If $P_{i}$ is a malware sample, it is also tagged with a family label $s_{i}$ that is useful for grouping samples with similar characteristics.
@@
 The predicted label $\hat{y_{i}}$ is $\hat{y_{i}}\!=\!1$ if $f(x_{i})\!\geq\!\tau$, or $\hat{y_{i}}\!=\!0$, otherwise.
-Here, $\tau$ is a threshold tuned based on the requirements, where a higher $\tau$ results in fewer false positives in exchange for fewer true positives.
-In Section 3.3, we discuss the tuning of $\tau$ in more detail.
-Figure 1. An overview of endpoint malware detection.
-3.1. Malware Detection Workflow
-Figure 1 provides an overview of the workflow for malware detection at endpoints.
-We focus on behavior-based detectors deployed in step 3.
-There are two main phases from the vendor’s perspective: training and testing.
+Here, $\tau$ is a tunable threshold, where a higher $\tau$ reduces false positives in exchange for fewer true positives.
+In Section III-C, we discuss the tuning of $\tau$ in more detail.
+Figure 1: An overview of endpoint malware detection.
+III-A Malware Detection Pipeline
+Figure 1 provides an overview of the pipeline for malware detection at endpoints.
+In practice, vendors deploy a chain of techniques [3, 9] to detect if a sample $P$ encountered by a host (e.g., from an e-mail attachment) is malware—step 2 and 3.
+If one technique in the pipeline fails to classify the sample as malware or benign conclusively, the next one is invoked.
+First, it deploys high-precision methods (blocklists and signatures), then less precise static ML approaches [49, 4], and finally, dynamic and behavioral analysis.
+We focus on ML-based behavioral malware detectors deployed in step 3.
+Static techniques operate without executing the sample and are deployed first as they are more efficient and secure—step 2.
+This is sufficient for classifying most existing samples [26] either as malware ($P$ is prevented from executing) or benign ($P$ is allowed to execute).
+If $P$ remains undetected after this step, it will be executed if the host demands it, e.g., by clicking on the attachment.
+If $P$ is malware, executing it leads to a real-world compromise—an infection.
+An endpoint detector is the last line of defense that operates at the host in real-time to identify whether $P$ is malware by consuming the trace resulting from this execution—step 3.
+If $P$ is detected as malware during this step, anti-malware products often apply remedies such as quarantining, deleting files, or alerting [8].
+We identified two issues regarding this pipeline in past research.
+First, prior work constructed evaluation sets for behavioral malware detectors without considering their position in an end-to-end pipeline.
+As a result, the challenges in performing behavioral detection on samples that bypassed static methods are unknown.
+Second, endpoint detectors must classify the sample based on an endpoint trace observed at the end host.
+It is not feasible to first execute the sample in a sandbox and then classify it based on the sandbox trace, as running in a sandbox (whether in the cloud [50] or on-premise [14]) introduces unacceptable delays.
+Prior work has only evaluated ML detectors on sandbox traces, but in deployment, the detectors will be applied to endpoint traces, so their actual performance has not been accurately evaluated.
+We tackle these issues by performing realistic evaluations using our endpoint dataset that contains only the traces of undetected samples—which reached the step 3 in the wild.
+Assumptions and Limitations.
+Security vendors such as VMRay [9], Palo Alto Networks [51], and Kaspersky [52] use multi-tiered solutions that chain static and dynamic analysis in their anti-malware products.
+We do not have access to these proprietary pipelines.
+Instead, we consider a streamlined one in Figure 1 based on public documentation by these vendors.
+Such a pipeline should ideally be evaluated end-to-end, starting with a large pool of samples from the wild, where each component sees only the samples that previous components cannot classify.
+We are unable to perform such an evaluation as we do not have raw binaries (to extract static features) of our endpoint samples (we could find less than 10% of samples from public sources, mostly malware).
+Prior malware detection datasets (e.g., SOREL [53]) only share vectorized static features and not the binaries, preventing the simulation of behavioral components.
+However, we can still realistically isolate and evaluate behavioral ML methods at the pipeline’s last stage as our endpoint dataset contains samples that bypassed a real-world product’s static components.
+We recommend future work to collect datasets (with static and dynamic features) that allow the evaluation of the whole pipeline end-to-end or each component according to its position.
+Moreover, a pipeline’s components constantly evolve in practice.
+For example, a sample from a new malware family might initially bypass static methods.
+Eventually, this family can be caught at step 2 with new static signatures or updates to static ML models.
+Evaluating this requires a dataset with a large temporal span, whereas our endpoint data covers only six months (insufficient for significant changes [25]).
+Consequently, we perform our evaluation with fixed behavioral ML models that are trained at a particular timestamp and tested on future samples that bypass static methods.
+III-B Model Training and Evaluation Scenarios
+There are two main phases for deploying an ML-based endpoint detector: training and testing.
 In the training phase, the vendor collects a set of samples.
-These samples are often collected from large-scale repositories, either public, such as MalwareBazaar (bazaar, ), or private (ugarte2019close, ).
-Most prior work in this domain has followed this practice (jindal2019neurlux, ; miller2016reviewer, ; karbab2019maldy, ; dambra2023decoding, ; kuchler2021does, ).
-The vendor can rely on a public platform, such as VirusTotal (virustotal, ) to label these samples, or for more recent or advanced malware, can allocate malware analysts for manual labeling (yong2021inside, ).
+These samples are often collected from large-scale repositories, either public, such as MalwareBazaar [27], or private [54].
+Most prior work in this domain has followed this practice [6, 55, 5, 26, 28].
+The vendor can rely on a public platform, such as VirusTotal [48], to label these samples or employ malware analysts in more advanced cases [56].
 Most commonly, the collected samples are executed in sandboxes that can conveniently produce large-scale datasets of traces for training.
 In contrast, collecting a large-scale dataset of endpoint traces from hosts requires a telemetry infrastructure, a permissive user agreement, and a large enough user base—conditions that commercial anti-malware vendors can meet but researchers typically cannot.
-The training phase concludes with training an ML model on the collected traces for detecting malware from runtime behaviors.
-During the test phase, the model is deployed locally at the hosts.
-Although we view them as sequential, these phases can be interwoven as vendors continuously update their models on new data (chen2023continuous, ).
-In practice, vendors deploy a chain of techniques (avcomparisons, ; vmray, ) to detect if a sample $P$ encountered by a host (e.g., from an e-mail attachment) is malware—step 2 in Figure 1.
-First the sample is analyzed with static techniques, which operate without executing the sample.
-This is sufficient for classifying most existing samples (dambra2023decoding, ) either as malware ($P$ is prevented from executing) or benign ($P$ is allowed to execute).
-If $P$ remains undetected after step 2, it will be executed if the host demands it, e.g., by clicking on the attachment.
-Critically, our endpoint dataset contains only the traces recorded in the wild after this step, i.e., the traces of undetected samples.
-The challenges in performing behavior-based detection on such samples are unknown, as prior work has not considered it a criterion for constructing evaluation sets.
-If $P$ is malware, executing it leads to a real-world compromise—an infection.
-An endpoint detector is the last line of defense that operates at the host in real time to identify whether $P$ is malware by consuming the trace resulting from this execution—step 3.
-If $P$ is detected as malware during this step, anti-malware products often apply remedies such as quarantining, deleting files, or alerting (quarantine, ).
-Endpoint detectors must classify the sample based on an endpoint trace observed on the end host.
-It is not feasible to first execute the sample in a sandbox and then classify it based on the sandbox trace, as running in a sandbox (whether in the cloud (cloudsandbox, ) or on-premise (onpremsandbox, )) introduces unacceptable delays.
-Prior research has only evaluated ML detectors on sandbox traces, but in deployment, the detectors will be applied to endpoint traces, so their actual performance has not been evaluated in the research literature to date.
-We tackle this issue in this paper.
-3.2. Model Training and Evaluation Scenarios
-Table 2 summarizes the detection scenarios, each defined along four dimensions, in our focus.
-The Samples dimensions indicate the step of the workflow training or testing samples originate from: 1—a broad distribution of all programs hosts encounter—or 3—a distribution of programs that remained undetected after 2.
+The collected traces are used to train an ML model that can detect malware from runtime behaviors.
+During the test phase, the model is deployed locally at each host.
+Training and testing phases can also be interwoven as vendors continuously update their models on new data [39].
+Table II summarizes the detection scenarios, each defined along four dimensions, considered in our work.
+The Samples dimensions indicate the step of the pipeline training or testing samples originate from: 1—a broad distribution of all programs hosts encounter—or 3—a distribution of programs that remained undetected after 2.
 The Env.
@@
 Deployment Scenario Training Phase Testing Phase Samples Env. Samples Env. SB->SB Step 1 Sandbox Step 1 Sandbox SB->EP Step 1 Sandbox Step 3 Endpoint EP->EP Step 3 Endpoint Step 3 Endpoint
-Table 2. Scenarios for behavior-based malware detection.
-SB->SB is the scenario most prior work in our domain considers: both training and testing samples are collected from sample repositories and corpora, which represent the distribution of all samples seen in the wild, and these samples are executed in a sandbox (karbab2019maldy, ; jindal2019neurlux, ; miller2016reviewer, ; kuchler2021does, ).
-It is already known (and also corroborated by our results) that detectors excel (90%+ TPR) in this scenario (dambra2023decoding, ).
+TABLE II: Scenarios for behavioral malware detection.
+SB->SB is the scenario most prior work in our domain considers: both training and testing samples are collected from sample repositories and corpora, which represent the distribution of all samples seen in the wild, and these samples are executed in a sandbox [5, 6, 55, 28].
+It is already known (and also corroborated by our results) that detectors excel (90%+ TPR) in this scenario [26].
 In SB->EP and EP->EP scenarios, detectors are deployed at 3 in Figure 1, i.e., on the endpoint traces of undetected samples.
@@
 This departs from the norm in published research, as most prior detectors have relied on sandbox traces for training.
-We study this scenario to assess the success of a model explicitly trained for endpoint detection.
-These scenarios represent a trade-off between practicality (SB->EP is easiest to implement) and expected detection performance (where EP->EP is favored).
-3.3. Success Metrics for Endpoint Detection
+We study this scenario to assess the success of a model trained explicitly for endpoint detection.
+These scenarios represent a trade-off between practicality (where SB->EP is favored) and detection performance (where EP->EP is favored).
+III-C Success Metrics for Endpoint Detection
 A conventional detector aims to detect as many malware samples as possible, i.e., to maximize the true-positive rate (TPR), while minimizing the number of benign samples misclassified, i.e., the false-positive rate (FPR).
 The defender can control the TPR and FPR by tuning the detection threshold $\tau$; a higher $\tau$ implies a lower TPR and FPR.
-Following prior work (arp2014drebin, ), we evaluate our models regarding their TPR @1% FPR on the test sets (reported as TPR).
+Following prior work [57], we evaluate our models regarding their TPR @1% FPR on the test sets (reported as TPR).
 Additionally, we report the area under the TPR-FPR curve, i.e., the receiver operating characteristic curve, to gauge a model’s overall ability to separate malware and benign classes (reported as AUC).
@@
 Measuring TPR and FPR in SB->SB is straightforward as there is a one-to-one mapping from samples to traces.
-However, in endpoint detection (SB->EP and EP->EP), there is a one-to-many mapping as multiple hosts may execute the same sample at different times, and these resulting traces often have high variability (avllazagaj2021malware, ).
-For example, a few malware samples have over 400 traces, each corresponding to a real-world infection.
-We present a histogram of the number of traces per sample in Figure 6(a).
-To obtain a one-to-one mapping, we consider “an average trace” by selecting one random endpoint trace of each sample and measuring the TPR (averaged over 100 runs).
-This gives us a fair way to compare performance across different scenarios.
+However, in endpoint detection (SB->EP and EP->EP), there is a one-to-many mapping as multiple hosts may execute the same sample at different times, and these resulting traces often have high variability [15].
+For example, a few malware samples have over 400 traces, each corresponding to a real-world infection (see Figure 7(a)).
+To obtain a one-to-one mapping, we consider “a representative trace” by randomly selecting one endpoint trace of each sample to create an EP test set on which we measure a model’s TPR.
+We repeat this process 100 times and report the average TPR over these runs as the final TPR.
+This gives us a fair and unambiguous way to compare performance across scenarios.
 We refine this metric further based on practical considerations.
 67% of traces of an average malware sample (52% for benign) are seen within 24 hours of its first execution at a host.
-As this ratio drops rapidly after the first day (see Figure 6(b)) and malware samples die quickly within days (yong2021inside, ), we only randomly select from the first-day traces—$\mathbf{x}^{0}_{i,(t<24)}$—unless stated otherwise.
-4. Technical Setup
-4.1. Datasets
+As this ratio drops rapidly after the first day (see Figure 7(b)) and malware samples die within days [56], we only randomly select from the first-day traces—$\mathbf{x}^{0}_{i,(t<24)}$—unless stated otherwise.
+IV Technical Setup
+IV-A Datasets
 Training Testing Dataset Mal. Ben. Mal. Ben. EP 0.5K 16.5K 0.4K 8.1K EP (Traces) 19.4K 531.6K 9.7K 412.5K SB1 46.4K 16.7K 31.0K 16.4K SB2 34.5K 7.8K 18.1K 7.3K SB1 $\cap$ SB2 9.9K 5.3K 8.9K 5.5K
-Table 3. A summary of the samples in our datasets.
+TABLE III: A summary of the samples in our datasets.
 Endpoint (EP) Dataset.
 Provided by Avllazagaj et al.
-(avllazagaj2021malware, ), we use a dataset of program traces recorded on real Windows hosts of a commercial anti-malware vendor from over 100 countries between January and July 2018.
+[15], we use a dataset of program traces recorded on real Windows hosts of a commercial anti-malware vendor from over 100 countries between January and July 2018.
 The vendor did not know whether the samples were benign or malicious at the time of execution.
-The samples were executed by the users, who interacted with them naturally, and the vendor’s behavior-based component recorded the traces in a last-ditch effort to discover unknown threats.
-We have reprocessed this dataset specifically for our problem and relabeled it by querying VirusTotal (virustotal, ) for more accurate labels.
-Our processed dataset contains around 1M execution traces from 900 malware samples and 25K benign samples.
+The samples were executed by the users, who interacted with them naturally, and the vendor’s behavioral component recorded the traces in a last-ditch effort to discover unknown threats.
+We have reprocessed this dataset for our problem and relabeled it by querying VirusTotal [48] for more accurate labels.
+Our processed dataset contains around 1M execution traces from 900 malware and 25K benign samples.
 Each trace includes high-level actions (file, registry, process, and mutex actions) of a sample that is executed until its termination.
@@
 Sandbox (SB) Dataset.
-Studying SB->EP realistically requires collecting a sandbox dataset contemporaneous to our EP dataset.
-Because the sandbox dataset will serve as the training set and the EP dataset as the test set, we must collect samples seen in or before 2018 to respect causality (pendlebury2019tesseract, ).
-Moreover, we cannot detonate samples in a sandbox, as it is commonly done (karbab2019maldy, ; jindal2019neurlux, ; miller2016reviewer, ; kuchler2021does, ), as behaviors of old programs today would differ from their original behaviors.
-This is because malware stops functioning when, for example, its remote infrastructure dies (yong2021inside, ) or it starts behaving differently over time (avllazagaj2021malware, ).
-Our solution is collecting traces from VirusTotal (virustotal, ), where third-party sandbox vendors publicly share behavior reports on samples.
-To this end, we curate a list of SHA-256 hashes of Windows samples from popular public malware detection corpora, including EMBER (anderson2018ember, ), and SOREL (harang2020sorel20m, ), released in 2017 or 2018.
+A realistic study of SB->EP requires a sandbox dataset contemporaneous to our EP dataset.
+Because the sandbox dataset will serve as the training set and the EP dataset as the test set, we must collect samples seen in or before 2018 to respect causality [38].
+Moreover, we cannot detonate samples in a sandbox, as it is commonly done [5, 6, 55, 28], as behaviors of old programs today would differ from their original behaviors.
+This is because malware stops functioning when, for example, its remote infrastructure dies [56] or it starts behaving differently over time [15].
+Our solution is collecting traces from VirusTotal [48], where third-party sandbox vendors publicly share behavior reports on samples.
+To this end, we curate a list of SHA-256 hashes of Windows samples from popular public malware detection corpora, including EMBER [49], and SOREL [53], released in 2017 or 2018.
 We then query VirusTotal with these hashes to collect their sandbox traces when available.
 We discard traces that came more than six months after the sample was first seen to capture close-to-original behaviors.
-This process results in traces from two sandbox vendors for around 110K malware and 40K benign samples.
+This results in traces from two sandbox vendors for around 110K malware and 40K benign samples.
 These vendors are well-known in their countries of origin (SB1 is from China, and SB2 is from the USA).
 There is a class imbalance as vendors on VirusTotal are more inclined to share traces of malware samples.
-Rank Test Set Trained on SB1 Trained on SB2 Trained on EP SB1->SB1 SB1->SB2 SB1->EP SB2->SB1 SB2->SB2 SB2->EP EP->EP NGR HYB ATT NGR HYB ATT NGR HYB ATT NGR HYB ATT NGR HYB ATT NGR HYB ATT NGR HYB ATT SB1 95.0 94.2 93.2 60.7 53.7 56.7 11.2 8.0 4.5 48.5 31.8 32.7 85.5 59.2 83.7 14.2 7.2 12.3 43.5 34.0 25.7 SB2 94.0 93.1 92.2 70.4 61.2 64.3 13.9 7.1 5.5 17.6 10.0 14.7 92.6 91.4 90.8 7.5 6.3 11.4 43.4 37.6 26.4 EP 93.2 93.1 87.8 63.5 48.0 52.9 16.7 10.2 8.3 34.5 16.6 15.7 79.3 79.5 88.7 17.3 10.9 13.1 49.5 43.5 42.8
-Table 4.
-The performance (TPR%) of three ML approaches (NGR, HYB, ATT) in seven detection scenarios (based on Table 2).
-In each row, the models are ranked based on the test set, and the average TPR of top-20 models is reported (see Table 15 for AUC).
+Sel.
+Test Set Trained on SB1 Trained on SB2 Trained on EP SB1->SB1 SB1->SB2 SB1->EP SB2->SB1 SB2->SB2 SB2->EP EP->EP NGR HYB ATT NGR HYB ATT NGR HYB ATT NGR HYB ATT NGR HYB ATT NGR HYB ATT NGR HYB ATT SB1 95.0 94.2 93.2 60.7 53.7 56.7 11.2 8.0 4.5 48.5 31.8 32.7 85.5 59.2 83.7 14.2 7.2 12.3 43.5 34.0 25.7 SB2 94.0 93.1 92.2 70.4 61.2 64.3 13.9 7.1 5.5 17.6 10.0 14.7 92.6 91.4 90.8 7.5 6.3 11.4 43.4 37.6 26.4 EP 93.2 93.1 87.8 63.5 48.0 52.9 16.7 10.2 8.3 34.5 16.6 15.7 79.3 79.5 88.7 17.3 10.9 13.1 49.5 43.5 42.8
+TABLE IV: The performance (TPR%) of three ML approaches (NGR, HYB, ATT) in seven detection scenarios (based on Table II).
+The models are selected using the test set in each row, and the average TPR of the top 20 models is reported.
 Trace Standardization.
-Our EP traces contain high-level information that the vendor can seamlessly collect and analyze at endpoint hosts, such as file, process, and mutex creations, as well as registry key creations and deletions.
-Sandboxes also record lower-level information (such as memory dumps (sandboxmemory, )) as they are less computationally constrained.
-Moreover, our traces from three sources (SB1, SB2, and EP) have different formats and conventions.
-This prompts us to implement routines to keep, in our sandbox traces, only the type of information available in EP traces and covert all traces to a standardized format (see Appendix A for details).
-As shown in Section 5, the information remaining in sandbox traces is still enough to obtain performance comparable to prior work ($\sim$95TPR and 99% AUC) in the SB->SB setting.
-This standardization is also critical for our leaderboard design (Section 8).
+Our traces from three sources (SB1, SB2, and EP) have different formats, contents, and conventions.
+We first pre-process our sandbox traces to keep only the type of behaviors recorded in our EP traces, such as file, process, mutex, and registry key creations.
+We then convert all traces to a single standardized format (see Section A-A for details).
+As shown in Section V, the resulting sandbox traces allow us to train models with comparable performance to prior work ($\sim$95TPR and 99% AUC) in the SB->SB scenario.
 Temporal Splitting and Labeling.
-We split our datasets into two portions based on the timestamps of the samples (the first-seen dates) to ensure that the train and test samples are temporally disjoint, avoiding a common pitfall in prior work (pendlebury2019tesseract, ; arp2020and, ).
-Samples seen before April 1st, 2018, along with their traces, are in the Training portion, and the samples seen after that date are in the Testing portion, on which we never train.
-As labels of samples are known to change mildly over time (labeldynamics, ), we make the best effort to assign the training samples historical labels that were available before April 1st, 2018.
+We split our datasets into two portions based on the timestamps of the samples (the first-seen dates) to ensure that the train and test samples are temporally disjoint, avoiding a common pitfall in prior work [38, 42].
+Samples seen before April 1st, 2018, and their traces are in the Training portion, and the samples seen after that are in the Testing portion, on which we never train.
+As labels of samples are known to change mildly over time [58], we make the best effort to assign the training samples historical labels that were available before April 1st, 2018.
 We could find such historical labels for $\sim$24and 100% of the samples in the EP and SB training portions, respectively.
-For testing samples, we query VirusTotal to obtain the most recent detection reports and label samples detected by over five anti-malware engines as malware, following the advice in (labeldynamics, ).
-Further, we use AVClass2 (sebastian2020avclass2, ) to assign family labels to our malware samples based on their detection reports.
+For testing samples, we query VirusTotal to obtain the most recent detection reports and label samples detected by over five anti-malware engines as malware, following the advice in [58].
+Further, we use AVClass2 [59] to assign family labels to our malware samples based on their detection reports.
 We tag malware samples that did not receive a family label as Generic malware.
-Overall, this methodology ensures that our evaluation is realistic and reflects the conditions the anti-malware vendor had to work under when our endpoint data was collected.
-4.2. Machine Learning Details
-We experiment with three ML approaches: (i) a bag-of-n-grams-based model (NGR), (ii) a hybrid model that combines convolution and attention (HYB), and (iii) a self-attention-based sequence model (ATT).
-These three approaches cover a wide range of designs proposed by prior work in behavior-based malware detection with ML, respectively, MalDy (karbab2019maldy, ), Neurlux (jindal2019neurlux, ) and, most recently, Nebula (trizna2023nebula, ).
-Please find the details on these models in Appendix B.
-We address the class imbalance in our training sets by oversampling the underrepresented classes (malware in EP and benign in SB datasets) (chawla2002smote, ).
-Note that our goal is not to develop a better model architecture or feature processing routine but to evaluate the effectiveness of state-of-the-art architectures at endpoint malware detection.
-We make our ML improvements without changing the fundamentals of these existing approaches.
+This methodology ensures that our evaluation is realistic and reflects the conditions the anti-malware vendor had to work under when our endpoint data was collected.
+IV-B Machine Learning Details
+We experiment with three deep-learning-based approaches: (i) a bag-of-n-grams-based model (NGR), (ii) a hybrid model that combines convolution and attention (HYB), and (iii) a self-attention-based sequence model (ATT).
+These approaches cover a wide range of designs proposed by prior work in behavioral malware detection with ML, respectively, MalDy [5], Neurlux [6] and, most recently, Nebula [7] (see Appendix B for details on these models).
+We address the class imbalance in our training sets by oversampling the underrepresented classes (malware in EP and benign in SB datasets).
+Note that our goal is not to develop a new architecture or feature processing routine but to evaluate the state-of-the-art ones at endpoint malware detection.
+We make our ML improvements without changing the fundamentals of existing approaches.
 […]
-5. Endpoint Detection Evaluations
-Table 4 presents the endpoint detection performance of ML models (based on the methodology in Section 4.2), using three different data sets (SB1, SB2, and EP) for training and testing, following the scenarios in Table 2.
+V Endpoint Detection Evaluations
+Table IV presents the detection performance of ML models (based on the methodology in Section IV-B), using three different data sets (SB1, SB2, and EP) for training and testing, following the scenarios in Table II.
 Here, for example, SB1->SB2 and SB1->EP refer to the scenarios where the models are trained on the SB1 training set and evaluated on the SB2 and EP testing sets, respectively.
-5.1. Evaluating the SB->EP Scenario
-First, we see significant gaps between SB->SB and SB->EP scenarios across the board.
-This gap is the largest when models are trained and ranked on the same sandbox, e.g., 95.0% vs.
-11.2% for NGR trained and ranked on SB1.
-When models are ranked on a different sandbox, it often shrinks, e.g., 94.0% vs.
-13.9% for NGR trained on SB1 and ranked on SB2.
-Nevertheless, the gap is still substantial even when we rank the models on their EP performance, e.g., 93.2% vs.
-16.7% for NGR trained on SB1 and ranked on EP.
-Although there is a gap between the performances on different sandboxes, e.g., 94.0% vs.
-70.4% for NGR trained on SB1 and ranked on SB2, it is significantly smaller than the SB->EP gap.
+See Table XIII for the AUC results.
+V-A Evaluating the SB->EP Scenario
+We see wide performance gaps between SB->SB and SB->EP scenarios.
+The gap widens when models are trained and selected using the same sandbox, e.g., 95.0% vs.
+11.2% for NGR trained and selected using SB1.
+When models are selected using a different sandbox, the gap shrinks, e.g., 94.0% vs.
+13.9% for NGR trained on SB1 and selected using SB2.
+The gap is still substantial even when we select the models based on their EP performance: e.g., 93.2% vs.
+16.7% for NGR trained on SB1 and selected using EP.
+Although there is a gap between a model’s performances on different sandboxes, e.g., SB1->SB1 vs.
+SB1->SB2, it is smaller than the SB->EP gap.
 Overall, this gap persists regardless of the training sandbox or the model type, hinting at the limitations of sandbox-trained models for endpoint detection.
-Notably, this gap is larger for models that perform better on the test traces from a sandbox that also produced the training set.
-In consequence, studies that rely on a single sandbox for evaluation in SB->SB (miller2016reviewer, ; karbab2019maldy, ; kuchler2021does, ; trizna2023nebula, ) are at a higher risk of producing worse models for SB->EP.
-Trained on SB1 Trained on SB2 SB1-SB2 SB1-EP SB2-EP SB1-SB2 SB1-EP SB2-EP NGR 0.47 0.03 0.45 --0.24 0.39 --0.46 HYB 0.88 0.84 0.75 --0.17 0.59 0.33 ATT 0.56 --0.17 0.31 0.51 0.58 0.35
-Table 5. Ranking correlations of sandbox-trained models according to their performances (TPR) on different test sets.
-Next, we verify the possibility of finding a better model for SB->EP by ranking the models using traces from an unseen sandbox (not used for training).
-We rank our models (100+ in each setting) based on their SB1, SB2, and EP performances (TPR) and compute Spearman’s correlation coefficients between these rankings.
-Table 5 reveals that (except for one setting) (i) the rankings based on the training sandbox correlate poorly (sometimes negatively) with EP rankings, (ii) rankings based on an unseen sandbox correlate more strongly with EP rankings.
-For example, for ATT trained on SB1, the rankings based on SB1 and SB2 have $0.03$ and $0.45$ correlation with the rankings based on EP, respectively.
+We observed that the gap widens when model training and selection use traces from the same sandbox.
+In Appendix C, we perform an experiment showing this is a generally valid observation.
 We believe this is because a model that performs better on the traces from an unseen sandbox is less likely to have overfitted to features specific to the training sandbox.
-We discuss such features in more detail in Section 7.
-As a result, to get higher performance in SB->EP, we recommend performing model selection using traces from a different, unseen sandbox.
-Finally, we experiment with combining traces from two sandboxes (SB1 and SB2) for training.
-However, this brings only minor improvements for SB->EP over training on traces from a single sandbox.
-When we use EP traces to rank the models, training on both sandboxes achieves at most 17.5% TPR on the EP test set for NGR (11.7% for HYB, and 13.9% for ATT).
-In Section 7, we show that the diversity between the traces of a sample from two sandboxes is much lower than between its traces from two endpoint hosts.
+We discuss such features in more detail in Section VII.
+This implies that studies that rely on a single sandbox for training and evaluation in SB->SB [55, 5, 28, 7] are at a higher risk of producing worse models for SB->EP.
+Consequently, we recommend performing model selection in SB->EP using traces from a different, unseen sandbox.
+Finally, we explore combining traces from two sandboxes (SB1 and SB2) for training.
+However, this brings only minor improvements for SB->EP over training on traces from one sandbox.
+When we use EP traces for model selection, training on both sandboxes achieves at most 17.5% TPR on the EP test set for NGR (11.7% for HYB, and 13.9% for ATT).
+In Section VII, we show that the diversity between the traces of a sample from two sandboxes is much lower than between its traces from two endpoint hosts.
 This lack of diversity makes combining SB1 and SB2 traces ineffective for increasing the EP performance.
+The Impact of Trace Standardization.
+The standardized trace format (see Section IV-A) removes features from sandbox traces unavailable in endpoint traces.
+In Section A-B, to understand whether this might hurt the SB->EP performance, we experiment with training and testing models on unstandardized sandbox traces.
+This shows that standardization moderately hurts the in-domain performance (test traces from the training sandbox) while improving generalization to other domains (traces from other sandboxes).
+Consequently, we believe standardization helps with improving out-of-domain performance in SB->EP.
 Applying Drift Detection.
 We study the low performance in SB->EP through the lens of distribution shift, which occurs if the process generating inputs, i.e., program traces, changes between the training and testing phases.
-Concept drift is a highly-studied type of distribution shift in malware detection that stems from malware evolution (pendlebury2019tesseract, ).
-In past work on malware concept drift, the distribution shift builds up over time in concept drift, e.g., over 50% performance drop in 2 years (barbero2022transcending, ).
-In contrast, we find that in SB->EP the shift manifests immediately after the model is deployed.
-One way to combat malware concept drift is to use a drift detector (DD) to identify samples that might have drifted and abstain from classifying them (jordaney2017transcend, ; barbero2022transcending, ; yang2021cade, ; chen2023continuous, ).
-DDs aim to prevent mistakes by rejecting out-of-distribution (OOD) samples that diverge from the training distribution while accepting in-distribution (ID) samples on which predictions are reliable (jordaney2017transcend, ; yang2021cade, ).
+Concept drift is a highly-studied type of distribution shift in malware detection that stems from malware evolution [38].
+In past work on malware concept drift, the distribution shift builds up over time in concept drift, e.g., over 50% performance drop in 2 years [25].
+In contrast, we find that in SB->EP, the shift manifests immediately after the model is deployed.
+One way to combat malware concept drift is to use a drift detector (DD) to identify samples that might have drifted from the training distribution [24, 25, 60, 39].
+The classifier then rejects such out-of-distribution (OOD) samples while accepting the in-distribution (ID) ones on which predictions are reliable [24, 60].
 We apply drift detection to quantify the distribution shift present in SB->EP.
-We employ a modern technique ReAct (sun2021react, ) to assign an outlierness score to each sample, among which the highest-scoring ones are rejected.
-ReAct rectifies a model’s penultimate layer activations, which is shown to make the prediction probabilities of a model a more calibrated indicator of outlierness.
-We selected ReAact due to its flexible approach that is agnostic to the underlying model as long as it’s a neural network.
-Figure 2. The rejection rates resulting from drift detection.
+We employ a recent technique ReAct [61] to assign an outlierness score to each sample, among which the highest-scoring ones are rejected.
+ReAct rectifies a model’s penultimate layer activations, which calibrates the prediction probabilities of a model to be a better indicator of outlierness.
+We selected ReAact due to its flexible approach, which can be applied to any neural network model.
+We experiment with a NGR model trained on the SB1 training set.
+We tune ReAct to reject $K\in[0,15]$of the traces in the SB1 test set (i.e., the ID rejection rate) and separately measure for EP and SB2 test sets (i) corresponding rejection rates on the malware and benign traces; and (ii) the TPR on the accepted traces.
+Table V presents the results.
 […]
-Finally, in Table 6, we present the TPR on the accepted (remaining) traces in EP and SB2 for increasing values of $K$.
-Rejecting the traces in EP or SB2 increases the performance, but this increase is less pronounced for EP, e.g., gaining $\sim$10TPR (over the baselines in Table 4) takes rejecting 67% and 22% of malware traces in EP and SB2, respectively.
+Regarding (ii), classification with rejection increases the performance for both EP and SB2 test sets.
+However, this increase is less pronounced for EP, e.g., gaining $\sim$10TPR (over the baselines in Table IV) takes rejecting 67% and 22% of malware traces in EP and SB2, respectively.
 This also distinguishes the distribution shift in SB->EP from concept drift; rejection rates must be high for meaningful performance gains due to the magnitude of the shift.
-($K$) ID RejSB1->EP SB1->SB2 TPR MalBenTPR MalBen5% 20.2% 31.1% 20.5% 79.2% 22.0% 21.5% 10% 23.5% 53.2% 39.0% 89.4% 38.2% 36.5% 15% 25.6% 67.7% 46.4% 88.3% 52.4% 53.6% 20% 42.6% 72.2% 49.0% 87.7% 60.8% 60.6% 25% 53.1% 78.5% 52.5% 82.9% 68.5% 70.8%
-Table 6. The TPR of an NGR model in SB1->SB2 and SB1->EP scenarios, with increasing ID rejection rates, and the corresponding rejection rates of malware and benign traces.
-5.2. Evaluating the EP->EP Scenario
+($K$) ID RejSB1->EP SB1->SB2 TPR MalBenTPR MalBen5% 20.2% 31.1% 20.5% 79.2% 22.0% 21.5% 10% 23.5% 53.2% 39.0% 89.4% 38.2% 36.5% 15% 25.6% 67.7% 46.4% 88.3% 52.4% 53.6%
+TABLE V: The TPR of an NGR model in SB1->SB2 and SB1->EP scenarios, with increasing ID rejection rates and the corresponding rejection rates of malw. and benign traces.
+V-B Evaluating the EP->EP Scenario
 For our models in EP->EP, we select up to 4 EP traces per sample to form our training set.
-As we will show in Section 7, the behavior variability among the real-world traces of a sample (avllazagaj2021malware, ) impacts the model’s predictions.
+As we will show in Section VII, the behavior variability among the real-world traces of a sample [15] impacts the model’s predictions.
 Consequently, including more traces per sample for training provides better coverage of the behaviors in the wild.
 We find that using more than 4 traces per sample provides diminishing returns, e.g., for 1, 2, 4, and 16 traces per sample, the TPR of NGR is 43.0%, 44.8%, 49.5%, and 49.9%, respectively.
-In Table 4, we see a much higher EP performance in EP->EP than in SB->EP, as expected.
+In Table IV, we see a much higher EP performance in EP->EP than in SB->EP, as expected.
 Although training and testing sets are from similar distributions in EP->EP (as opposed to SB->EP), the performance still lags behind the SB->SB performance, e.g., 95% vs.
 50% TPR and 99% vs.
-88% AUC (Table 15).
+88% AUC (Table XIII).
 In the following sections, we shed light on this discrepancy and what makes endpoint detection particularly more challenging than a sandbox-only scenario.
-6. Endpoint Detection Deals With Difficult-To-Classify Samples
-As described in Section 3.1, an endpoint detector operates on samples that cannot be classified statically—a subset of all samples that exist in the wild.
+Takeaways: Distribution shift causes ML methods trained on sandbox traces to perform poorly on endpoint traces (70+TPR drop).
+Even when trained on endpoint traces, these methods’ endpoint performance still cannot reach their sandbox-based performance ($\sim$20vs.
+95% TPR), hinting at the inherent challenges in endpoint detection.
+VI Endpoint Detection Deals With Difficult-To-Classify Samples
+As described in Section III-A, an endpoint detector operates on samples that cannot be classified statically—a subset of all samples in the wild.
 The filtering effect this introduces on the distribution of samples relevant to endpoint detection has not been systematically measured before.
-To illustrate, we explore (based on Table 13) the top malware families in our EP test set (undetected samples) and the SB test set (samples from public corpora) Although both datasets include only the samples first seen in 2017-2018, they encode vastly different priors over samples.
+To illustrate, we examine (based on Table XIV) the top malware families in our EP test set (undetected samples) and the SB test set (the samples from public corpora).
+Although both datasets include only the samples first seen in 2017-2018, they encode different priors over samples.
 For example, generic malware (i.e., malware that could not be placed into any known family) covers 23.8% of the EP samples vs.
 4.8% of the SB samples.
-In the EP set, there are families such as Khalesi and Emotet that were rampant in 2018 (emotet, ).
+In the EP set, there are families such as Khalesi and Emotet that were rampant in 2018 [62].
 Conversely, in the SB set, we see older families such as Sivis and Upatre (circa 2013-2014), whose variants still spread to this day.
-Moreover, in Table 13, we share top benign sample publishers (extracted from the code-signing certificates when available), indicating a similar distribution difference in benign samples.
-Overall, the typical distribution of training and testing samples used by prior work in SB->SB (karbab2019maldy, ; jindal2019neurlux, ; dambra2023decoding, ) is substantially different than the distribution of testing samples in SB->EP and EP->EP.
+Moreover, in Table XIV, we share top benign sample publishers (extracted from the code-signing certificates when available), indicating a similar distribution difference in benign samples.
+Overall, the typical distribution of training and testing samples used by prior work in SB->SB [5, 6, 26] is substantially different than the distribution of testing samples in SB->EP and EP->EP.
 Borderline Samples and Label Noise.
-A common practice in related work in malware detection with ML is discarding borderline samples on which the anti-malware engines deployed by VirusTotal are not in consensus.
-Different works use different criteria; for example, they discard a sample if it is detected by less than 20 engines (mantovani2020prevalence, ) or five engines (kuchler2021does, ), or by more than 0 and less than 40 engines (lucasadversarial, ).
-Even for works that do not explicitly discard borderline samples, their data collection results in few of them.
-For example, recent related work has suggested that dynamic features have limited benefits over static features (dambra2023decoding, ).
+A common practice in related work in malware detection with ML is discarding borderline samples on which the anti-malware engines on VirusTotal are not in consensus.
+Different works use different criteria; for example, they discard a sample if it is detected by more than zero and less than five engines [28], 20 engines [63] or 40 engines [64].
+Even works that do not explicitly discard borderline samples may follow a biased data collection process.
+For example, recent related work has suggested that dynamic features have limited benefits over static features [26].
 The benign samples used in this work were from a trusted Windows software repository, and 93% of their malware samples were detected by 20+ engines.
-Based on prior definitions, we consider samples detected by more than 0 and less than 20 engines borderline.
+We consider the samples detected by more than zero and less than 20 engines borderline.
 This results in 27.7% borderline samples in our EP test set (vs.
@@
 Next, we will show that this artificially inflates the measured success of a model.
-As described in Section 4.1, we do not discard any samples and label a sample as malware if more than five engines detect it.
+As described in Section IV-A, we do not discard any samples and label a sample as malware if more than five engines detect it.
 This threshold poses a trade-off for the ground truth: as it increases, more malware samples may be labeled as benign, and as it decreases, more benign samples may be labeled as malware.
-We present a case study in Section 6.2 to demonstrate how this trade-off plays out.
-6.1. The Impact of Sample Distributions
-Here, we measure how the distribution of samples in the test set changes a model’s measured performance.
+We present a case study in Section VI-B to demonstrate how this trade-off plays out.
+VI-A The Impact of Sample Distributions
+Here, we measure how the distribution of samples in the test set changes a model’s measured (and perceived) performance.
 To this end, we resample our test sets according to some criteria, e.g., following the malware family distribution in the EP test set, and evaluate our models on these new test sets.
 This simulates different distributions over a model’s test samples and allows us to do controlled experiments.
-We focus on NGR models due to their superior EP performance in Section 5.
-EX# SB1->SB1 SB1->EP EP->EP
-TPR AUC TPR AUC TPR AUC
-EX0 No Resampling - Orig. Distributions
-93.2 99.0 16.7 78.4 49.5 87.5
-EX1 Orig. Distributions w/o Generic Malware
-95.6 99.5 18.6 81.4 56.8 91.4
-EX2 Only Generic Malware
-38.4 85.8 9.8 67.8 23.9 74.2
-EX3 Discard Borderline Samples
-96.7 99.5 22.6 83.9 66.9 94.2
-EX4 Malw. Resampled Following EP Test (w/ Generic)
-63.2 94.2 16.7 78.4 49.5 87.5
-EX5 Malw. Resampled Following EP Test (w/o Generic)
-72.1 97.2 18.6 81.4 56.8 91.4
-Table 7. The impact of sample distributions in the test set on model evaluations. Each experiment is denoted as EX0-5.
-Table 7 presents the results of our resampling experiments.
+We focus on NGR models due to their superior EP performance in Section V.
+EX# SB1->SB1 SB1->EP EP->EP TPR AUC TPR AUC TPR AUC EX0 No Resampling - Orig.
+Distributions 93.2 99.0 16.7 78.4 49.5 87.5 EX1 Orig.
+Distributions w/o Generic Malware 95.6 99.5 18.6 81.4 56.8 91.4 EX2 Only Generic Malware 38.4 85.8 9.8 67.8 23.9 74.2 EX3 Discard Borderline Samples 96.7 99.5 22.6 83.9 66.9 94.2 EX4 Malw.
+Resampled Following EP Test (w/ Generic) 63.2 94.2 16.7 78.4 49.5 87.5 EX5 Malw.
+Resampled Following EP Test (w/o Generic) 72.1 97.2 18.6 81.4 56.8 91.4
+TABLE VI: The impact of sample distributions in the test set on model evaluations. EX0-5 denote different experiments.
+Table VI presents the results of our resampling experiments.
 We resample the test sets according to five criteria—denoted as EX0--5 where EX0 is testing on the original test set of each respective scenario.
 In EX1, leaving out generic malware samples from original test sets leads to a significant boost in performance, whereas keeping only generic malware in EX2 causes a massive drop.
-This aligns with prior claims that generic malware is at the borderline and much harder to classify (ugarte2019close, ).
+This aligns with prior claims that generic malware is at the borderline and harder to classify [54].
 The fact that Generic has 5$\sim$ 5$\times$ 5 × higher coverage in the EP test set than in the SB test set (23.8% vs.
 4.8%) also demonstrates the filtering effect that funnels more difficult samples to the endpoint detector.
-Next, in EX3, we discard the borderline samples (identified in the previous section) from all test sets, simulating a common prior practice (kuchler2021does, ; mantovani2020prevalence, ; dambra2023decoding, ).
+Next, in EX3, we discard the borderline samples (identified in the previous section) from all test sets, simulating a common prior practice [28, 63, 26].
 This yields a significant boost in all scenarios: 5.9% TPR boost in SB1->EP and 17.4% in EP->EP.
@@
 Note that, for SB1->EP and EP->EP, EX4 is equivalent to EX0 and EX5 is equivalent to EX1.
-We observe a great decrease in the original SB1->SB1 performance, e.g., from 93.2% TPR to 63.2% in EX4 and to 72.1% in EX5.
-These results suggest that the difference in distribution of malware families explains some of the performance drop when applying these classifiers to EP traces, perhaps because EP traces contain more harder-to-classify malware families, but there remains a large gap that is not explained by differences in the distribution of families.
+We observe a large decrease in the original SB1->SB1 performance, e.g., from 93.2% TPR to 63.2% in EX4 and to 72.1% in EX5.
 Despite neither including any generic malware, the SB1->SB1 performance in EX5 is significantly lower than EX1.
-This suggests that non-generic malware samples relevant to endpoint detection are generally more difficult to classify correctly, whether based on their sandbox or endpoint traces.
-Overall, these experiments reveal how the performance of prior behavior-based detectors may be over-estimated because they are evaluated on samples that are not representative of the distribution of samples faced in endpoint detection.
-6.2. Case Studies
-Following our measurements, we present two case studies illustrating the sample distribution challenges in endpoint detection.
+This suggests that non-generic malware samples relevant to endpoint detection are generally harder to classify correctly, whether based on their sandbox or endpoint traces.
+Consequently, the performance of prior behavioral detectors may be overestimated as they are evaluated on samples that are not representative of the distribution of samples faced in endpoint detection.
+Although the distribution of malware samples explains most of the performance loss when ML classifiers are applied to EP traces, an unexplained gap remains (63.2% vs.
+49.5%), which we attempt to explain in Section VII.
+VI-B Case Studies
+Next, we present two case studies illustrating sample distribution challenges in endpoint malware detection.
 Difficult Benign Samples.
-Table 13 shows that Microsoft is the top benign publisher in our EP test set, covering 7.8% of all benign samples.
+Table XIV lists Microsoft as the top benign publisher in our EP test set, covering 7.8% of all benign samples.
 This is counterintuitive as Microsoft is a trusted publisher, and our EP dataset only records samples that could not be classified statically as malware or benign.
-We discovered that almost all (98%) Microsoft samples in our EP test set share the AM_Delta prefix in their filenames, corresponding to periodic patches to Windows Defender Anti-Malware.
+We discovered that almost all (98%) Microsoft samples in our EP test set share the AM_Delta prefix in their filenames, which correspond to periodic patches to Windows Defender.
 The NGR model in EP->EP outputs an average score of 37% on AM_Delta endpoint traces, almost 2.5$\times$ of the average score across all benign traces.
@@
 In EP->EP, on a test set containing only AM_Delta samples as the benign samples, the model achieves only 62.6% AUC, compared to 87.5% with all benign samples.
-Moreover, there are online reports (amdeltareport, ) about anti-malware software falsely flagging AM_Delta files.
-A popular configuration repository for Sysmon—a tool to monitor process activities—includes an allow-list rule specifically for AM_Delta (amdelta_syzmon, ).
-Our evidence suggests that AM_Delta samples are difficult to classify correctly due to making sensitive modifications on hosts.
-Conversely, our SB test set contains only two AM_Delta samples (less than 0.01% of all benign samples).
-This highlights that an endpoint detector often faces difficult—false positive prone—benign samples in addition to difficult malware families, which is a root cause behind lower performance in SB->EP and EP->EP.
+Moreover, there are reports [65] about anti-malware software falsely flagging AM_Delta files.
+A popular configuration repository for Sysmon—a tool to monitor process activities—includes an allow-list rule specifically for AM_Delta [66].
+These suggest that AM_Delta samples are difficult to classify correctly as they make sensitive modifications on hosts.
+In contrast, our SB test set contains only two AM_Delta samples (less than 0.01% of all benign samples).
+This highlights that an endpoint detector often faces difficult—false positive prone—benign samples in addition to difficult malware families, which is a root cause behind poor performance in SB->EP and EP->EP.
 Label Noise in Borderline Samples.
@@
 This demonstrates how borderline samples such as Roblox can introduce label noise, particularly in endpoint detection, where they are more common.
-Before After NGR HYB NGR HYB TPR AUC TPR AUC TPR AUC TPR AUC Soft Labeling (EP->EP) 49.5 87.5 43.5 86.8 50.6 87.8 49.2 88.2 Resampled Training Set — Uniform (SB1->EP) 16.7 78.4 10.2 74.1 15.9 77.2 8.6 74.0 Resampled Training Set — EP Training (SB1->EP) 16.7 78.4 10.2 74.1 20.0 78.6 11.4 74.9
-Table 8. Our strategies to improve the performance against sample distribution challenges in endpoint detection.
-6.3. Combating Sample Distribution Challenges
-Building an intuition from our previous experiments, we propose two strategies to improve the performance in SB->EP and EP->EP.
+VI-C Combating Sample Distribution Challenges
+Building on our observations, we propose two strategies to improve the performance in SB->EP and EP->EP.
 Soft-Labeling Against Label Noise in EP->EP.
 Malware detectors are generally trained using hard binary labels—0 for benign and 1 for malware.
-However, hard labels are hazardous when noisy as they force the model to overfit to a wrong prediction (li2019learning, ).
+However, hard labels are hazardous when noisy as they force the model to overfit to a wrong prediction [67].
 The high frequency of borderline, i.e., potentially noisy, samples exacerbates this problem in endpoint detection.
-Research suggests that soft labels can be effective against label noise (lukasik2020does, ) by preventing the model from getting too confident on noisy labels during training.
-Hoping to improve the EP->EP performance, we implement a function to assign soft labels to our borderline training samples.
+Research suggests that soft labels can be effective against label noise [29] by preventing the model from getting too confident on noisy labels during training.
+To improve the EP->EP performance, we implement a function to assign soft labels to our borderline training samples.
 […]
+Before After NGR HYB NGR HYB TPR AUC TPR AUC TPR AUC TPR AUC Soft Labeling (EP->EP) 49.5 87.5 43.5 86.8 50.6 87.8 49.2 88.2 Resampled Training Set — Uniform (SB1->EP) 16.7 78.4 10.2 74.1 15.9 77.2 8.6 74.0 Resampled Training Set — EP Training (SB1->EP) 16.7 78.4 10.2 74.1 20.0 78.6 11.4 74.9
+TABLE VII: Our strategies to improve the performance against sample distribution challenges in endpoint detection.
 More Accurate Training Distributions in SB->EP.
 Our models in SB->EP were trained on the distribution of samples captured in public corpora.
-Although this is standard for obtaining the best SB->SB performance (jindal2019neurlux, ; trizna2023nebula, ), it is not ideal for SB->EP where the test set follows a different distribution.
-ML models are heavily influenced by the priors encoded in their training sets due to the empirical risk minimization (ERM) paradigm, which minimizes the average loss over all training samples.
-For example, 10% of the malware in the SB training set is from the Virut family, and, therefore, that family will be prioritized over Emotet, which accounts for 0.3%.
-We first attempt to alleviate this discrepancy in SB->EP by resampling the SB training set to include an equal number of samples from each family.
-Unfortunately, as presented in Table 8 (second segment), this uniform sampling strategy ends up hurting the performance in SB->EP.
-Considering over 700 malware families are in the SB training set (vs.
-around 100 in the EP training set), families relevant to endpoint detection remain a minority in the new family-balanced training set.
-This explains the low performance and raises doubts about whether such an uninformed strategy could work.
+Although this is standard for obtaining the best SB->SB performance [6, 7], it is not ideal for SB->EP where the test set follows a different distribution.
+ML models are heavily influenced by the priors encoded in their training sets because the empirical risk minimization (ERM) paradigm minimizes the average loss over all training samples.
+For example, 10% of the malware in the SB training set is from the Virut family, which will be prioritized over Emotet, which accounts for 0.3%.
+To alleviate this discrepancy in SB->EP, we turn our attention to ML research, where group robustness techniques [31, 69, 70] have been effective against a similar problem.
+These techniques aim to train models that perform well on each subgroup (e.g., demographic groups) instead of favoring groups over-represented in the training set.
+We implement a prior group robustness method [71] by creating an SB training set with an equal number of samples from each malware family.
+Unfortunately, as presented in Table VII (second segment), this uniform sampling strategy ends up hurting the performance in SB->EP.
+We attribute this shortcoming to the large number of groups (families) in our problem, compared to standardized ML benchmarks, over 600 vs.
+typically less than 10 [69, 72].
+Moreover, malware family labels are often incomplete or noisy [59, 73].
+For example, in our SB1 training set, there are many noisy malware family tags such as vbkryjetor (13 samples).
+As a result, the model is forced to balance hundreds of groups, which also might be inaccurately tagged.
+These results highlight a key difference between real-world security applications and simplified ML benchmarks, on which most solutions are developed and evaluated, raising doubts about whether an uninformed training strategy could work.
 Finally, we experiment with an informed strategy where the SB->EP training set is resampled to follow the family distribution in the EP training set.
-Widely used threat intelligence platforms can guide practitioners about which families to include in training (ciscothreatintel, ).
-Table 8 (third segment) shows that this strategy brings moderate performance improvements: 3.3% and 1.2% higher TPR for NGR and HYB, respectively.
-We will continue using this strategy to train models in the rest of our experiments.
-We leave further improvements to future work, such as handling notably noisy family labels (sebastian2020avclass2, ; kotzias2019mind, ).
-7. Environment-Sensitive Program Behaviors Hurt Endpoint Detection
+Widely used threat intelligence platforms can guide practitioners about which families to include in training [74].
+Table VII (third segment) shows that this strategy brings moderate improvements: 3.3% and 1.2% higher TPR for NGR and HYB, respectively.
+We will use this strategy to train models in the rest of our paper.
+Takeaways: The distribution of samples selected evaluate malware detectors must account for the detector’s position in a pipeline.
+Accordingly, creating realistic test sets for ML-based behavioral detectors (i.e., the final step in the pipeline) reveals the overestimated performance in prior evaluations done on broad distributions (63% vs.
+93% TPR).
+VII Environment-Sensitive Program Behaviors Hurt Endpoint Detection
 We have shown the performance implications of the distribution of samples encountered by endpoint detectors.
-However, even when we account for distribution differences (EX4 in Table 7), the performance in SB->SB (63.2% TPR) is higher than the performance in SB->EP (16.7% TPR) and in EP->EP (49.5% TPR).
-In this section, we aim to illuminate the remaining performance gap between by studying the impact of variable program behaviors.
-It is known that program behaviors are environment sensitive and, as a result, a sample can exhibit varying behaviors in different environments (balzarotti2010efficient, ; lindorfer2011detecting, ; avllazagaj2021malware, ).
-Moreover, malware samples are more environment-sensitive than benign samples because, for example, they execute only when specific vulnerabilities are present in the host (avllazagaj2021malware, ) or do not execute if they detect they are running in a sandbox (lindorfer2011detecting, ; miramirkhani2017spotless, ).
-In this section, we build on these insights to understand the impact of variable program behaviors on endpoint detection performance.
+However, even when we accounted for distribution differences (EX4 in Table VI), the performance in SB->SB (63.2% TPR) is higher than the performance in SB->EP (16.7% TPR) and in EP->EP (49.5% TPR).
+In this section, we aim to illuminate the remaining performance gap by studying the impact of variable program behaviors on ML-based approaches.
+It is known that program behaviors are environment sensitive and, as a result, a sample can exhibit varying behaviors in different environments [16, 34, 15].
+Malware samples are more environment-sensitive than benign samples because, for example, they execute only if the host has specific software vulnerabilities [15] or do not execute if they detect they are running in a sandbox [34, 21].
+Building on these insights, we aim to characterize the impact of variable program behaviors on ML-based detectors.
 We focus on the variability stemming from environment differences between sandboxes and endpoint hosts (for the SB->EP scenario) and between different endpoint hosts (for the EP->EP scenario).
-We perform our experiments using the improved NGR models from Section 6.3.
-7.1. The Diversity Among Endpoint Traces
-Table 9 shows how much traces vary when similar sets of samples (e.g., from the same family) are run in different environments.
-We aggregate all the traces from three popular malware families and two benign publishers from our SB1, SB2, and EP datasets.
-To cover different types of samples, we select Wannacry (ransomware), Emotet (banking trojan), and Khalesi (info-stealer) as the families, and Opera (web browser), Roblox (game platform) as the publishers.
-We then measure the average pairwise distances between these traces within an environment, e.g., between the traces from SB1 (SB1-SB1), and between two environments, e.g., between the traces from SB1 and EP (SB1-EP).
-We use normalized compression distance metric, which has been used in similar contexts (bailey2007automated, ; avllazagaj2021malware, ).
-Within Environment Between Environments SB1-SB1 SB2-SB2 EP-EP SB1-SB2 SB1-EP SB2-EP Wcry 0.16 0.11 0.44 0.28 0.67 0.50 Emot 0.48 0.35 0.50 0.55 0.65 0.53 Khal 0.33 0.29 0.48 0.52 0.62 0.57 Oper 0.05 0.07 0.69 0.43 0.70 0.75 Rblx 0.18 0.04 0.63 0.38 0.67 0.69
-Table 9. Average pairwise distances between traces of samples from three malware families and two benign publishers.
-We observe: (i) traces from the same sandbox are almost always very similar; (ii) traces from two endpoint hosts are dissimilar (also shown in (avllazagaj2021malware, )); (iii) traces from different sandboxes are dissimilar but much less so than one trace from a sandbox and one trace from an endpoint host.
-The observation (iii) is potentially detrimental for SB->EP, as the features learned from SB traces might be irrelevant to EP traces, and (ii) is potentially detrimental for both scenarios.
-The dissimilarity between SB and EP traces implies that the training set of a model in SB->EP follows a different distribution than its test set, which explains the high drift detection in Section 5.1.
+We perform our experiments using the improved NGR models from Section VI-C.
+VII-A The Diversity Among Endpoint Traces
+Table VIII shows how much traces vary when a set of related samples (e.g., from the same malware family) is run in different environments.
+We aggregate all the traces from three malware families and two benign publishers in our SB1, SB2, and EP datasets.
+To cover a range of samples, we select Wannacry (ransomware), Emotet (banking trojan), and Khalesi (info-stealer) as the families, and Opera (web browser), Roblox (game platform) as the publishers.
+We report the average pairwise distances between these traces within an environment, e.g., between the traces from SB1 (SB1-SB1), and between two environments, e.g., between the traces from SB1 and EP (SB1-EP).
+We use normalized compression distance, which has been used in similar contexts [75, 15].
+Within Envs.
+Between Envs.
+SB1-SB1 SB2-SB2 EP-EP SB1-SB2 SB1-EP SB2-EP Wcry 0.16 0.11 0.44 0.28 0.67 0.50 Emot 0.48 0.35 0.50 0.55 0.65 0.53 Khal 0.33 0.29 0.48 0.52 0.62 0.57 Oper 0.05 0.07 0.69 0.43 0.70 0.75 Rblx 0.18 0.04 0.63 0.38 0.67 0.69
+TABLE VIII: Avg. pairwise distances between traces of samples from three malware families and two benign publishers.
+We observe: (i) traces from the same sandbox are almost always very similar; (ii) traces from two endpoint hosts are dissimilar (also shown in [15]); (iii) traces from different sandboxes are dissimilar but much less so than one trace from a sandbox and one trace from an EP host.
+The observation (iii) can be detrimental for SB->EP, as the features learned from SB traces might be irrelevant to EP traces, and (ii) is potentially detrimental for both scenarios.
+The dissimilarity between SB and EP traces implies that a model’s training set in SB->EP follows a different distribution than its test set, explaining the high drift detection in Section V-A.
 Next, we demonstrate how (iii)—the distance between SB traces and EP traces—affects a model in SB1->EP.
@@
 Consequently, the less similar the input EP trace is to the SB traces in the training set, the worse decisions the model makes and performs poorly in SB->EP.
-Figure 3. Relationship between the average distance of an input EP trace to the SB traces and the model’s output score in SB->EP. Each marker represents the avg. of 30 input traces.
+Figure 2: Relationship between the avg. distance of an input EP trace to the SB traces and the model’s output score in SB->EP. Each marker represents the avg. of 30 input traces.
 […]
+Figure 3: Prediction score std. deviations on the EP traces of each sample in the EP test set; in SB1->EP and EP->EP.
 […]
-Figure 4. Prediction score std. deviations on the endpoint traces of each sample in the EP test set; in SB1->EP and EP->EP.
-7.2. Case Studies
-After our quantitative results on behavior variability and its impact on the model, we present two case studies demonstrating its nature.
-Behavior Variability in Benign Samples.
-RobloxLauncher and OperaPatcher are two benign samples on the opposite ends of the variability spectrum.
+The Impact of Sandbox Configurations.
+Our experiments use sandbox traces obtained from commercial sandboxes with configurations unknown to us, likely optimized to identify malware from its sandbox behaviors.
+The configuration differences between sandboxes and endpoint hosts (e.g., operating systems, software versions) lead to the large dissimilarity between SB and EP traces we observed.
+This raises the question of whether a sandbox can be configured to obtain traces that approximate those from real-world hosts.
+Training on such sandbox traces will likely result in higher performance in SB->EP.
+We acknowledge this as a valid possibility, but it remains an open research question in the current landscape.
+There is no accepted standard to configure sandboxes in such a manner, and some research suggests that it would be challenging to replicate endpoint systems, e.g., simulating user interactions [35] or diverse real-world configurations [22].
+Nonetheless, currently, the research community collects training sets mainly from a single sandbox for ML-based detection solutions [55, 5, 76].
+We are unable to collect our own sandbox traces with various configurations to analyze how sandbox configurations affect detection performance, as this would cause a disparity between our endpoint and sandbox data (explained in Section IV-A).
+We highlight this as an important direction for future work.
+VII-B Case Studies
+Next, we present two case studies to illustrate the nature of behavior variability and its impact on ML models.
+Benign Behavior Variability.
+RobloxLauncher and OperaPatcher are two benign samples on the opposite ends of the behavior variability spectrum.
 RobloxLauncher accepts user input through a graphical interface, such as mouse clicks on menu items.
@@
 In its SB2 trace, the sample downloads a similar list of assets, whereas, in its SB1 trace, it creates only a log file and stops, likely because it could not access the Internet.
-Neither sandbox triggers the execution of the game launch code path, so we expect that a model trained on sandbox traces would struggle to classify endpoint traces that launch the game.
-Applied to a model trained on SB1 traces, a drift detector (Section 5.1) consistently rejects the endpoint traces of this sample.
-We believe this illustrates a broader challenge: emulating user interactions in sandboxes is challenging (liu2022enhancing, ), causing major differences in the data distribution of sandbox vs endpoint traces from benign samples.
-On the other hand, OperaPatcher performs almost the same actions in all its SB1, SB2, and EP traces: creating localization and library files to update the browser.
-The fact that this sample executes without any user input eliminates behavior (and prediction) variability across traces.
+Neither sandbox triggers the game launch execution path, so we expect a model trained on sandbox traces to struggle to classify endpoint traces that launch the game.
+Applied to a model trained on SB1 traces, a drift detector (Section V-A) consistently rejects the EP traces of this sample.
+We believe this illustrates a broader challenge: emulating user interactions in sandboxes is challenging [35], causing major differences in the data distribution of sandbox vs endpoint traces from benign samples.
+Conversely, OperaPatcher performs almost the same actions in all its SB1, SB2, and EP traces: creating localization and library files to patch a browser.
+This sample executing without any user interaction eliminates behavior (and score) variability across traces.
 Non-Malicious Malware Traces.
-In behavior-based detection, most commonly, the label of a sample (e.g., obtained from VirusTotal) is transferred to all traces from this sample, which are then used for training and testing.
+In behavioral malware detection, most commonly, the label of a sample (e.g., obtained from VirusTotal) is transferred to all traces from this sample, which are then used for training and testing.
 This practice implicitly assumes that all traces of a malware sample contain some discernible malicious activity.
-However, this can be problematic in cases where the sample has failed to execute, e.g., because its remote infrastructure is down (yong2021inside, ).
+However, this can be problematic in cases where the sample has refused or failed to execute, e.g., because its remote infrastructure is down [56].
 To assess this approach, we study the Wannacry ransomware family.
@@
 Using multiple sources, we create a list of IOCs for Wannacry and look for them in our SB1, SB2, and EP traces.
-We separate the traces into two sets based on whether they contain any IOC and then measure the average prediction scores of our models on each set.
-Based on Table 14, we observe (i) most traces contain at least one IOC (e.g., 97.7% of the EP traces), (ii) the model’s average prediction score is lower on the traces with no IOC (e.g., 97% vs.
-52% in EP->EP); however, it is still significantly higher than the averages on benign traces (13%).
+We split the traces into two sets based on whether they contain any IOC and then measure the average prediction scores of our models on each set.
+Based on Table XV, we observe (i) most traces contain at least one IOC (e.g., 97.7% of the EP traces), (ii) the model’s average prediction score is lower on the traces with no IOC (e.g., 97% vs.
+52% in EP->EP), which is still significantly higher than the averages on benign traces (13%).
 The ability to associate malware traces with no IOC with maliciousness is a strength of ML-based methods over IOC-based detectors.
-Any malware that reaches the execution stage is a potential danger, regardless of whether it successfully compromises the host.
-7.3. Sandbox Features That Do Not Generalize
-We have shown that sandbox and endpoint traces for a given sample are different. In this section, we examine why they are different and how this difference might hurt the model in the SB->EP scenario.
-Sandbox-Specific Features.
-We found several sandbox-specific features that are highly prevalent (seen in many traces) and predictive (seen only in malware traces), but occur only in one sandbox and not in the other environments.
-ML models typically exploit such features to minimize the loss and may overlook other predictive features (pezeshki2021gradient, ), causing the model to perform poorly on traces collected from a different sandbox or from endpoints.
-Table 10 presents four of these features (called artifacts) that we have found in SB1 and SB2 traces.
-We have identified over 100 such artifacts, most of which frequently co-occur.
+Any malware that reaches the execution stage is hazardous, regardless of whether it successfully compromises the host.
+VII-C Spurious Features From Sandboxes
+We have shown that sandbox and endpoint traces for a given sample are different.
+Next, we examine why they are different and how this difference may trigger unreliable predictions from the ML models in the SB->EP scenario.
+Sandbox-Specific Artifacts.
+We found several sandbox-specific features that are highly prevalent (seen in many traces) and predictive (seen only in malware traces) but occur only in one sandbox and not in the other environments.
+ML models typically exploit such features to minimize the loss and may overlook other predictive features [77], causing the model to perform poorly on traces collected from a different sandbox or from endpoint hosts.
+Table IX presents four of these features (called artifacts) that we have found in SB1 and SB2 traces.
+We have identified over 100 such artifacts, most frequently co-occurring in traces.
 For each artifact, we report its prevalence (Prv.) and its malware ratio (MalR.) in SB1, SB2, and EP training sets.
@@
 is the percentage of these samples labeled as malware.
-File Name SB1 Traces SB2 Traces EP Traces Prv. MalR. Prv. MalR. Prv. MalR. SogouExp. 9.0% 100.0% 0.2% 0.0% PersonalBankP. 3.3% 100.0% Spotify 0.1% 0.0% 3.0% 100% 0.2% 0.0% Python 3.1% 100% 0.1% 0.0%
-Table 10. Some strong malware features found in specific sandboxes that do not generalize to other environments.
-These artifacts appear to be a result of the specific configuration and particular apps pre-installed on the specific sandbox (see Appendix C).
+File Name SB1 Traces SB2 Traces EP Traces Prv.
+MalR.
+Prv.
+MalR.
+Prv.
+MalR.
+SogouExp.
+9.0% 100.0% 0.0% 0.2% 0.0% PersonalB.
+3.3% 100.0% 0.0% 0.0% Spotify 0.1% 0.0% 3.0% 100% 0.2% 0.0% Python 0.0% 3.1% 100% 0.1% 0.0%
+TABLE IX: Some strong malware features found in specific sandboxes that do not generalize to other environments.
+These artifacts appear to be a result of the specific configuration and particular apps pre-installed on the specific sandbox (see Appendix D).
 As real-world hosts and other sandboxes do not share the same configuration, this explains why an endpoint classifier trained on traces from one sandbox might fail to generalize to endpoint traces seen in the wild.
 The Impact of Sandbox Evasion.
-Recent works have measured that 40-80% of malware uses at least one evasive technique (galloro2022systematical, ) to avoid analysis or terminate early if it is running in a sandbox (mcafeeevasion, ; miramirkhani2017spotless, ).
+Sandbox evasion is a common mechanism malware authors use to make a sample’s sandbox trace dissimilar to its real-world traces.
+If running in a sandbox, evasive malware attempts to avoid analysis or terminate early [78, 21].
+In Appendix E, we present a series of quantitative experiments to understand the implications of sandbox evasion for ML-based detectors.
+We find that an ML-based detector trained on sandbox traces predicts a higher score of maliciousness for the shortest traces.
+This length bias boosts the performance in the SB->SB scenario, where very short traces are more likely from malware.
+The same bias, however, hurts the performance in the SB->EP scenario, where very short traces are not more likely from malware.
+This makes trace length a spurious correlation learned from sandbox traces that fails to generalize to endpoint traces.
+Our evidence suggests that this correlation is introduced by evasive malware, which tends to produce short sandbox traces [28].
+VII-D Invariant Learning to Counter Environment-Sensitivity
+We have shown that the environment-sensitivity of program behaviors impacts the model.
+When the classifier exploits features that are specific to the training environment (e.g., a sandbox artifact), its predictions fail to generalize.
+Table IX suggests that the artifacts of SB1 are absent in SB2, and vice versa, as they are configured sufficiently differently.
+This highlights an opportunity: learning features that are invariant across two environments (e.g., both sandboxes) might yield better generalization to endpoint detection.
+Prior work has studied invariant learning in self-supervised learning, e.g., viewpoint-invariant visual features [79].
+We employ the Siamese loss [80], which forces the model to produce similar embeddings (measured by a distance metric) on pairs of related inputs.
+In our case, each pair consists of two traces of the same sample from different sandboxes (in SB->EP) or from different EP hosts (in EP->EP).
+We aim to make the model invariant to the differences in these pairs (such as sandbox-specific artifacts) to gain robustness to variable behaviors.
+The critical question is whether invariance across SB traces translates to invariance across EP traces.
+We represent a pair of traces of a sample $P_{i}$ as ($x_{i,0},x_{i,1})$ and the set of all pairs as $\mathbf{D}$; and define the following loss function that is added to the standard ERM loss as a regularizer:
+$\mathcal{L}_{\text{inv}}=\dfrac{1}{\lvert\mathbf{D}\rvert}\sum_{(x_{i,0},x_{i,1})\in\mathbf{D}}-\mathrm{cos}(\text{enc}(x_{i,0}),\text{enc}(x_{i,1}))$
+$\mathcal{L}_{\text{final}}=\mathcal{L}_{\text{ERM}}+\alpha\mathcal{L}_{\text{inv}}$
+[…]
+[…]
+TABLE X: The impact of $\mathcal{L}_{\text{inv}}$ on endpoint detection TPR.
+[…]
+Figure 4: The model’s embeddings with increasing values of $\alpha$ for invariant learning. Sep quantifies the separability of the SB1 and SB2 embeddings. t-SNE [81] visualization.
+[…]
+Takeaways: An execution environment’s specific configuration might introduce artifacts into program traces, hurting behavioral detectors’ generalizability to other environments.
+Learning invariant features over a sample’s traces from multiple environments (e.g., two sandboxes) can enhance ML models’ robustness to such artifacts, improving their performance in real-world hosts with diverse configurations.
+VIII Discussion and Limitations
+Dataset Bias, Size and Label Noise.
+Our study has some limitations due to our endpoint data.
+First, this data was collected from hosts that use a single vendor’s anti-malware product.
+Although we cannot rule out selection bias, the fact that these hosts are located in over 100 countries in both enterprise and consumer settings suggests that our results have broad applicability.
+Second, our data was collected six years ago.
+This is a common limitation in malware studies as collecting large-scale, up-to-date real-world data is infeasible for researchers [60, 25, 39].
+Nevertheless, to our knowledge, our endpoint dataset (originally collected in [15]) is the only dataset of endpoint malware behavior analyzed in the literature.
+We focus on the gaps between sandbox-based and endpoint detection and not on the specifics of the threats of the time, making our observations still relevant today.
+Third, our data consists of only Windows hosts and lacks network-related actions.
+As 95% of malware is aimed at Windows [82] and ML-based detectors can perform well without network actions [6], we do not expect this to affect our findings.
+Moreover, our endpoint dataset is much smaller than our sandbox dataset and smaller than what is available to a security vendor, potentially causing us to underestimate the EP->EP performance in practice.
+However, consider Figure 6 where we train models in EP->EP using increasing proportions of our endpoint dataset.
+Here, the performance resulting from using 25%, 50%, and 100% of the data for training is 35%, 41%, and 46% TPR, respectively.
+This hints that even with larger endpoint dataset sizes, the performance in EP->EP is unlikely to catch SB->SB due to fundamental challenges we exposed (difficult-to-classify samples and behavior variability).
+Finally, for a given sample, labels obtained from VirusTotal (VT) are initially noisy and stabilize after around a year [83].
+In our experiments (see Section IV-A), we use the oldest label available (e.g., from old VT reports) for the training samples and the labels from VT reports collected four years later (in 2022) for the testing samples.
+This methodology approximates the realistic conditions for deploying a malware model.
+To gauge the impact of label noise on our measurements, we compare the labels of training samples from recent VT reports to the labels we used to train our models.
+In the SB and EP training sets, old and new labels agree on 98% and 97% of the samples, suggesting that the effect of label noise is minimal.
+Early Detection of Malware.
+We study detectors that use the whole execution trace of a sample, and thus can only detect malware once it terminates.
+Although vendors offer ways to undo the damage from malware after its execution, such as quarantining [8], it is more desirable to catch malware in its tracks, e.g., before a ransomware sample starts encrypting personal files.
+In that task, a detector has strictly less information available to classify a sample, which makes our results an upper bound on the performance of early malware detection.
+Fine-Tuning a Sandbox-Based Model on Endpoint Traces.
+In Appendix F, we present an experiment where we first train a NGR model only on sandbox traces; we then fine-tune this model on increasing amounts of data from our EP training set.
+This hybrid approach simulates a realistic scenario where low-cost sandbox data is supplemented with high-cost endpoint traces from the wild.
+We implement two strategies: fine-tuning all layers of the model and fine-tuning only the last layer.
+We find that adding a few EP traces improves performance, and when very few EP traces are available, fine-tuning outperforms training from scratch (3–5% higher TPR).
+Attacks Against ML.
+All ML-based detectors, especially deep-learning-based ones, are subject to adversarial attacks due to their sensitivity to input perturbations [84, 85].
+We expect our models will also be vulnerable to such attacks, e.g., an adversary can inject dummy actions into the behaviors of their malware sample designed to fool a model.
+We consider a defense against these attacks to be out of scope for this work; instead, we aim to show that naturally occurring pressures, such as environment variability, also greatly degrade the performance of ML-based malware detectors in the wild, posing a critical security and trustworthiness challenge in practice.
+Promoting Future Research.
+Our investigation relies on a dataset of endpoint traces from the wild provided by Avllazagaj et al., who collected and analyzed it in their prior work [15].
+Using this data, we have pinpointed the salient challenges in ML-based endpoint malware detection.
+Performing evaluations only on lab-based data can obscure these challenges and produce biased or unreliable solutions inapplicable to the real world [44, 43].
+Although its terms of use prevent us from sharing Avllazagaj et al.’s dataset, we built a pipeline that allows the community to perform realistic evaluations of ML-based behavioral malware detectors, aiming to minimize this bias.
+Anyone may submit to us a pre-trained detector and a feature extractor that converts a trace in the standardized format (Section IV-A) into an input to the detector.
+Upon receiving a submission, we will evaluate it on our endpoint traces dataset and return the submitter a detailed performance report to guide their research.
+Further, if the submitter wishes, we will publicly list their results and submission details on the leaderboard on our website (https://malwaredetectioninthewild.github.io).
+To assist participants, we will release the sandbox dataset we collected and the metadata for over 200K samples used in our work.
+The details of these artifacts can be found in Appendix G.
+Please refer to our website for technical details, data access requests, and detector submission instructions.
+This pipeline offers a blueprint for security vendors to drive research without publicly sharing their sensitive data.
+We hope to spearhead it into more applications of ML for security where evaluation problems frequently hinder real progress.
+Implications for Future Work.
+As described in Section III-A, malware detection in practice involves a pipeline of techniques.
+We believe ignoring this pipeline and treating each component as a standalone solution causes critical evaluation problems.
+Consider ML-based static malware detectors, an active research area [49, 4, 26].
+They are ideally evaluated after filtering out the test samples that can already be detected by the pipeline components preceding static ML, such as existing blocklists, allowlists, and static signatures.
+However, to our knowledge, this is not an established practice, which might cause an overestimation of real-world performance (similar to our findings in Section VI).
+Our endpoint data allowed us to isolate and realistically evaluate ML-based behavioral detectors, revealing previously unknown challenges.
+This, however, still falls short of an ideal end-to-end evaluation.
+Instead of measuring the success of individual components, quantifying the overall effectiveness of a pipeline yields more relevant insights for users.
+For example, if static components detect 95% of the active malware at a given time and the dynamic components detect 80% of the remaining malware, the overall pipeline would be at 99% TPR.
+Moreover, the inter-dependence in components introduces poorly understood challenges, e.g., the trade-off between minimizing the pipeline’s false positives and classifying more samples statically to reduce infections.
+Due to our dataset limitations, we are unable to perform such analyses.
+We recommend that future research approach malware detection holistically, treating it as a pipeline, and collect suitable datasets for performing end-to-end evaluations.
+Furthermore, prior work on ML-based behavioral detectors has focused strictly on a sandbox-based scenario where training and testing traces are collected from the same sandbox.
+This obscures the impact of sandbox configurations on models’ robustness when deployed on traces from diverse environments, such as real-world hosts, as shown in Section VII.
+Our work had to rely on traces from two third-party sandboxes with unknown configurations, which prevented us from conducting controlled experiments to study this impact, e.g., to increase the performance in SB->EP.
+Instead, we show that invariant learning over traces on these two sandboxes might increase robustness.
+While sandbox configuration is a well-studied topic in malware analysis [35, 21, 22], its role in generating training sets for ML-based detectors has been overlooked.
+To address this gap, we recommend that future work tackle realistic scenarios where behavioral models are deployed in unpredictable environments, emphasizing the need for robustness.
+Developing principled methods for configuring sandboxes for this scenario is a promising research direction.
+IX Conclusion
+Behavioral malware detection serves as a last line of defense at endpoint hosts, providing security when all other measures have failed.
+Malware samples that breach this last line cause real-world infections and harm.
+A long-standing ambition is to find the best way to detect such samples.
+Through a systematic exploration of different scenarios, we provide some clarity.
+State-of-the-art ML-based detectors trained on sandbox execution traces degrade severely when deployed at real-world endpoints.
+Though training on data from endpoints leads to better performance, it still falls far short of expectations established in prior evaluations.
+We characterize various challenges ML approaches face in this domain and explore promising techniques targeting them, achieving moderate improvements.
+Ultimately, this task remains challenging.
+By exposing evaluation pitfalls and ML shortcomings, we call on the community to seek new solutions to these security-threatening problems.
+X Acknowledgements
+This research was supported by the US Intelligence Community Postdoctoral Fellowship (Kaya), the US Department of Defense (Dumitraş), UK EPSRC Grant EP/X015971/1 (Pierazzi), US National Science Foundation Grants CNS-2154873 (Wagner) and CNS-2327427 (Botacin), and generous research awards from Google (Cavallaro and Wagner) and Amazon (Dumitraş and Wagner).
+We would like to thank VirusTotal for granting us academic access to their platform.
+Finally, we thank Omer Faruk Akgul and David Acs for their support in the earlier phases of our research and Erin Avllazagaj for providing us access to their program traces in the wild dataset.
+Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the supporting organizations.
+Appendix A Trace Standardization
+A-A Technical Details
+Our EP traces contain high-level information the vendor can seamlessly collect and analyze at endpoint hosts, including file, process, and mutex creations, registry key creations and deletions, and process injections.
+Sandboxes also record lower-level information (such as memory dumps [86]) as they are less computationally constrained.
+In a standardized trace, first, we only keep the action types that all formats in our datasets share: file creation, registry key creation and deletion, process creation and injection, and mutex creation.
+Second, we clean up the strings (e.g., file names) in each trace by removing white spaces, capitalization, punctuation, non-ASCII characters, and so on.
+Third, we replace certain file and directory names to minimize differences caused by operating system versions or logging conventions.
+Fourth, we use regular expressions to replace MAC addresses, Windows security and resource identifiers, hashes (SHA-256, SHA-1, MD5), and epoch timestamps with special tokens (e.g.
+<macaddress>), to prevent introducing artifacts to our ML models.
+Fifth, we tokenize the entries in each trace (e.g., split a full file path into its components) and save a trace as a sequence of tokens, following [6].
+We will release our sandbox dataset in this standardized format.
+Moreover, this format enables researchers to submit their detectors to be evaluated against our endpoint dataset (detailed in Section VIII) by providing them with an expected input format their detectors should accept.
+When we train a model, we only keep the top-10K tokens in terms of frequency in the training traces and replace the remaining tokens with special tokens, such as <rare_file_name>, again following [6, 7].
+This makes the task more suitable for learning by eliminating uninformative tokens and reducing the feature dimensionality.
+A-B The Performance Impact
+We run the following experiment to understand whether our standardization hurts detection performance.
+We trained NGR models on unstandardized SB1 traces, which contain more action types (e.g., file reads, registry reads, and modules loaded) than standardized traces.
+We also kept the top-25K tokens (instead of the top-10K) from these traces for tokenization.
+The best model achieves 97.8% TPR on the SB1 test set (up from 95.0% when trained on standardized traces).
+However, its performance is 53.4% TPR on the SB2 test set (down from 60.7% when trained on standardized traces).
+This suggests that trace standardization moderately hurts the in-domain performance (on the test traces from the training sandbox) while improving generalization to other domains (to traces from other sandboxes).
+We hypothesize that standardization reduces overfitting to domain-specific features (discussed in Section VII-C), which is desirable in our study as we aim to maximize the performance of a sandbox-based model on endpoint traces.
+Note that, even after removing some features from sandbox traces, the classifier’s performance is still competitive with prior work (e.g., Neurlux [6]) that uses the full sandbox report.
+This is because different features (e.g., registry reads and registry creations) are already correlated and have redundancy, making each additional feature less effective.
+Appendix B Further Machine Learning Details
+NGR (based on Maldy [5]).
+We convert each trace into a list of 2-grams.
+For example, the file path a/b/c.jpg is turned into three 2-grams: <a/b>, <b/c> and <c.jpg>.
+As this results in an intractable number of unique 2-grams (mostly very rare), we apply the hashing trick [87] that assigns a numerical value up to $2^{14}$ to each 2-gram.
+The final feature vector—$x$—for a trace is a $2^{14}$-dimensional vector and each dimension is set to the number of occurrences of the corresponding 2-gram in the trace.
+We use a ResNet-based architecture [88] to train on these features.
+HYB (based on Neurlux [6]).
+We treat a trace as a natural language document and train an attention-convolution-hybrid sequence classification model.
+This approach eliminates the need for feature engineering (unlike the n-gram approach) and can extract useful features from long sequences thanks to the attention mechanism.
+ATT (based on Nebula [7]).
+We treat a trace the same way as HYB (a sequence) and train a self-attention transformer-based architecture that is claimed to be robust to heterogeneous information (  e.g., different report format).
+Overall, these approaches represent an increasing level of complexity, NGR being the simplest and most traditional and ATT being the most advanced.
+Although more advanced models seem to perform better in sandbox-based scenarios, we are interested in whether this trend changes in the endpoint scenario.
+Appendix C How to Do Model Selection in SB->EP
+Following our observation in Section V-A, in this section, we assess whether finding a better model for SB->EP is possible by selecting the models using traces from an unseen sandbox (not used for training).
+We rank our models (100+ in each setting) based on their SB1, SB2, and EP performances (TPR) and compute Spearman’s correlation coefficients between these rankings.
+Table XI reveals that (except for one setting) (i) the rankings based on the training sandbox correlate poorly (sometimes negatively) with EP rankings, (ii) rankings based on an unseen sandbox correlate more strongly with EP rankings.
+For example, for ATT trained on SB1, the rankings based on SB1 and SB2 have $0.03$ and $0.45$ correlation with the rankings based on EP, respectively.
+These results support our claim that it yields better results when model selection in SB->EP is performed using traces from a different, unseen sandbox that is not used for training.
+Trained on SB1 Trained on SB2 SB1-SB2 SB1-EP SB2-EP SB1-SB2 SB1-EP SB2-EP NGR 0.47 0.03 0.45 --0.24 0.39 --0.46 HYB 0.88 0.84 0.75 --0.17 0.59 0.33 ATT 0.56 --0.17 0.31 0.51 0.58 0.35
+TABLE XI: Ranking correlations of sandbox-trained models according to their performances (TPR) on different test sets.
+Appendix D Case Studies on Sandbox-Specific Artifacts
+Here, we dive deeper into the sandbox-specific artifacts we identified in Table IX.
+SogouExplorer is a Chinese web browser that exists only in malware traces (100% MalR.) in SB1; whereas it does not exist in any SB2 traces and very few EP traces.
+The samples that interact with it mainly belong to families such as Sivis and Memery, all tagged as file infectors that attach their code to other programs.
+Considering that a Chinese vendor developed SB1, we believe they pre-installed this browser on their sandboxes to generate an analysis environment representative of Chinese hosts.
+This, however, creates features specific to SB1 as samples interact with the programs in the environment.
+Although this artifact exists in a few endpoint traces from hosts in China, its prevalence is almost zero.
+PersonalBankPortal, according to our research, is a program distributed by a Chinese bank to its customers.
+The samples that inject into this program belong to families such as Tinba and Ramnit, all considered as banking trojans that specifically ex-filtrate banking data.
+We believe the vendor pre-installs this program to lure malware samples into exhibiting their behaviors.
+Although this practice is useful for analyzing a sample [56] (and for SB->SB), it causes artifacts that are rarely observed in the wild.
+Among the artifacts found in SB2, Spotify is a popular music streaming service, and Python is the interpreter for Python programming language.
+Both programs are targeted and injected by file infectors, similar to SogouExplorer in SB1.
+These programs are much less prevalent in endpoint traces than in SB2 traces.
+We believe the SB2 vendor, based in the US, pre-installs them to create an environment representative of the hosts in the US.
+Appendix E The Impact of Sandbox Evasion
+Recent works have measured that 40-80% of malware uses at least one evasive technique [17] to avoid analysis or terminate early if it is running in a sandbox [78, 21].
 Evasion makes sandbox traces dissimilar to endpoint traces.
@@
 We explore the impact of sandbox evasion.
-We use a standard heuristic and treat a sample as evasive if the number of actions it performs is too low (kuchler2021does, ; balzarotti2010efficient, ; miramirkhani2017spotless, ).
+We use a standard heuristic and treat a sample as evasive if the number of actions it performs is too low [28, 16, 21].
 We first find the malware families common between our EP and SB1 test sets (30 total).
@@
 Families #Fams #EP #SB1 SB1->EP SB1->SB1 All 30 169 5.8K 23.1% 86.4% Evasive 9 49 2.9K 20.4% 89.8% Non-Eva. 21 120 2.9K 23.3% 83.0%
-Table 11. The performance (TPR) on malware families more (Evasive) or less (Non-Eva) likely to be evading sandboxes.
+TABLE XII: The performance (TPR) on malw. families more (Evasive) or less (Non-Eva) likely to be evading sandboxes.
 The trace length of a sample monotonically increases as its execution continues.
-This means our measurements might be confounded by the discrepancies between the execution durations in SB1 and endpoint hosts.
+This means the discrepancies between the execution durations in SB1 and endpoint hosts might confound our measurements.
 Before presenting our results, we confirm this is unlikely to be the case: 82% of our EP traces are from executions that lasted less than 90 seconds.
-Although we do not know the exact configuration of SB1, allowing one to two minutes of execution is standard for most sandboxes in practice (kuchler2021does, ).
-Based on the results in Table 11, we observe: (i) SB1->EP performance is higher (by 3%) on non-evasive families than on evasive families; (ii) SB1->SB1 performance is significantly higher (by 7%) on evasive families than on non-evasive families.
+Although we do not know the exact configuration of SB1, allowing one to two minutes of execution is standard for most sandboxes in practice [28].
+Based on the results in Table XII, we observe: (i) SB1->EP performance is higher (by 3%) on non-evasive families than on evasive families; (ii) SB1->SB1 performance is significantly higher (by 7%) on evasive families than on non-evasive families.
 This suggests that the classifier exploits short traces (evasiveness) as a feature, which is beneficial for SB->SB, though it does not generalize to SB->EP.
@@
 Evidence suggests this is introduced by evasive malware that tends to produce short sandbox traces in the same execution time.
-Methods preventing the model from learning such correlations (kirichenko2023last, ; pezeshki2021gradient, ) offer a promising next step.
-7.4. Combatting Environment-Sensitivity With Environment-Invariant Features
-We have shown that the environment-sensitivity of program behaviors impacts the model.
-When the classifier exploit features that are specific to the training environment (e.g., trace length), its predictions fail to generalize.
-Table 10 suggests that the artifacts of SB1 are absent in SB2, and vice versa, as they are configured sufficiently differently.
-This highlights an opportunity: learning features that are invariant across two environments (e.g., both sandboxes) might yield better generalization to endpoint detection.
-Prior work has studied invariant learning in self-supervised learning, e.g., viewpoint-invariant visual features (purushwalkam2020demystifying, ).
-We employ the Siamese loss (chen2021exploring, ), which forces the model to produce similar embeddings (measured by a distance metric) on pairs of semantically related inputs.
-In our case, each pair consists of two traces of the same sample from different sandboxes (in SB->EP) or from different EP hosts (in EP->EP).
-To gain robustness to variable behaviors, we aim to make the model invariant to the differences in these pairs (such as sandbox-specific artifacts).
-The critical question is whether invariance across SB traces translates to invariance across EP traces.
-We represent a pair of traces of a sample $P_{i}$ as ($x_{i,0},x_{i,1})$ and the set of all pairs as $\mathbf{D}$; and define the following loss function that is added to the standard ERM loss as a regularizer:
-$\mathcal{L}_{\text{inv}}=\dfrac{1}{\lvert\mathbf{D}\rvert}\sum_{(x_{i,0},x_{i,1})\in\mathbf{D}}-\mathrm{cos}(\text{enc}(x_{i,0}),\text{enc}(x_{i,1}))$
-$\mathcal{L}_{\text{final}}=\mathcal{L}_{\text{ERM}}+\alpha\mathcal{L}_{\text{inv}}$
-[…]
-[…]
-Table 12. The impact of $\mathcal{L}_{\text{inv}}$ on endpoint detection TPR.
-[…]
-Figure 5. The model’s embeddings with increasing values of $\alpha$ for invariant learning. Sep quantifies the separability of the SB1 and SB2 embeddings. t-SNE (van2008visualizing, ) visualization.
-[…]
-8. Endpoint Malware Detection Leaderboard
-Our investigation has relied on a dataset of endpoint traces from the wild provided to us by Avllazagaj et al., who collected and analyzed it in their prior work (avllazagaj2021malware, ).
-We have pinpointed the salient challenges in ML-based endpoint malware detection using this data.
-Performing evaluations only on lab-based data can obscure these challenges and produce biased solutions inapplicable to the real world (cherubin2022online, ; jacobs2022ai, ).
-To minimize this bias, we designed a leaderboard for realistic evaluations of behavior-based malware detectors (malwaredetectioninthewild.github.io (leaderboard, )).
-Prior leaderboards such as RobustBench (croce2021robustbench, ) or TrojAI (trojai, ) have motivated research into critical and difficult problems and fueled many innovations in ML.
-Anyone may submit a pre-trained detector and a feature extractor that converts a trace in the standardized format into an input to the detector.
-We evaluate each submitted detector on our endpoint traces and list it on our leaderboard.
-Although we are unable to share Avllazagaj et al.’s dataset, we will release the sandbox dataset we collected and the metadata for over 200K samples used in our work to assist the participants.
-The details of these artifacts can be found in §E.
-This format offers a blueprint for security vendors to guide researchers without publicly sharing their sensitive data.
-We hope to spearhead it into more applications of ML for security.
-9. Discussion
-Threats to Validity and Limitations.
-Our study has some limitations due to the nature of our endpoint data.
-First, this data was collected at the hosts of a single security vendor.
-Although we cannot rule out selection bias, the fact that these hosts are located in over 100 countries in both enterprise and consumer settings suggests that our results have broad applicability.
-Second, our data was collected five years ago.
-This is a common limitation in malware detection studies as collecting large-scale up-to-date data has not been feasible for researchers (yang2021cade, ; barbero2022transcending, ; chen2023continuous, ).
-We focus on the gaps between sandbox-based and endpoint detection and not on the specifics of the threats that existed at the time, making our observations relevant to today’s malware detection landscape.
-Third, our data consists of only Windows hosts and lacks any network-related actions.
-As 95% of malware is aimed at Windows (windowsmalware, ) and ML-based detectors can perform well without network actions (jindal2019neurlux, ), we do not expect this to affect our findings.
-Early Detection of Malware.
-In our study, a detector uses the whole execution trace of a sample that terminates in a host.
-Although vendors offer ways to undo the damage from malware after its execution, e.g., quarantining (quarantine, ), it is more desirable to catch malware in its tracks.
-This would allow for containing the harm of a malware sample by cutting its execution short, e.g., before a ransomware sample starts encrypting personal files.
-As the detector has strictly less information available in this task, our results can be viewed as an upper bound on the performance of early malware detection.
-In consequence, we do not expect a detector that is trained only on sandbox traces to perform well in this setting.
-Fine-Tuning a Sandbox-Based Model on Endpoint Traces.
-In Appendix D, we present an experiment where we first train a NGR model on sandbox traces and then fine-tune this model on increasing amounts of data from our EP training set.
-This hybrid approach simulates a realistic scenario where low-cost sandbox data is supplemented with high-cost endpoint traces from the wild.
-We evaluate two strategies: fine-tuning all layers of the model and fine-tuning only the last layer.
-We find that adding a few EP traces improves performance, and when very few EP traces are available, fine-tuning outperforms training from scratch (3–5% higher TPR).
-Attacks on ML.
-All ML-based detectors, especially deep learning ones, are subject to adversarial attacks due to their sensitivity to input perturbations (szegedy2014intriguing, ; grosse2017adversarial, ).
-We expect our models will also be vulnerable to such attacks, e.g., an adversary can inject dummy actions into the behaviors of their malware sample designed to fool a model.
-We consider a defense against these attacks to be out of scope for this work; instead, we aim to show that naturally occurring factors, such as the environment variability, also degrade the performance of ML-based detectors in the wild.
-10. Conclusion
-Malware detection with dynamic features serves as a last line of defense at endpoints, providing security when all other measures have failed.
-Malware samples that breach this last line cause real-world infections and harm.
-A long-standing ambition is to find the best way to detect such samples.
-Through a systematic exploration of different scenarios, we provide some clarity.
-Existing detectors trained on sandbox executions degrade when deployed at endpoints.
-Training on real-world behaviors performs much better, though it necessitates collecting data from real hosts.
-We pinpoint various challenges in endpoint detection and propose ML techniques targeting them, yielding moderate improvements.
-Ultimately, this task remains challenging, and through our leaderboard, we hope to guide the community to develop effective and applicable solutions.
-11. Acknowledgements
-This research was supported by the Intelligence Community Postdoctoral Fellowship, the U.S.
-Department of Defense, EPSRC Grant EP/X015971/1, National Science Foundation Grant CNS-2154873, and generous gifts from Google and Amazon.
-We would like to thank VirusTotal for granting us academic access to their platform, which allowed us to curate and label our datasets.
-Finally, we thank Omer Faruk Akgul and David Acs for their support in the earlier phases of our research and Erin Avllazagaj for providing us with access to their program traces in the wild dataset.
-Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the supporting organizations.
-Appendix A Trace Standardization
-First, we only keep the action types that all formats in our datasets share, which are file creation, registry key creation and deletion, process creation and injection, and mutex creation.
-Second, we clean up the strings (e.g., file names) in each trace by removing white spaces, capitalization, punctuation, non-ASCII characters, and so on.
-Third, we replace certain file and directory names to minimize differences caused by operating system versions or logging conventions.
-Fourth, we use regular expressions to replace MAC addresses, Windows security and resource identifiers, hashes (SHA-256, SHA-1, MD5), and epoch timestamps with special tokens (e.g.
-<macaddress>), to prevent introducing artifacts to our ML models.
-Fifth, we tokenize the entries in each trace (e.g., split a full file path into its components) and save a trace as a sequence of tokens, following (jindal2019neurlux, ).
-This standardized format allows us to implement a leaderboard for endpoint malware detection (detailed in Section 8) by providing all participants with an expected input format that their detectors should accept.
-When we train a model, we only keep the top-10K tokens in terms of frequency in the training traces and replace the remaining tokens with special tokens, such as <rare_file_name>, again following (jindal2019neurlux, ; trizna2023nebula, ).
-This makes the task more suitable for learning by eliminating uninformative tokens and reducing the feature dimensionality.
-Appendix B Further Machine Learning Details
-NGR (based on Maldy (karbab2019maldy, )).
-We convert each trace into a list of 2-grams.
-For example, the file path a/b/c.jpg is turned into three 2-grams: <a/b>, <b/c> and <c.jpg>.
-As this results in an intractable number of unique 2-grams (mostly very rare), we apply the hashing trick (weinberger2009feature, ) that assigns a numerical value up to $2^{14}$ to each 2-gram.
-The final feature vector—$x$—for a trace is a $2^{14}$-dimensional vector and each dimension is set to the number of occurrences of the corresponding 2-gram in the trace.
-We use a ResNet-based architecture (gorishniy2021revisiting, ) to train on these features.
-HYB (based on Neurlux (jindal2019neurlux, )).
-We treat a trace as a natural language document and train an attention-convolution-hybrid sequence classification model.
-This approach eliminates the need for feature engineering (unlike the n-gram approach) and can extract useful features from long sequences thanks to the attention mechanism.
-ATT (based on Nebula (trizna2023nebula, )).
-We treat a trace the same way as HYB (a sequence) and train a self-attention transformer-based architecture that is claimed to be robust to heterogeneous information (  e.g., different report format).
-Overall, these approaches represent an increasing level of complexity, NGR being the simplest and most traditional and ATT being the most advanced.
-Although more advanced models seem to perform better in sandbox-based scenarios, we are interested in whether this trend changes in the endpoint scenario.
-Appendix C Case Studies on Sandbox-Specific Artifacts
-Here, we dive deeper into the sandbox-specific artifacts we identified in Table 10.
-SogouExplorer is a Chinese web browser that exists only in malware traces (100% MalR.) in SB1; whereas it does not exist in any SB2 traces and very few EP traces.
-The samples that interact with it mostly belong to families such as Sivis and Memery, all tagged as file infectors that attach their code to other programs.
-Considering that SB1 was developed by a Chinese vendor, we believe that they pre-installed this browser on their sandboxes to generate an analysis environment representative of Chinese hosts.
-This, however, creates features specific to SB1 as samples interact with the programs in the environment.
-Although this artifact exists in a few endpoint traces from hosts in China, its prevalence is almost zero.
-PersonalBankPortal, according to our research, is a program distributed by a Chinese bank to its customers.
-The samples that inject into this program belong to families such as Tinba and Ramnit, all considered as banking trojans that specifically ex-filtrate banking data.
-We believe the vendor pre-installs this program to lure malware samples into exhibiting their behaviors.
-Although this practice is useful for analyzing a sample (yong2021inside, ) (and for SB->SB), it causes artifacts that are rarely observed in the wild.
-Among the artifacts found in SB2, Spotify is a popular music streaming service, and Python is the interpreter for Python programming language.
-Both programs are targeted and injected by file infectors, similar to SogouExplorer in SB1.
-These programs are much less prevalent in endpoint traces than in SB2 traces.
-We believe the SB2 vendor, which is based in the US, pre-installs them to create an environment representative of the hosts in the US.
-Appendix D Fine-Tuning a Sandbox-Based Model on Endpoint Traces
+Methods preventing the model from learning such correlations [72, 77] offer a promising next step.
+Figure 5: Comparing the correlations between trace length and malware-ness prediction scores of the model.
+Malware ratio is the ground truth ratio of malware traces of a certain length among all traces of that length in a dataset.
+Appendix F Fine-Tuning a Sandbox-Based Model on Endpoint Traces
 For the experiments in Figure 6, we select two endpoint traces per sample and implement two fine-tuning strategies: (i) freezing the encoder layers $enc$ of our model and tuning only the classification layer $g$, and (ii) tuning all layers without freezing.
@@
 These experiments show that starting from a well-performing model in SB->EP is beneficial in regimes with limited EP data, highlighting a promising direction for future work.
-Appendix E More Details on the Leaderboard
-The following artifacts will be released to the participants:
+Figure 6: The results of fine-tuning a sandbox-based model in SB->EP on endpoint traces. The dashed line indicates our best model SB->EP. Experiments on NGR.
+Sel.
+Test Set Trained on SB1 Trained on SB2 Trained on EP SB1->SB1 SB1->SB2 SB1->EP SB2->SB1 SB2->SB2 SB2->EP EP->EP NGR HYB ATT NGR HYB ATT NGR HYB ATT NGR HYB ATT NGR HYB ATT NGR HYB ATT NGR HYB ATT SB1 99.1 99.0 98.9 93.8 91.3 92.6 76.1 70.7 72.8 89.5 83.4 89.1 97.7 94.1 96.9 72.0 68.2 70.8 85.6 85.2 84.3 SB2 99.0 98.6 98.8 94.9 93.2 93.5 76.6 71.3 74.4 81.7 71.5 86.4 98.7 98.5 98.2 71.6 66.7 72.7 85.4 85.9 84.8 EP 99.0 98.6 98.7 92.9 91.2 92.9 78.4 74.1 74.9 84.4 71.8 88.3 98.0 97.8 98.0 74.6 72.0 73.4 87.5 86.8 86.8
+TABLE XIII: The performance (AUC%) of three ML approaches (NGR, HYB, ATT) in seven detection scenarios (based on Table II).
+The models are selected using the test set in each row, and the average AUC of the top 20 models is reported.
+Appendix G More Details on the Artifact Release
+We will release the following artifacts to the community (visit our website for more details https://malwaredetectioninthewild.github.io/):
 Sandbox Dataset.
-We will release the training portion of our sandbox datasets in Table 3, stored in the standardized format discussed in Section 4.1.
-We avoid releasing the testing portion to prevent participants from obtaining an impractical advantage by training their detectors on it (will be released at a later date).
+We will release the training portion of our sandbox datasets in Table III, stored in the standardized format discussed in Section IV-A.
+We avoid releasing the testing portion publicly to prevent researchers who wish to participate in our realistic evaluation leaderboard from obtaining an impractical advantage by training their detectors on it (will be released if requested in exchange for being excluded from the leaderboard).
 This will level the playing field for participants and ensure they all have access to the same sandbox data for development.
-Further, as this data contains traces from two sandboxes, participants can leverage our observation in Section 5 and tune their models on a second sandbox, not seen during training, for potentially higher endpoint performance.
-Moreover, they can also apply invariance learning techniques, which have been shown promising in Section 7.4.
+Further, as this data contains traces from two sandboxes, participants can leverage our observations in Section V and tune their hyper-parameters on a second sandbox, not seen during training, for improving the endpoint performance or train their models on data from both sandboxes.
+They can also apply invariant learning techniques, which have shown promising results in Section VII-D.
 Sample Metadata.
-We will release the metadata relating to the training samples, including their SHA-256 hashes, ground truth labels, family tags (if malware), and first-seen timestamps.
-Moreover, our metadata also tags the source of a sample, e.g., EMBER (anderson2018ember, ), SOREL (harang2020sorel20m, ), or our endpoint dataset.
-This allows participants to easily obtain realistic priors over malware families for endpoint detection, and make use of our findings and improvements in Section 6.3.
-Moroever, they can also create realistic testing distributions over samples using these malware family priors (and avoid the problems we discussed in Section 6.
-Appendix F Additional Tables and Figures
+We will release the metadata relating to the training samples in all datasets, including their SHA-256 hashes, ground truth labels, family tags (if malware), and first-seen timestamps.
+Additionally, our metadata also annotates the source of a sample, e.g., EMBER [49], SOREL [53], or our endpoint dataset.
+This allows researchers to compute realistic priors over malware families for endpoint detection and make use of our findings and improvements in Section VI-C.
+Moreover, they can also create realistic testing distributions over samples using these malware family priors (and avoid the problems we discussed in Section VI.
+Appendix H Additional Tables and Figures
 Statistics on the Endpoint Traces
@@
 (b) The average ratio of seen traces per sample as a function of time.
-Figure 6. Statistics on the endpoint traces in our dataset.
+Figure 7: Statistics on the endpoint traces in our dataset.
+Soft Labeling Function Refer to caption
+Figure 8: Our function for assigning probabilistic (soft) labels to each sample based on the number of VirusTotal detections.
+Histogram of Prediction Score Standard Deviations on Endpoint Traces of the Same Sample for Models Trained With the Invariance Loss.
+Figure 9: Prediction score standard deviations on the endpoint traces of each sample in the EP test set; for a model in SB1->EP (left), and EP->EP (right).
 Top Malware Families and Benign Publishers
@@
 2.2% 0.0% Yandex 1.0% 2.7% Zoom 2.0% 0.3%
-Table 13. Top malware families (top) and benign (bottom) publishers in our EP (left) and SB (right) test sets, along with their shares in each dataset.
-Soft Labeling Function Refer to caption
-Figure 7. Our function for assigning probabilistic (soft) labels to each sample based on the number of VirusTotal detections.
+TABLE XIV: Top malware families (top) and benign (bottom) publishers in our EP (left) and SB (right) test sets, along with their shares in each dataset.
 Case Study on Wannacry and Its Indicators-Of-Compromise
 Trace Type Ratio Avg. Pred. Score SB1 SB2 EP SB1->EP SB2->EP EP->EP Any IOC 89.9% 100% 97.7% 82.0% 75.6% 97.4% No IOC 10.1% 0.0% 2.3% 77.7% 67.7% 51.8%
-Table 14.
-For Wannacry traces in our SB1, SB2, and EP datasets, we first measure the ratio of traces with at least one and no IOC.
+TABLE XV: For Wannacry traces in our SB1, SB2, and EP datasets, we first measure the ratio of traces with at least one and no IOC.
 We then measure the average prediction score of our NGR models in SB1->EP, SB2->EP and EP->EP on these traces.
-Rank Test Set Trained on SB1 Trained on SB2 Trained on EP SB1->SB1 SB1->SB2 SB1->EP SB2->SB1 SB2->SB2 SB2->EP EP->EP NGR HYB ATT NGR HYB ATT NGR HYB ATT NGR HYB ATT NGR HYB ATT NGR HYB ATT NGR HYB ATT SB1 99.1 99.0 98.9 93.8 91.3 92.6 76.1 70.7 72.8 89.5 83.4 89.1 97.7 94.1 96.9 72.0 68.2 70.8 85.6 85.2 84.3 SB2 99.0 98.6 98.8 94.9 93.2 93.5 76.6 71.3 74.4 81.7 71.5 86.4 98.7 98.5 98.2 71.6 66.7 72.7 85.4 85.9 84.8 EP 99.0 98.6 98.7 92.9 91.2 92.9 78.4 74.1 74.9 84.4 71.8 88.3 98.0 97.8 98.0 74.6 72.0 73.4 87.5 86.8 86.8
-Table 15.
-The performance (AUC%) of three ML approaches (NGR, HYB, ATT) in seven detection scenarios (based on Table 2).
-In each row, the models are ranked based on the test set, and the average AUC of top-$20$ models is reported.
-Trace Length Bias of the Sandbox-Based Model
-Figure 8.
-Comparing the correlations between trace length and malware-ness prediction scores of the model.
-Malware ratio is the ground truth ratio of malware traces of a certain length among all traces of that length in a dataset.
-Histogram of Prediction Score Standard Deviations on Endpoint Traces of the Same Sample for Models Trained With the Invariance Loss.
-Figure 9. Prediction score standard deviations on the endpoint traces of each sample in the EP test set; for a model in SB1->EP (left), and EP->EP (right).
-Fine-Tuning a Sandbox-Based Model on Endpoint Traces Refer to caption
-Figure 10. The results of fine-tuning a sandbox-based model in SB->EP on endpoint traces. The dashed line indicates our best model SB->EP. Experiments on NGR.